### v1

Abstract
In general-sum stochastic games, a stationary Stackelberg equilibrium (SSE) does not always exist, in which the leader maximizes leader’s return for all the initial states when the follower takes the best response against the leader’s policy. Existing methods of determining the SSEs require strong assumptions to guarantee the convergence and the coincidence of the limit with the SSE. Moreover, our analysis suggests that the performance at the fixed points of these methods is not reasonable when they are not SSEs. Herein, we introduced the concept of Pareto-optimality as a reasonable alternative to SSEs. We derive the policy improvement theorem for stochastic games with the best-response follower and propose an iterative algorithm to determine the Pareto-optimal policies based on it. Monotone improvement and convergence of the proposed approach are proved, and its convergence to SSEs is proved in a special case.
1 Introduction
A Markov decision process (MDP) is a mathematical framework modeling the decision-making of the agent in dynamic environments, which serves as the theoretical foundation of reinforcement learning (RL). The evaluation value of the agent’s policy is determined by assessing the state-action sequences generated by the policy in the environment with the reward function. This evaluation represents the expectation of cumulative discounted reward. An optimal policy is defined as one that maximizes the evaluation value.
One of the extensions of MDPs for the multi-agent is a stochastic game, also known as a Markov game, where multiple agents with individual reward functions simultaneously try to develop policies that maximize their expected cumulative discounted rewards in one environment. Stochastic games are initially proposed in game theory, and the solution is defined as a tuple of policies that form an equilibrium, such as a Nash equilibrium.The problem setting is characterized by the relationship in the reward functions, such as the zero-sum setting and the fully cooperative setting.
In this study, we addressed the problem of determining a Stackelberg equilibrium in 2-player stochastic games with general relationships in the rewards. A Stackelberg equilibrium is defined as a tuple of strategies of two types of agents, “leader” and “follower,” such that the leader’s strategy maximizes the leader’s payoff when the follower always takes the strategy that is the best response with respect to the follower’s payoff against the leader’s strategy. We aim to find the stationary policy that forms the Stackelberg equilibrium in general-sum stochastic games, that is called SSE. We assume that the follower’s best responses are always computable and the environment is known.
Under an identical setting as ours, Bucarey et al. (2022) proposed an algorithm for finding the SSE. However, this algorithm requires strong assumptions to guarantee convergence to the SSEs. Moreover, this algorithm can converge to low-quality policies when it does not converge to the SSEs Bucarey et al. (2019). As an SSE does not always exist in the general-sum setting Bucarey et al. (2022), the algorithm should converge to the SSE if it exists; otherwise, the algorithm must converge to policies that achieve reasonably good performance. We have developed such an algorithm.
Our contributions are summarized as follows.
First, theoretical analyses of the method proposed by Bucarey et al. (2022) reveal that its fixed point is not necessarily a reasonable policy when it is not guaranteed to converge to the Stackelberg equilibrium. (Section 5.1)
Second, a new sufficient condition for the existence of the SSE includes the one derived in Bucarey et al. (2022). A necessary and sufficient condition for the case of the myopic leader is derived. (Section 5.2)
Third, a concept of the Pareto-optimality is introduced as a reasonable alternative to the SSE. There always exist policies that realize the Pareto-optimal value functions or their neighborhood with arbitrary precision. Moreover, it agrees with the SSE if the equilibrium exists. (Section 6.1)
Lastly, a policy improvement theorem for general-sum SSGs is derived (Section 6.2), based on which a novel algorithm is proposed. Monotone improvement of the state values and convergence of our algorithm are proven in this study (Section 6.3).
2 Preliminaries
2.1 Markov Decision Process
An MDP is a stochastic process where the state transition depends on the agent’s action and the rewards are defined. MDPs are defined as a tuple $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$: a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, an initial state distribution $\rho\in\Delta(\mathcal{S})$, a bounded reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a discount rate $\gamma\in[0,1)$, where $\Delta(\mathcal{S})$ is a set of probability distributions on $\mathcal{S}$.
The agent determines its action following its Markov policy. Let $\mathcal{W}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ be a set of (stochastic) decision rules, then at the time step $t\in\mathbb{N}$, the agent has the decision rule $f_{t}\in\mathcal{W}$ and selects the action $a_{t}\in\mathcal{A}$ in the current state $s_{t}\in\mathcal{S}$ as $a_{t}\sim f_{t}(s_{t})$. The conditional probability is denoted as $f_{t}(a_{t}|s_{t})$. If the decision rule is equal at any time step ($f_{t}=f$), the Markov policy is called a stationary policy and denoted by a single decision rule $f\in\mathcal{W}$.
The next state $s_{t+1}$ is stochastically determined as $s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})$ when the agent selects its action $a_{t}$ in the state $s_{t}$. Repeating such transitions with one stationary policy $f\in\mathcal{W}$ generates the state-action sequence $\{(s_{t},a_{t})\}_{t=0}^{\infty}$ under $f$. Then, the performance of $f$ can be defined by evaluating the state-action sequence with the reward function, and the optimal policies are defined as policies with the highest performance. Optimal policies are defined as stationary policies that maximize the expectation of the cumulative discounted reward under $f\in\mathcal{W}_{A}$ conditioned by an initial state $s\in\mathcal{S}$, which is given by
$V^{f}(s):=\mathbb{E}^{f}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\middle|s_{0}=s\right],$
for all $s\in\mathcal{S}$, where $\mathbb{E}^{f}\left[\cdot\right]$ is the expectation over $\{(s_{t},a_{t})\}_{t=0}^{\infty}$ generated under $(f,p,\rho)$. $V^{f}(s)$ is called a state value function of $f$.
There always exists a deterministic optimal policy Sutton & Barto (2018) such that, for all $s\in\mathcal{S}$,
$f^{*}(s)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\bigg{\{}r(s,a)\\ +\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\max_{f\in\mathcal{W}}V^{f}(s^{\prime})\right]\bigg{\}}.$ (1)
The main algorithms used to determine the optimal policy are value iteration and policy iteration when the transition function and the reward function are known and RL when they are unknown Sutton & Barto (2018).
2.2 Stochastic Game and Stackelberg Game
Stochastic games Shapley (1953); Solan (2022), also called Markov games, are the extension of MDPs for the multi-agent. When there are two agents, A and B, they have finite action spaces $\mathcal{A}$ and $\mathcal{B}$, bounded reward functions $r_{A}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$ and $r_{B}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$, discount rates $\gamma_{A}\in[0,1)$ and $\gamma_{B}\in[0,1)$ and sets of decision rules $\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ and $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\}$, respectively. The state $s\in\mathcal{S}$ is shared among the agents, and the transition function is defined as $p:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta(\mathcal{S})$, that is, the transition probability depends on the actions of all the agents.
The expectation of the cumulative discounted reward under stationary policies $f\in\mathcal{W}_{A}$ and $g\in\mathcal{W}_{B}$ is defined for $r_{A}$ and $r_{B}$, respectively. Then, a pair of policies that are Nash equilibrium or other equilibria can be defined as the solutions to the problem by considering each cumulative reward as the payoff of the agent.
The relationship of the reward functions between agents characterizes the game. In game theory, the game where $r_{A}=-r_{B}$ is called zero-sum game, and the others are called non-zero-sum game. Both are called general-sum games. In the context of a multi-agent RL, situations where $r_{A}\propto-r_{B}$ and $r_{A}\propto r_{B}$ are called competitive and cooperative, respectively.
The Stackelberg game is a game model where a leader first selects its strategy and then a follower selects its strategy such that it maximizes the follower’s payoff against the leader’s strategy. This follower’s strategy is called the best response strategy against the leader’s strategy. Then, when the leader’s strategy maximizes the leader’s payoff under the follower’s best response, the pair of strategies is called a Stackelberg equilibrium. In this study, we consider the extension of Stackelberg games to stochastic games, such as stochastic Stackelberg game (SSG). An example of the application of SSG is provided in Appendix M. Unlike the single-agent MDP, the existence of the optimal policy, the SSE policy defined below, is not guaranteed, and the SSE policy is not deterministic even if it exists.
3 Problem Setting
A general-sum SSG is formulated. Its solutions are defined as the policies that are Stackelberg equilibria.
In a 2-player stochastic game represented by $(\mathcal{S},(\mathcal{A},\mathcal{B}),\hat{p},\rho,(\hat{r}_{A},\hat{r}_{B}),(\gamma_{A},\gamma_{B}))$, agent A is called a “leader” and agent B is called a “follower.” The sets of decision rules are defined by
$\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\},$ $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\},$
and the sets of deterministic decision rules are defined by
$\mathcal{W}_{A}^{d}:=\{f:\mathcal{S}\to\mathcal{A}\},$ $\mathcal{W}_{B}^{d}:=\{g:\mathcal{S}\to\mathcal{B}\}.$
Let $\mathcal{F}_{\mathcal{S}}:=\{v:\mathcal{S}\to\mathbb{R}\}$ be the set of real-valued functions with a state as input.
For conciseness, we define the marginalization of $\hat{r}$ and $\hat{p}$ over the leader’s action $a\in\mathcal{A}$ under an action distribution $f_{s}\in\Delta(\mathcal{A})$ as
$r_{i}(s,f_{s},b)$ $:=\sum_{a\in\mathcal{A}}f_{s}(a)\hat{r}_{i}(s,a,b),\quad(\forall i\in\{A,B\})$
$p(s^{\prime}|s,f_{s},b)$ $:=\sum_{a\in\mathcal{A}}f_{s}(a)\hat{p}(s^{\prime}|s,a,b),$
where $s\in\mathcal{S}$ is a state and $b\in\mathcal{B}$ is a follower’s action. Hereafter, the players are indexed by $i\in\{A,B\}$.
The state value functions of the leader and the follower are defined as follows.
Definition 3.1 (State value function).
Let $s\in\mathcal{S}$ be a state and $f\in\mathcal{W}_{A},g\in\mathcal{W}_{B}$ be stationary policies of the leader and the follower. The state value functions $V_{i}^{fg}\in\mathcal{F}_{\mathcal{S}}$ of player $i\in\{A,B\}$ for a pair $(f,g)$ are given by
$V_{i}^{fg}(s):=\mathbb{E}^{fg}\left[\sum_{t=0}^{\infty}\gamma_{i}^{t}\hat{r}_{i}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right],$
where $\mathbb{E}^{fg}\left[\cdot\right]$ is the expectation over the state-action sequence $\{(s_{t},a_{t},b_{t})\}_{t=0}^{\infty}$ generated under $(f,g,\hat{p},\rho)$.
The follower’s best response is defined by employing the state value functions.
Definition 3.2 (Follower’s best response).
Let $s\in\mathcal{S}$ be a state and $f\in\mathcal{W}_{A}$ a leader’s decision rule. The set of the follower’s best responses $R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$ against $f$ is defined as
$R_{B}^{*}(f):=\{g\in\mathcal{W}_{B}^{d}~{}|~{}g(s)\in R_{B}^{*}(s,f)~{}\forall s\in\mathcal{S}\},$
where
$R_{B}^{*}(s,f):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f(s),b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),b)}\left[\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s^{\prime})\right].$
As the optimal policies in single-agent MDPs are represented as in Equation 1, we can observe that the set of best responses $R_{B}^{*}(f)$ is the set of the optimal policies with respect to single-agent MDPs when the leader’s policy is fixed to the stationary policy $f$. It is also shown that $R_{B}^{*}(f)$ is non-empty for all $f\in\mathcal{W}_{A}$.
The follower that always adopts the best response against the leader’s policy is called the “best-response follower.” If the leader adopts a stationary policy $f\in\mathcal{W}_{A}$, the best-response follower adopts the stationary policy $g\in R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$.
We assume that $R_{B}^{*}(s,f)$ is a singleton for all $s\in\mathcal{S},f\in\mathcal{W}_{A}$. It follows that $R_{B}^{*}(f)$ is also singleton for all $f\in\mathcal{W}_{A}$; thus, the follower’s best response against $f$ is unique. For simplicity, let $R_{B}^{*}(f)$ be the best response. This enables us to define a state value function $V_{A}^{fR_{B}^{*}(f)}$ of $f$ under the best-response follower. For simplicity, we denote
$V_{A}^{f{\dagger}}:=V_{A}^{fR_{B}^{*}(f)}.$
We aim to determine an SSE policy, which is defined as the leader’s stationary policy that maximizes the leader’s state value function for all states under the best-response follower. This policy and the SE value function are defined below.
Definition 3.3 (SE value function).
Let $s\in\mathcal{S}$ be a state. An SE value function is defined as
$V_{A}^{*}(s):=\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$
Definition 3.4 (SSE policy).
If a leader’s stationary policy $f^{*}\in\mathcal{W}_{A}$ satisfies $V_{A}^{f^{*}{\dagger}}(s)=V_{A}^{*}(s)$ for all $s\in\mathcal{S}$, then $f^{*}$ is an SSE policy.
Notation
[…]
4 Related Works
Bucarey et al. (2022) proposed an algorithm for “strong” SSEs in the 2-player general-sum stochastic game. Assuming the uniqueness of the follower’s best response, the problem setting of Bucarey et al. (2022) is the same as ours. However, their algorithm requires strong assumptions to guarantee convergence to the SSEs, and there is little assurance of the performance of the obtained policy when it does not converge to the SSEs or when there is no SSE. This problem is because the fixed point of the operator used in Bucarey et al. (2022) is different from the SE value function. We discuss this in detail in Section 5.1.
A case where the algorithm of Bucarey et al. (2022) converges to the SSE is when the follower is myopic, which maximizes its immediate reward. There are some existing studies in the SSGs under the myopic follower. For example, Zhao et al. (2023) derived the upper bound of the regret of the leader’s payoff in the cooperative tasks, where the reward function is shared between the leader and the follower, with information asymmetry under the myopic follower. Zhong et al. (2023) proposed the guaranteed algorithm in the SSG where the group of myopic followers forms a Nash equilibrium among them. Different from these works, we focus on the situation where the follower is not myopic.
Some studies concentrate on the cooperative setting. Kao et al. (2022) and Zhao et al. (2023) analyzed the cooperative tasks with information asymmetry, where the leader cannot observe the follower’s action Kao et al. (2022) and the reward functions are known only to the follower Zhao et al. (2023). Kononen (2004) and Zhang et al. (2020a) point out Pareto efficiency and uniqueness as advantages of Stackelberg equilibria to Nash equilibria in cooperative tasks.
Zhang et al. (2020a) is similar to ours and Bucarey et al. (2022) in that the problem setting is the general-sum and the follower is not myopic. The differences are that the transition function is unknown (i.e., in the RL setting), the follower’s policy takes the leader’s action as input, and only deterministic stationary policies are considered the leader’s policies. The algorithm proposed by Zhang et al. (2020a) shares its principal foundation with that of Bucarey et al. (2022) (see Appendix L). This means that these methods use the same operator, implying that the limitation of the method of Bucarey et al. (2022) applies to that of Zhang et al. (2020a).
We proposed a novel algorithm for general-sum SSGs where the transition function and the reward functions are known and the follower is not myopic, which is the same as Bucarey et al. (2022). Different from the previous work, we proposed a novel approach with a convergence guarantee even when no SSE policy exists to deal with such a case.
5 Fixed Point Analysis of Operators
5.1 Operator defined by Bucarey et al. (2022)
Bucarey et al. (2022) addressed the same problem as ours under the uniqueness of $R_{B}^{*}$ [^1] Bucarey et al. (2022), which defines Strong SSEs to avoid the ambiguity of the best responses and utilizes it as the optimal solution of the problem. In this study, instead, we assume that there is no ambiguity in the best responses; hence, the SSEs are indeed strong SSEs. . They defined an operator whose fixed point coincides with the SE value function under certain conditions. Then, they proposed a method to construct a stationary equilibrium policy corresponding to the fixed point of the defined operator, which is called fixed point equilibrium (FPE).
However, they required strong assumptions that can potentially limit the applications of their approach (see for example Appendix M). First, they assumed that the follower is “myopic” or the reward functions are either cooperative or competitive to guarantee the convergence of their algorithm to the FPE Bucarey et al. (2022); Zhang et al. (2020a). Second, to guarantee the coincidence of the FPE and the SSE, the myopic follower was assumed Bucarey et al. (2022). Notably, the situation where the follower is “myopic” in Bucarey et al. (2022) is equivalent to the one where the follower’s discount rate $\gamma_{B}=0$ or the one where the transition function is not dependent on the follower’s actions Bucarey et al. (2022). That is, if the follower is not myopic, the algorithm may not converge and the fixed point of the algorithm may be different from the SSE.
Even if the convergence of the algorithm is not guaranteed, it is of great significance from a practical viewpoint to understand the characteristics of the FPE if exists. Particularly, it is interesting to examine how effective the FPE is if it does not coincide with the SSE. Thus, we analyzed the fixed point of the operator defined in Bucarey et al. (2022).
The operator $T:\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ defined by Bucarey et al. (2022) for the strong SSE policies can be simplified as one for the SSE policies as: for $v:=(v_{A},v_{B})\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$, $i\in\{A,B\}$ and $s\in\mathcal{S}$,
$(Tv)_{i}(s):=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))\\ +\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))}\left[v_{i}(s^{\prime})\right],$ (2)
where $R_{A}(s,v)$ is
$R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$
and is assumed to exist for all $s$ and $v$, [^2] This assumption is necessary to derive a policy corresponding to a value function and Bucarey et al. (2022) implicitly assume it. However, it is not guaranteed because the inside of $\operatornamewithlimits{argmax}$ is not necessarily continuous with respect to $f_{s}$, even though $\Delta(\mathcal{A})$ is compact. $R_{B}$ always exists because $\mathcal{B}$ is finite, and
$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right]$
is assumed to be a singleton for all $s$, $f_{s}$, and $v_{B}$. [^3] Bucarey et al. (2022) defines the “strong” version of $R_{B}$ similarly to the strong SSE. Thus, we assume that $R_{B}$ is a singleton to simplify the discussion in accordance with our goal to find the SSE policies. As the core of our analysis does not depend on the uniqueness of $R_{B}$, the result can be easily extended to the method of Bucarey et al. (2022).
If there exists a fixed point $V:=(V_{A},V_{B})$ of $T$, the pair of the value functions of the FPE $(\bar{f},\bar{g})$ defined as follows coincides with $V$ Bucarey et al. (2022): for any $s\in\mathcal{S}$,
$\bar{f}(s)\in R_{A}(s,V),$ $\bar{g}(s)\in R_{B}(s,\bar{f}(s),V_{B}).$
$T$ is a contraction mapping for $v_{A}$ under certain conditions [^4] Bucarey et al. (2022). An example is the case when $r_{A}$ is an affine transformation of $r_{B}$. In this situation, the game is either cooperative or competitive.. Then, we can obtain the fixed point asymptotically by repeatedly applying such $T$ to an arbitrary initial function $v_{A}^{(0)}\in\mathcal{F}_{\mathcal{S}}$ Bucarey et al. (2022). In this case, a single update by $T$ can be interpreted as solving a normal-form Stackelberg game on each state $s\in\mathcal{S}$ by considering the current Q functions
$q_{i}^{(t)}(s,a,b):=r_{i}(s,a,b)+\gamma_{i}\mathbb{E}_{s^{\prime}}\left[v_{i}^{(t)}(s^{\prime})\right]\quad(i\in\{A,B\})$
as the gain tables on $s$ for the leader and the follower. This operation is also used in Nash-Q Learning Hu & Wellman (2003), which is guaranteed to converge to the Nash equilibrium in stochastic games. However, the convergence to the Stackelberg equilibrium by the same operation is not guaranteed even if the equilibrium exists.
Let us consider the characteristics of the fixed point $(V_{A},V_{B})$ of the operator $T$ in general situations where the fixed point exists but it is different from the SE value functions. Theorem 5.1 reveals a general property of $(V_{A},V_{B})$. The proof is in Appendix A.
Theorem 5.1.
For $V:=(V_{A},V_{B})\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$, let $R_{A}(V)\in\mathcal{W}_{A}$ be a stationary policy whose action distribution on a state $s\in\mathcal{S}$ is in $R_{A}(s,V)$ and $R_{B}(f,V_{B})$ be a deterministic stationary policy under $f\in\mathcal{W}_{A}$ whose action on a state $s$ is in $R_{B}(s,f(s),V_{B})$. Then, if $(TV)_{i}(s)=V_{i}(s)$ holds for all $i\in\{A,B\}$ and for all $s\in\mathcal{S}$, it holds that
$V_{A}(s)$ $=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)\quad\forall s\in\mathcal{S};$
$V_{B}(s)$ $=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s)\quad\forall s\in\mathcal{S}.$
If $V$ is the fixed point, because $\bar{f}=R_{A}(V)$ and $\bar{g}=R_{B}(\bar{f},V_{B})$ , it holds that
$V_{B}^{\bar{f}\bar{g}}(s)=V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{\bar{f}g}(s),$ (3)
$V_{A}^{\bar{f}\bar{g}}(s)=V_{A}(s)=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)$ (4)
for all $s\in\mathcal{S}$ in light of Theorem 5.1. Equation 3 shows that the follower’s FPE $\bar{g}=R_{B}(\bar{f},V_{B})$ is the best response to the leader’s FPE $\bar{f}$. However, $R_{B}(f,V_{B})$ is not a best response to an arbitrary $f\in\mathcal{W}_{A}$, except when $V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s)~{}\forall s\in\mathcal{S}$. In addition, determining any meaning from $R_{B}(f,V_{B})$ is difficult. Therefore, $V_{A}^{\bar{f}\bar{g}}$ does not coincides with the SE value function. Moreover, there the performance of the FPE is not guaranteed, which is also empirically demonstrated in Bucarey et al. (2019).
5.2 State-dependent Markov policy and SSE Policy
An attempt to design another operator that has an SE value function as a fixed point is made. From this attempt, we derive a novel sufficient condition for the existence of the SSE policy that is also a necessary condition for a special case. The game class that is guaranteed to have SSE policies by our sufficient condition includes the game class guaranteed by Bucarey et al. (2022).
First, we define a new policy class, state-dependent Markov policy, in which the whole decision rule for each state is determined depending on the state.
Definition 5.2 ((Leader’s) State-dependent Markov policy).
Let $s\in\mathcal{S}$ be a state visited by the agent. A decision rule that the agent follows at $s$ is determined depending on $s$ when the agent follows a state-dependent Markov policy. In other words, let $\hat{\mathbf{f}}:\mathcal{S}\times\mathcal{S}\to\Delta(\mathcal{A})$ be the agent’s state-dependent Markov policy and $\hat{\mathbf{f}}(\cdot;s)\in\mathcal{W}_{A}$ be the agent’s decision rule at $s$. Then, the agent’s action on $s$ is sampled as $a\sim\hat{\mathbf{f}}(s;s)$ and the follower’s best response action is $R_{B}^{*}(s,\hat{\mathbf{f}}(\cdot;s))\in\mathcal{B}$.
Let $\mathcal{W}_{A}^{\mathcal{S}},\mathcal{W}_{B}^{\mathcal{S}}$ be the sets of state-dependent Markov policies for the leader and the follower, respectively. For simplicity, let $\hat{f}_{s}:=\hat{\mathbf{f}}(\cdot;s)$.
State-dependent Markov policies are an extension of the stationary policy concept, such that $\mathcal{W}_{A}^{\mathcal{S}}\supset\mathcal{W}_{A}$. Similar to the stationary policies, we can define the leader’s state value function under the best-response follower as
$V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$ $:=V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s)$
$:=\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right],$
where $R_{B}^{*}(\hat{\mathbf{f}})\in\mathcal{W}_{B}^{\mathcal{S}}$ is the follower’s state-dependent Markov policy in which the decision rule on the visited state $s\in\mathcal{S}$ is $R_{B}^{*}(\hat{f}_{s})\in\mathcal{W}_{B}$, and $\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}[\cdot]$ is an expectation for the trajectory $\{(s_{t},a_{t},b_{t})\}_{t=0}^{\infty}$, which is generated under $(\hat{\mathbf{f}},R_{B}^{*}(\hat{\mathbf{f}}),\hat{p},\rho)$.
Let the maximization of $V_{A}^{\hat{\mathbf{f}}{\dagger}}$ on $\mathcal{W}_{A}^{\mathcal{S}}$ be defined as
$\hat{V}_{A}^{*}(s):=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$
for each $s\in\mathcal{S}$. Subsequently, Theorem 5.3 shows that $\hat{V}^{*}$ is the unique fixed point of the operator $\hat{T}_{A}$ defined in Equation 5. The proof is in Appendix B.
Theorem 5.3.
Let $\hat{T}_{A}:\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}$ be defined as
$(\hat{T}_{A}v)(s):=\sup_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\quad\forall s\in\mathcal{S}.$ (5)
Then, $\hat{T}_{A}$ is a contraction mapping and its unique fixed point is $\hat{V}_{A}^{*}$.
We define the set of decision rules $R_{A}^{*}(s)$ for each state $s\in\mathcal{S}$ as
$R_{A}^{*}(s):=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (6)
and we suppose $R_{A}^{*}(s)$ is non-empty for all $s\in\mathcal{S}$. Lemma 5.4 provides a way to construct a state-dependent Markov policy by using $R_{A}^{*}$ that realizes $\hat{V}_{A}^{*}$. This proof is in Appendix C.
Lemma 5.4.
Let $\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}$ be a leader’s state-dependent Markov policy that satisfies $\hat{\mathbf{f}}(\cdot;s)\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. Then, $V_{A}^{\hat{\mathbf{f}}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}\hat{V}_{A}^{*}$ holds.
Finally, Theorem 5.5 provides a sufficient condition for the existence of SSE policies. The proof is in Appendix D.
Theorem 5.5.
Let $R_{A}^{*}(s)$ be a subset of $\mathcal{W}_{A}$ defined in Equation 6 for each $s\in\mathcal{S}$. Then, if there exists $\hat{f}\in\mathcal{W}_{A}$ such that $\hat{f}\in\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$, $\hat{f}$ is an SSE policy.
If we assume the sufficient conditions for the existence of SSE policies in Bucarey et al. (2022), $\gamma_{B}=0$ holds, or the transition does not depend on the follower’s action. Then, for $f\in\mathcal{W}_{A}$, because $R_{B}^{*}(s,f)$ depends only on $s\in\mathcal{S}$ and the leader’s action distribution $f(s)\in\Delta(\mathcal{A})$ on state $s$, the part to the right of the $\max$ operator in the right-hand side of Equation 6 depends only on $f(s)\in\Delta(\mathcal{A})$ and not on all $f(s^{\prime})\in\Delta(\mathcal{A})$ such that $s^{\prime}\neq s$. Thus, $R_{A}^{*}(s)$ is a restricted set with respect only to the action distribution $f(s)\in\Delta(\mathcal{A})$, which shows that there exists $f^{\prime}\in\mathcal{W}_{A}$ such that $f^{\prime}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. Then, the existence of SSE policies is guaranteed by Theorem 5.5. This means that the game class whose existence of SSE policies is guaranteed by Theorem 5.5 includes the game class whose existence is guaranteed by Bucarey et al. (2022).
Rather than considering a myopic follower, if we consider the situation that the leader’s discount rate $\gamma_{A}=0$, the sufficient condition in Theorem 5.5 becomes a necessary condition. Its proof is provided in Appendix E.
Corollary 5.6.
If $\gamma_{A}=0$, then an SSE policy exists if and only if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is non-empty.
Even if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is empty, we can construct stationary policies utilizing $R_{A}^{*}$, which is asymptotically obtained using $\hat{T}_{A}$ in light of Theorem 5.3. A possible approach is to choose $f$ contained in $R_{A}^{*}(s)$ for most of the states or to extract one non-dominated by stationary policies in $R_{A}^{*}$. However, the performance is not guaranteed for the stationary policies obtained as aforementioned. This is because any stationary policy not included in $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ does not achieve $\hat{V}_{A}^{*}$, and its performance may be significantly worse than $\hat{V}_{A}^{*}$ as $R_{B}^{*}(f)$; hence, $\hat{V}_{A}^{*}$ is not necessarily continuous with respect to $f$.
6 Policy Iteration for Pareto-Optimal Policies
SSE policies do not always exist in general-sum stochastic games. Therefore, it is desired that an algorithm converges to an SSE policy if it exists, and to a policy that is “reasonable” provided an SSE policy does not exist. First, we introduce the notion of Pareto-optimality as the reasonable target. A policy that approximates a Pareto-optimal value function with an arbitrary precision always exists and it admits an SSE policy if an SSE policy exists. Then, we derive a policy improvement theorem for SSGs. Based on the policy improvement theorem for SSGs, we design an algorithm that monotonically improves the policy toward Pareto-optimal policies.
6.1 Pareto-Optimal Policy
SSE policies exist if and only if there exists $f\in\mathcal{W}_{A}$ that maximizes $V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$. We consider $V_{A}^{f{\dagger}}(s)$ as $|\mathcal{S}|$ objective functions with $f\in\mathcal{W}_{A}$ as a common input. From this multi-objective optimization viewpoint, we define the Pareto-optimal policies. Hereafter, we view a value function $v$ as a vector in $\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and apply topological argument in the standard sense in $\mathbb{R}^{\lvert\mathcal{S}\rvert}$.
Definition 6.1 (Pareto Optimality).
Let $\mathcal{V}=\{v\in\mathcal{F}_{\mathcal{S}}:v\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}\ \exists f\in\mathcal{W}_{A}\}$ be the set of a reachable value function. Let $\partial\mathcal{V}$ be the boundary of $\mathcal{V}$ and $\mathrm{cl}\mathcal{V}=\mathcal{V}\cup\partial\mathcal{V}$ be the closure of $\mathcal{V}$. A Pareto-optimal (PO) value function $v^{*}\in\mathrm{cl}\mathcal{V}$ is such that there exists no $v\in\mathrm{cl}\mathcal{V}$ satisfying $v\succ v^{*}$. The set of PO value functions is denoted as $\mathcal{PV}$. A stationary policy $f\in\mathcal{W}_{A}$ satisfying $V_{A}^{f{\dagger}}\in\mathcal{PV}$ is a PO policy.
There always exists $\mathcal{PV}\neq\emptyset$. Moreover, $\mathcal{PV}\subseteq\partial\mathcal{V}$. A PO policy exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$. For any PO value function $v^{*}\in\mathcal{PV}$, there always exists a policy whose value function is arbitrarily close to $v^{*}$. Meanwhile, there is no stationary policy that realizes a value function better than a PO value function. Moreover, an arbitrary PO policy is an SSE policy if there exist SSE policies, as shown in Proposition 6.2. Therefore, PO value functions are a reasonable alternative to the SE value function. The proof is presented in Appendix F.
Proposition 6.2.
The SE value function $V_{A}^{*}\in\mathcal{PV}$ if and only if $\mathcal{PV}$ is a singleton. Consequently, the SSE policies exist if and only if $\mathcal{PV}$ is a singleton and $\mathcal{PV}\subseteq\mathcal{V}$.
6.2 Policy Improvement Theorem
We derive three important lemmas based on which we design the proposed approach.
Lemma 6.3 is an extension of the policy improvement theorem Sutton & Barto (2018) in single-agent MDPs. With this theorem, we can confirm if a policy $f$ is better than a baseline $f^{\prime}$ without computing the value function for $f$. The proof is shown in Appendix G.
Lemma 6.3.
Let $f\in\mathcal{W}_{A},f^{\prime}\in\mathcal{W}_{A}$ be stationary policies, and a Q-function $Q_{A}^{f^{\prime}{\dagger}}(s,f):\mathcal{S}\times\mathcal{W}_{A}\to\mathbb{R}$ is expressed as
$Q_{A}^{f^{\prime}{\dagger}}(s,f):=r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right],$
where $\mathbb{E}_{s^{\prime}}$ is considered under $p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))$. Then, the following conditions hold:
  • (a)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\succcurlyeq Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}$;
  • (b)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}$;
  • (c)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}\iff\\ V_{A}^{f{\dagger}}-V_{A}^{f^{\prime}{\dagger}}(s)\succ\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$.
A necessary condition for PO policies is provided below. It shows that any improvement by Lemma 6.3 (a) is not possible from PO policies. This proof is in Appendix H.
Lemma 6.4 (Necessary condition for PO policies).
If $f^{\prime}\in\mathcal{W}_{A}$ is a PO policy, it holds that for any $f\in\mathcal{W}_{A}$
$Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}.$ (7)
The following result shows a sufficient condition for PO policies. This proof is presented in Appendix I.
Lemma 6.5 (Sufficient condition for PO policies).
$f^{\prime}\in\mathcal{W}_{A}$ is a PO policy if, for any $f\in\mathcal{W}_{A}$, either of the following two conditions hold:
  • i)
    $Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$, $\exists s\in\mathcal{S}$, where $\delta(f,f^{\prime})=\max\limits_{s\in\mathcal{S}}\mathbb{E}_{s^{\prime}}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ and the expectation is considered for $s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))$;
  • ii)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}$.
6.3 Policy Iteration for PO Policies
We propose an approach that monotonically improves a policy toward PO policies and is guaranteed to converge.
PO value functions are not always unique. To guide the algorithm, we introduce scalarization. Given a stationary policy $f\in\mathcal{W}_{A}$, we maximize a Pareto-compliant scalarization $\mathcal{L}$ of the state value on $f$, namely, $\mathcal{L}[V_{A}^{f{\dagger}}]$, where the Pareto-compliant scalarization555 For example, a weighted sum scalarization $\mathcal{L}[v]=\sum_{s\in\mathcal{S}}\alpha_{s}v(s)$ is a Pareto-compliant scalarization for $\boldsymbol{\alpha}:=\{\alpha_{s}\in\mathbb{R}_{>0}\}_{s\in\mathcal{S}}$ such that $\sum_{s\in\mathcal{S}}\alpha_{s}=1$. If we choose $\alpha$ such that $\alpha_{s}=\rho(s)$ for all $s\in\mathcal{S}$, then $\mathcal{L}[v]$ coincides with the expected return under the initial state distribution $\rho$. $\mathcal{L}:\mathrm{cl}\mathcal{V}\to\mathbb{R}$ is a continuous function such that $v\succ v^{\prime}\implies\mathcal{L}[v]>\mathcal{L}[v^{\prime}]$. Then, the maximum of the scalarized value in $\mathrm{cl}\mathcal{V}$ is obtained by a PO value function, namely, $\operatornamewithlimits{argmax}_{v\in\mathrm{cl}\mathcal{V}}\mathcal{L}\left[v\right]\subseteq\mathcal{PV}$.
Directly maximizing $\mathcal{L}[V_{A}^{f{\dagger}}]$ is, however, intractable. This is because the computation of $V_{A}^{f{\dagger}}$ for each $f\in\mathcal{W}_{A}$ requires repeatedly applying the Bellman expectation operator under $f$ and $R_{B}^{*}(f)$,
$(T_{A}^{fR_{B}^{*}(f)}V)(s):=r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[V(s^{\prime})\right],$
to an initial function $V_{0}\in\mathcal{F}_{\mathcal{S}}$ until it converges, where the convergence is guaranteed as in a single-agent MDP.
To alleviate this difficulty and obtain a sequence $\{f_{t}\}_{t\geq 0}^{\infty}$ satisfying $\mathcal{L}[V_{A}^{f_{t+1}{\dagger}}]\geq\mathcal{L}[V_{A}^{f_{t}{\dagger}}]$, we employ the policy improvement theorem, as in the policy iteration in single-agent MDPs. Rather than computing $V_{A}^{f{\dagger}}$ for each $f$, we calculate $Q_{A}^{f_{t}{\dagger}}(\cdot,f)$, requiring only $V_{A}^{f_{t}{\dagger}}$. Then, we construct the set $\mathcal{W}_{\succcurlyeq}(f_{t})$ of policies satisfying the policy improvement condition from $f_{t}$, namely,
$\mathcal{W}_{\succcurlyeq}(f_{t}):=\left\{f\in\mathcal{W}_{A}~{}|~{}Q_{A}^{f_{t}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f_{t}{\dagger}}\right\}.$ (8)
Selecting the next policy $f_{t+1}$ from $\mathcal{W}_{\succcurlyeq}(f_{t})$ guarantees $\mathcal{L}[V_{A}^{f_{t+1}{\dagger}}]\geq\mathcal{L}[V_{A}^{f_{t}{\dagger}}]$ in light of Lemma 6.3 (a). Ideally,666 In practice, $\operatornamewithlimits{argmax}$ is not necessary, and sometimes it does not exist or is not computable. An alternative approach is to select a policy $f_{t+1}$ from a set $\mathcal{W}_{1-\epsilon}(f_{t})$ of policies that are $\epsilon$-optimal in terms of $\mathcal{L}$, such as $\mathcal{W}_{1-\epsilon}(f_{t}):=\bigg{\{}f\in\mathcal{W}_{\succcurlyeq}(f_{t})\mid\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\\ \geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\bigg{(}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\bigg{)}\bigg{\}}$ (9) for some $\epsilon\in[0,1)$. Notably, $\mathcal{W}(f_{t})=\mathcal{W}_{1}(f_{t})$. Therefore, the update in Equation 10 is a special case of Equation 9. If it is still intractable, selecting $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$ s.t. $Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\succ V_{A}^{f_{t}{\dagger}}$ ensures that (a) and (b) in Theorem 6.6 hold. the proposed algorithm selects $f_{t+1}$ from the set of policies that maximizes the scalarized value, namely,
$f_{t+1}\in\mathcal{W}(f_{t}):=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right].$ (10)
Theorem 6.6 guarantees that the proposed algorithm always converges and that the limit satisfies the necessary condition for PO policies (i.e., $v_{\infty}\in\partial\mathcal{V}$). The proof is shown in Appendix J.
Theorem 6.6.
Let $f_{0}\in\mathcal{W}_{A}$ be an arbitrary initial policy and a policy sequence $\{f_{t}\in\mathcal{W}(f_{t-1})\}_{t=1}^{\infty}$ be obtained by Equation 9. Let $\{v_{t}=V_{A}^{f_{t}{\dagger}}\}_{t=0}^{\infty}$ be the corresponding value functions. Then, the following statements hold:
  • a)
    $\{v_{t}\}_{t=0}^{\infty}$ monotonically increases in the sense that $v_{t+1}\succcurlyeq v_{t}$ and converges to $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$;
  • b)
    $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$ if and only if $v_{t}\stackrel{{\scriptstyle s}}{{=}}v_{\infty}$;
  • c)
    $v_{\infty}\in\partial\mathcal{V}$ and for all $v\in\mathcal{V}$,
    $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$;
  • d)
    $v_{\infty}\in\mathcal{PV}$ if $\gamma_{A}=0$.
As shown in the proof (Appendix J), the algorithm ceases to improve if and only if $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. The algorithm terminates when it holds and returns $f_{t}$. The whole proposed algorithm summarizing the above is shown in Algorithm 1.
Algorithm 1 Policy Iteration for PO Policies (PIPOP)
0:  Maximum number of iterations $M$.
1:  Randomly initialize $f_{0}\in\mathcal{W}_{A}$.
2:  for $t=0$ to $M-1$ do
3:     Compute $V_{A}^{f_{t}{\dagger}}$ by repeatedly applying $T_{A}^{f_{t}R_{B}^{*}(f_{t})}$
4:     Compute $\mathcal{W}_{\succcurlyeq}(f_{t})$ in Equation 8
5:     if $Q_{A}^{f_{t}{\dagger}}(s,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$ then
6:        return $f^{*}\leftarrow f_{t}$
7:     else
8:        Compute $\mathcal{W}(f_{t})$ in Equation 10 or Equation 9
9:        Randomly sample $f_{t+1}$ from $\mathcal{W}(f_{t})$
10:     end if
11:  end for
12:  return $f^{*}\leftarrow f_{M}$
12:  A stationary policy $f^{*}$
The convergence guarantee and the monotone improvement of Algorithm 1 (Statements (a) and (b) in Theorem 6.6) show the advantages of the proposed approach. However, compared with the policy iteration approach in single-agent MDPs, where the convergence toward the optimal policy is guaranteed, the fixed point of Algorithm 1 is not assured to be PO value functions. Rather, Theorem 6.6 (c) guarantees that the limit is at the boundary $\partial\mathcal{V}$, including all the PO value functions. In a special case of $\gamma_{A}=0$, the limit is guaranteed to be a PO value function (Theorem 6.6 (d)).
Furthermore, there exists an algorithmic difference to the policy iteration for the single-agent MDP. In single-agent MDPs, a state-wise maximization of the action value function of $f_{t}$, $a_{s}\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}Q^{f_{t}}(s,a)$ is performed to construct the next (deterministic) policy. This is invalid in SSGs because such a policy does not necessarily satisfy the condition of Lemma 6.3 (a) and may not improve the values. Moreover, the optimal policy is not necessarily a deterministic one. Because of these differences, Equation 10 cannot be simplified as the state-wise maximization. This can be a limitation from a practical viewpoint when Equation 10 (or Equation 9) is intractable.
In practice, it is helpful to restart the policy iteration process from other initial policies to obtain a better policy unless the obtained policy satisfies the sufficient condition Lemma 6.5. In this situation, one can randomly pick an initial policy or select a policy $f\in\mathcal{W}(f_{t-1})$ that is not yet explored. Because $f\in\mathcal{W}(f_{t-1})$ is already improved from the previous initial policy, we can expect such an initialization to be more efficient. The detail of the proposed approach with the restart strategy is provided in Appendix K.
7 Limitation and Future Work
This study proposed a novel algorithm with a convergence guarantee in general-sum SSGs. We introduced the notion of the Pareto-optimal value function and developed an algorithm that monotonically improves a policy toward the Pareto-optimal value functions. While existing methods have little guarantee, especially in games where the SSEs do not exist, our proposed approach can monotonically improve the leader’s performance in general, which is of great importance in practice. A potential application of the proposed approach is the intervention to an adaptive or learning follower agent as described in Appendix M, where the follower is not usually myopic; hence the existence of the SSEs is not guaranteed. On such an application, the proposed approach is advantageous to the existing approaches discussed in this paper.
However, there is room for improvement both from theoretical and practical viewpoints. Further research must focus on addressing the limitations of the current work listed below.
First, the Pareto-optimality of the proposed algorithm is not guaranteed. A necessary and sufficient condition for the Pareto-optimality of the algorithm must be derived. The development of a restart strategy to satisfy such a sufficient condition is desired.
Second, a computationally efficient and algorithmic update rule to select the next policy (cf., Equation 10 or Equation 9) is required. Practically, such an update similar to the policy iteration in single-agent MDPs, is desired.
Finally, our approach requires the knowledge of the follower’s best response and transition probability as in the previous studies. A sample approximation of the algorithm by reinforcement learning is desired to widen the applicability of the proposed approach.
Appendix A Proof of Theorem 5.1
Proof.
First, letting $V_{B}$ be fixed, we prove that $(TV)_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}$ implies $V_{A}\stackrel{{\scriptstyle s}}{{=}}\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
Suppose that $V_{A}\stackrel{{\scriptstyle s}}{{=}}(TV)_{A}$. Given a policy $f\in\mathcal{W}_{A}$, let $f_{s}=f(s)$ for each $s\in\mathcal{S}$. Then, we obtain for all $s\in\mathcal{S}$
$V_{A}(s)$ $=(TV)_{A}(s)$ (11a)
$=\sup_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},V_{B}))}\left[V_{A}(s^{\prime})\right]$ (11b)
$\geq r_{A}(s,f_{s},R_{B}(s,f_{s},V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},V_{B}))}\left[V_{A}(s^{\prime})\right]$ (11c)
$=r_{A}(s,f(s),R_{B}(s,f(s),V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}(s,f(s),V_{B}))}\left[V_{A}(s^{\prime})\right].$ (11d)
By repeatedly applying this inequality, we obtain
$V_{A}(s)$ $\geq r_{A}(s,f(s),R_{B}(s,f(s),V_{B}))+\gamma_{A}\mathbb{E}_{s_{1}}\left[V_{A}(s_{1})\right]$ (12a)
=$\end{multlined}r_{A}(s,f(s),R_{B}(s,f(s),V_{B}))+\gamma_{A}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},f(s_{1}),R_{B}(s_{1},f(s_{1}),V_{B}))+\gamma_{A}\mathbb{E}_{s_{2}}\left[V_{A}(s_{2})\right]\Bigl{]}$ (12c)
[…]
$=\mathbb{E}^{fR_{B}(f,V_{B})}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}(f,V_{B})}\Bigl{[}V_{A}(s_{n})\Bigl{|}s_{0}=s\Bigl{]}.$ (12d)
In the limit for $n\to\infty$, the first term of the right-hand side of Equation 12d converges to $V_{A}^{fR_{B}(f,V_{B})}(s)$, while the second term converges to 0777Since $\hat{r}_{A}$ is bounded and $\gamma_{A}\in[0,1)$, the inside of the expectation of Equation 12d is always bounded, which enables the exchange of the extreme and the expectation. Therefore, the same exchanges are performed in Equation 16d and Equation 37e.. As the choice of $f$ is arbitrary, we obtain $V_{A}\succcurlyeq\sup_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
We show $V_{A}\stackrel{{\scriptstyle s}}{{=}}\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$. It suffices to show that there exists a policy $f^{*}\in\mathcal{W}_{A}$ such that $V_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{*}R_{B}(f^{*},V_{B})}$. Because $R_{A}(s,V)$ is non-empty, let $f_{s}^{*}\in R_{A}(s,V)$ for each $s\in\mathcal{S}$ and $f^{*}$ be a policy whose action distribution at $s$ is $f^{*}(s)=f_{s}^{*}$ for each $s$, we have for all $s\in\mathcal{S}$
$V_{A}(s)$ $=(TV)_{A}(s)$ (13a)
$=\sup_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},V_{B}))}\left[V_{A}(s^{\prime})\right]$ (13b)
$=r_{A}(s,f_{s}^{*},R_{B}(s,f_{s}^{*},V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s}^{*},R_{B}(s,f_{s}^{*},V_{B}))}\left[V_{A}(s^{\prime})\right]$ (13c)
$=r_{A}(s,f^{*}(s),R_{B}(s,f^{*}(s),V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f^{*}(s),R_{B}(s,f^{*}(s),V_{B}))}\left[V_{A}(s^{\prime})\right].$ (13d)
Therefore, with the repeated application of this equality, analogously to the above argument, we obtain $V_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{*}R_{B}(f^{*},V_{B})}$.
Second, letting $V_{A}$ be fixed, we prove that $(TV)_{B}\stackrel{{\scriptstyle s}}{{=}}V_{B}$ implies $V_{B}\stackrel{{\scriptstyle s}}{{=}}\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}$ for $f=R_{A}(V)$.
Suppose $V_{B}\stackrel{{\scriptstyle s}}{{=}}(TV)_{B}$ holds. Then, by constructing $R_{B}$, we obtain
$V_{B}(s)=(TV)_{B}(s)\geq r_{B}(s,R_{A}(s,V),b_{s})+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),b_{s})}\left[V_{B}(s^{\prime})\right]$ (14)
for all $s\in\mathcal{S}$ and for all $b_{s}\in\mathcal{B}$. It implies that
$V_{B}(s)\geq r_{B}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),g(s))}\left[V_{B}(s^{\prime})\right]$ (15)
for all $s\in\mathcal{S}$ and for all $g\in\mathcal{W}_{B}^{d}$. By recursively applying Equation 15 for $V_{B}$ on the right-hand side, we have for all $s\in\mathcal{S}$
$V_{B}(s)$ $\geq r_{A}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s_{1}}\left[V_{B}(s_{1})\right]$ (16a)
$\end{multlined}r_{A}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},R_{A}(s_{1},V),g(s_{1}))+\gamma_{B}\mathbb{E}_{s_{2}}\left[V_{B}(s_{2})\right]\Bigl{]}$ (16c)
[…]
$\geq\mathbb{E}^{R_{A}(V)g}\left[\sum_{t=0}^{n-1}\gamma_{B}^{t}\hat{r}_{B}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]+\gamma_{B}^{n}\mathbb{E}^{R_{A}(V)g}\Bigl{[}V_{B}(s_{n})\Bigl{|}s_{0}=s\Bigl{]}.$ (16d)
In the limit for $n\to\infty$, the first term of the right-hand side of Equation 16d converges to $V_{B}^{R_{A}(V)g}(s)$, and the second term converges to 0. Therefore, $V_{B}(s)\geq V_{B}^{R_{A}(V)g}(s)$ holds for all $s\in\mathcal{S}$. Since $g\in\mathcal{W}_{B}^{d}$ is arbitrary, it holds that
$V_{B}(s)\succcurlyeq\max_{g\in\mathcal{W}_{B}^{d}}V_{B}^{R_{A}(V)g}.$ (17)
Owing to the existence of a deterministic optimal policy for any single-agent MDP, we obtain
$V_{B}\succcurlyeq\max_{g\in\mathcal{W}_{B}^{d}}V_{B}^{R_{A}(V)g}\stackrel{{\scriptstyle s}}{{=}}\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}.$ (18)
We show that $V_{B}(s)\stackrel{{\scriptstyle s}}{{=}}\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}$. It suffices to show that there exists $g^{*}$ such that $V_{B}(s)\stackrel{{\scriptstyle s}}{{=}}V_{B}^{R_{A}(V)g^{*}}$. Let $g^{*}$ such that $g^{*}(s)=R_{B}(s,R_{A}(s,V),V_{B})$ for each $s\in\mathcal{S}$. Then, by the definition of $R_{B}$, for any $s\in\mathcal{S}$, we have
$V_{B}(s)=r_{B}(s,R_{A}(s,V),g^{*}(s))+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),g^{*}(s))}\left[V_{B}(s^{\prime})\right].$ (19)
With the repeated application of this equality, analogous to the above argument, we obtain $V_{B}\stackrel{{\scriptstyle s}}{{=}}V_{B}^{R_{A}(V)g^{*}}$. ∎
Appendix B Proof of Theorem 5.3
Proof.
First, we prove that $\hat{T}_{A}$ is a contraction mapping; hence it has a unique fixed point. For an arbitrary $(v,u)\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ and $s\in\mathcal{S}$, we obtain
$(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)=\sup_{f\in\mathcal{W}_{A}}\left\{r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\right\}\\ -\sup_{f\in\mathcal{W}_{A}}\left\{r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[u(s^{\prime})\right]\right\}$ (20)
from the definition. Let the first term on the right-hand side of the above equation be $v^{*}(s)$. Then, for an arbitrary $\delta>0$, there exists $f\in\mathcal{W}_{A}$ such that
$r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\geq v^{*}(s)-\delta.$ (21)
Let $\delta=\frac{1-\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}$ and $f_{v}$ satisfy the above condition. Then,
$(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)$ $\leq v^{*}(s)-\left\{r_{A}(s,f_{v}(s),R_{B}^{*}(s,f_{v}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{v}(s),R_{B}^{*}(s,f_{v}))}\left[u(s^{\prime})\right]\right\}$ (22a)
[…]
$=\delta+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{v}(s),R_{B}^{*}(s,f_{v}))}\left[v(s^{\prime})-u(s^{\prime})\right]$ (22e)
$\leq\delta+\gamma_{A}\lVert v(s)-u(s)\rVert_{\infty}$ (22f)
$=\frac{1+\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}.$ (22g)
Analogously, we have
$(\hat{T}_{A}u)(s)-(\hat{T}_{A}v)(s)\leq\frac{1+\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}.$ (23)
Therefore, since $s$ is arbitrary, we have
$\lVert(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)\rVert_{\infty}\leq\frac{1-\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}$ (24)
for any $(v,u)\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$. This implies that $\hat{T}_{A}$ is a contraction mapping. According to the Banach fixed-point theorem, $\hat{T}_{A}$ has a unique fixed point.
Subsequently, we prove that $\hat{V}_{A}^{*}$ is the fixed point of $\hat{T}_{A}$. From the definition, we obtain
$\hat{V}_{A}^{*}(s)$ $=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$ (25a)
$=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[V_{A}^{\hat{\mathbf{f}}{\dagger}}(s^{\prime})\right]$ (25b)
$\leq\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (25c)
$=\sup_{\hat{f}_{s}\in\mathcal{W}_{A}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (25d)
$=(\hat{T}_{A}\hat{V}_{A}^{*})(s)$ (25e)
for all $s\in\mathcal{S}$. We show that the equality in Equation 25c holds for all $s\in\mathcal{S}$. Suppose that there exists $s_{0}\in\mathcal{S}$ such that the above inequality strictly holds. Let the discrepancy between the left-hand side and the right-hand side of the above inequality at state $s_{0}$ be $2\epsilon$. Then, there must exists $\hat{\mathbf{f}}\in\mathcal{W}_{\mathcal{A}}^{\mathcal{S}}$ such that
$\hat{V}_{A}^{*}(s)$ $\leq r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]-\epsilon\quad\text{for }s=s_{0}$ (26)
$\hat{V}_{A}^{*}(s)$ $\leq r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]+\epsilon(1-\gamma_{A})\quad\text{for }s\neq s_{0}$ (27)
holds, where $\hat{f}_{s}$ is the policy derived from $\hat{\mathbf{f}}$ at $s$. Then, starting $s_{0}$ with Equation 26 and repeatedly applying Equation 27, we obtain
$\hat{V}_{A}^{*}(s_{0})$
$\leq r_{A}(s_{0},\hat{f}_{s_{0}}(s_{0}),R_{B}(s_{0},\hat{f}_{s_{0}})+\gamma_{A}\mathbb{E}_{s_{1}}\left[\hat{V}_{A}^{*}(s_{1})\right]-\epsilon$ (28a)
$\end{multlined}r_{A}(s_{0},\hat{f}_{s_{0}}(s_{0}),R_{B}(s_{0},\hat{f}_{s_{0}}))+\gamma_{A}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},\hat{f}_{s_{1}}(s_{1}),R_{B}(s_{1},\hat{f}_{s_{1}}))+\gamma_{A}\mathbb{E}_{s_{2}}\left[\hat{V}_{A}^{*}(s_{2})\right]\Bigl{]}-\epsilon+\epsilon\gamma_{A}(1-\gamma_{A})$ (28c)
[…]
$\leq\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]-\epsilon+\epsilon(1-\gamma_{A})\sum_{t=0}^{\infty}\gamma_{A}^{t+1}$ (28d)
$=V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)-\epsilon(1-\gamma_{A}).$ (28e)
That is, $V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)>\hat{V}_{A}^{*}(s_{0})$, contradicting $\hat{V}_{A}^{*}\stackrel{{\scriptstyle s}}{{=}}\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}$. Therefore, the equality must hold in Equation 25c for all $s\in\mathcal{S}$. This implies
$\hat{V}_{A}^{*}(s)=\max_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]=(\hat{T}_{A}\hat{V}_{A}^{*})(s).$ (29)
This completes the proof. ∎
Appendix C Proof of Lemma 5.4
Proof.
Let $\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}$ satisfy $\hat{f}_{s}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. When the follower follows the best response against $\hat{\mathbf{f}}$, from the definition of $R_{A}^{*}$, we have
$r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]\\ =\max_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (30)
for all $s\in\mathcal{S}$. Therefore, from Equation 29, it holds that
$\hat{V}_{A}^{*}(s)=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (31)
for all $s\in\mathcal{S}$. Let an operator $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}:\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}$ be defined as
$(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}v)(s)=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[v(s^{\prime})\right],$ (32)
then $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$ is a contraction mapping, and as it holds that
$V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s)$ $=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s^{\prime})\right]$ (33a)
$=(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})})(s)$ (33b)
for all $s\in\mathcal{S}$ from the definition of $V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$, $V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$ is the unique fixed point of $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$.
From Equation 31, since $\hat{V}_{A}^{*}$ satisfies $\hat{V}_{A}^{*}(s)=(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\hat{V}_{A}^{*})(s)$ for all $s\in\mathcal{S}$, it holds that $V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)=\hat{V}_{A}^{*}(s)$ for all $s\in\mathcal{S}$. ∎
Appendix D Proof of Theorem 5.5
Proof.
From Lemma 5.4, it follows that if there exists $f^{\prime}\in\mathcal{W}_{A}$ such that $\hat{f}\in\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$, it holds that $V_{A}^{\hat{f}{\dagger}}(s)=\hat{V}_{A}^{*}(s)$ for all $s\in\mathcal{S}$. Then, because it holds that
$\hat{V}_{A}^{*}(s)=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)\geq\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ (34)
from $\mathcal{W}_{A}^{\mathcal{S}}\supset\mathcal{W}_{A}$, it holds that
$\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)\geq V_{A}^{\hat{f}{\dagger}}(s)=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)\geq\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ (35)
for all $s\in\mathcal{S}$. Therefore, it holds that $V_{A}^{\hat{f}{\dagger}}(s)=\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$. ∎
Appendix E Proof of Corollary 5.6
Proof.
From Theorem 5.5, if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is non-empty, then a SSE policy exists.
If $\hat{f}\in\mathcal{W}_{A}$ is a SSE policy, it is holds that $\hat{f}\in\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$ from the definition. If $\gamma_{A}=0$, because
$R_{A}^{*}(s)=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))$
holds, $\hat{f}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. It implies that $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is nonempty. ∎
Appendix F Proof of Proposition 6.2
Proof.
First, suppose that $V_{A}^{*}\in\mathcal{PV}$. Then, by the definition of the SE value function, we have $V_{A}^{*}\succcurlyeq v$ for all $v\in\mathrm{cl}\mathcal{V}$. Therefore, no other PO value function exists. Hence, $\mathcal{PV}=\{V_{A}^{*}\}$.
[…]
Altogether, we have $V_{A}^{*}\in\mathcal{PV}$ if and only if $\mathcal{PV}$ is a singleton. Because PO policies exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$, the SE policy exists if and only if $\mathcal{PV}=\{v^{*}\}$ is a singleton and $v^{*}\in\mathcal{V}$. ∎
Appendix G Proof of Lemma 6.3
Proof.
We prove each statement one by one.
Proof of (a)
The definition of $Q_{A}^{f^{\prime}{\dagger}}(s,f)$ is equivalent to
$Q_{A}^{f^{\prime}{\dagger}}(s,f)=\mathbb{E}^{fR_{B}^{*}(f)}\left[\hat{r}_{A}(s,a,b)\middle|s\right]+\gamma_{A}\mathbb{E}^{fR_{B}^{*}(f)}\left[V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\middle|s\right].$ (36)
Then, if it holds that $V_{A}^{f^{\prime}{\dagger}}(s)\leq Q_{A}^{f^{\prime}{\dagger}}(s,f)$ for all $s\in\mathcal{S}$, we have, for all $s_{0}\in\mathcal{S}$,
$V_{A}^{f^{\prime}{\dagger}}(s_{0})$ $\leq\mathbb{E}^{fR_{B}^{*}(f)}\left[\hat{r}_{A}(s_{0},a_{0},b_{0})\middle|s_{0}\right]+\gamma_{A}\mathbb{E}^{fR_{B}^{*}(f)}\left[V_{A}^{f^{\prime}{\dagger}}(s_{1})\middle|s_{0}\right]$ (37a)
[…]
$\leq\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}^{*}(f)}\left[V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}\right].$ (37e)
In the limit for $n\to\infty$, the first term converges to $V_{A}^{fR_{B}^{*}(f)}(s_{0})$ and the second term converges to 0. Therefore, it holds that $V_{A}^{f^{\prime}{\dagger}}(s_{0})\leq Q_{A}^{f^{\prime}{\dagger}}(s,f)\leq V_{A}^{fR_{B}^{*}(f)}(s_{0})=V_{A}^{f{\dagger}}(s_{0})$ for all $s_{0}\in\mathcal{S}$.
Proof of (b)
If it holds that $Q_{A}^{f^{\prime}{\dagger}}(s,f)=V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$, by the similar derivation of Equation 37, we obtain
$V_{A}^{f^{\prime}{\dagger}}(s_{0})=\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}^{*}(f)}\left[V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}\right]$ (38)
for all $s_{0}\in\mathcal{S}$. Then, in the limit for $n\to\infty$, we have $V_{A}^{f^{\prime}{\dagger}}(s_{0})=V_{A}^{f{\dagger}}(s_{0})$ for all $s_{0}\in\mathcal{S}$.
Conversely, if it holds that $V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$, we have
$V_{A}^{f{\dagger}}(s)$ $=r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[V_{A}^{f{\dagger}}(s^{\prime})\right]$ (39a)
$=r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ (39b)
$=Q_{A}^{f^{\prime}{\dagger}}(s,f)$ (39c)
for all $s\in\mathcal{S}$, which follows that $Q_{A}^{f^{\prime}{\dagger}}(s,f)=V_{A}^{f{\dagger}}(s)=V_{A}^{f^{\prime}{\dagger}}(s)$ for all $s\in\mathcal{S}$.
Proof of (c)
By the definition of $V_{A}^{f{\dagger}}(s)$, we can rewrite the reward value as
$r_{A}(s,f(s),R_{B}^{*}(s,f))=V_{A}^{f{\dagger}}(s)-\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})\right].$ (40)
Plugging it into the definition of $Q_{A}^{f^{\prime}{\dagger}}(s,f)$ and subtracting $V_{A}^{f^{\prime}{\dagger}}(s)$, we have
$Q_{A}^{f^{\prime}{\dagger}}(s,f)-V_{A}^{f^{\prime}{\dagger}}(s)$ $=r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]-V_{A}^{f^{\prime}}(s)$ (41)
$=V_{A}^{f{\dagger}}(s)-V_{A}^{f^{\prime}{\dagger}}(s)-\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$ (42)
The policy improvement is possible if and only if $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}$. Because of the above equality, equivalently we can say that the policy improvement from $f^{\prime}$ is possible if and only if there exists a policy $f$ whose value function satisfies
$V_{A}^{f{\dagger}}(s)-V_{A}^{f^{\prime}{\dagger}}(s)\geq\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ (43)
for all $s\in\mathcal{S}$, and there exists a state $s$ where the inequality strictly holds. ∎
Appendix H Proof of Lemma 6.4
Proof.
We prove the counterpart of Lemma 6.4. Taking the negation of Equation 7 for all $f\in\mathcal{W}_{A}$, we have
$\neg\forall f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}\implies V_{A}^{f{\dagger}}(s)=V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}\right\}$
$\iff\neg\forall f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)~{}\exists s\in\mathcal{S}\lor V_{A}^{f{\dagger}}(s)=V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}\right\}$ (44a)
$\iff\exists f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}\land V_{A}^{f{\dagger}}(s)\neq V_{A}^{f^{\prime}{\dagger}}(s)~{}\exists s\in\mathcal{S}\right\}.$ (44b)
From Lemma 6.3 (a), since $V_{A}^{f{\dagger}}(s)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$ holds and $Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$ holds, we have
$\exists f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}\land V_{A}^{f{\dagger}}(s)\neq V_{A}^{f^{\prime}{\dagger}}(s)~{}\exists s\in\mathcal{S}\right\}$
$\implies\exists f\in\mathcal{W}_{A}\left\{V_{A}^{f{\dagger}}(s)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}\land V_{A}^{f{\dagger}}(s)\neq V_{A}^{f^{\prime}{\dagger}}(s)~{}\exists s\in\mathcal{S}\right\}$ (45)
$\iff\exists f\in\mathcal{W}_{A}\left\{V_{A}^{f{\dagger}}\succ V_{A}^{f^{\prime}{\dagger}}\right\}$ (46)
$\implies\exists v\in\mathcal{V}\subseteq\mathrm{cl}\mathcal{V}\left\{v\succ V_{A}^{f^{\prime}{\dagger}}\right\}$ (47)
$\implies V_{A}^{f^{\prime}{\dagger}}\notin\mathcal{PV}.$ (48)
Therefore, $f^{\prime}$ is not a PO policy. This completes the proof. ∎
Appendix I Proof of Lemma 6.5
Proof.
First, we prove that (i) $\implies V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~{}\exists s\in\mathcal{S}$. Given $f\in\mathcal{W}_{A},f^{\prime}\in\mathcal{W}_{A}$, by the definition of $\delta(f,f^{\prime})$, we have
$\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]\leq\delta(f,f^{\prime})$ (49)
for all $s\in\mathcal{S}$. Thus, we have
$Q_{A}^{f^{\prime}{\dagger}}(s,f)$ $=r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ (50a)
$\geq r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\Bigl{(}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)\right]-\delta(f,f^{\prime})\Bigl{)}$ (50b)
[…]
$\geq\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}r_{A}(s_{t},f(s_{t}),R_{B}^{*}(s_{t},f))+\gamma_{A}^{n}V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}=s\right]-\left(\sum_{t=0}^{n-1}\gamma_{A}^{t}-\gamma_{A}^{0}\right)\delta(f,f^{\prime}).$ (50c)
for all $s\in\mathcal{S}$. In the limit for $n\to\infty$, the first term converges to $V_{A}^{f{\dagger}}(s)$, the second term converges to 0, and the third term converges to $-\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$. Therefore, we have
$Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$ (51)
for all $s\in\mathcal{S}$. Then, if (i) holds, because there exists $s\in\mathcal{S}$ such that $Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)-\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$, it holds that
$V_{A}^{f^{\prime}{\dagger}}(s)-\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,$ $f^{\prime})>Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})\quad\exists s\in\mathcal{S}.$ (52)
$\therefore V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)\quad\exists s\in\mathcal{S}.$ (53)
If (ii) holds, because of Lemma 6.3 (b), it holds that
$V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)\quad\forall s\in\mathcal{S}.$ (54)
Therefore,
Considering this, it holds that
$\forall f\in\mathcal{W}_{A}\{\mathrm{(i)~{}or~{}(ii)}\}\implies\forall f\in\mathcal{W}_{A}\left\{V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~{}~{}\exists s\in\mathcal{S}\mathrm{~{}~{}or~{}~{}}V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)~{}~{}\forall s\in\mathcal{S}\right\},$ (55)
where the right-hand side is equivalent to the definition of PO policies (Definition 6.1). ∎
Appendix J Proof of Theorem 6.6
Proof.
We prove each statement one by one.
Proof of (a)
As $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$, it holds that
$v_{t+1}=V_{A}^{f_{f+1}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\succcurlyeq V_{A}^{f_{t}{\dagger}}(\cdot)=v_{t}.$ (56)
Thus, $v_{t+1}\succcurlyeq v_{t}$ holds for all $t\in\mathbb{N}$.
Because $r_{A}$ is bounded, so is $V_{A}^{f{\dagger}}$, implying $\mathrm{cl}\mathcal{V}$ is compact. Considering the sequence $\{v_{t}(s)\}_{t=0}^{\infty}$ for each $s\in\mathcal{S}$, it is a monotonically increasing sequence. $\{v_{t}(s)\}_{t=0}^{\infty}$ converges in $\mathrm{cl}\mathcal{V}$ because $\mathrm{cl}\mathcal{V}$ is compact. Let $v_{\infty}(s)=\lim_{t\to\infty}v_{t}(s)$ for each $s\in\mathcal{S}$. Then, it holds that $v_{\infty}=\lim_{t\to\infty}v_{t}$.
Proof of (b)
First, $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}\impliedby v_{t}\stackrel{{\scriptstyle s}}{{=}}v_{\infty}$ is apparent.
Second, we prove $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}\implies v_{t}\stackrel{{\scriptstyle s}}{{=}}v_{\infty}$. As $f_{t+1}\in\mathcal{W}_{1-\epsilon}(f_{t})$, it holds that
$\mathcal{L}\left[V_{A}^{f_{t+1}{\dagger}}\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\geq\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\Bigl{(}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\Bigl{)}.$ (57)
If $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$, we have $\mathcal{L}[v_{t+1}]-\mathcal{L}[v_{t}]=0$; hence, the above inequality implies
$\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]=\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right];$ (58)
therefore, for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]=\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$. Because $\mathcal{L}$ is Pareto-compliant, the above equality implies $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\not\succ V_{A}^{f_{t}{\dagger}}$. However, because $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f_{t}{\dagger}}$, implying $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$. From Statement (b) of Lemma 6.3, we have $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. (Conversely, if $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$.) Thus, $\mathcal{W}_{\succcurlyeq}(f)=\mathcal{W}_{\succcurlyeq}(f_{t})$ holds for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. It implies that $\mathcal{W}_{\succcurlyeq}(f_{t+1})=\mathcal{W}_{\succcurlyeq}(f_{t})$ since $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$; thus, $f_{t+2}\in\mathcal{W}_{\succcurlyeq}(f_{t+1})=\mathcal{W}_{\succcurlyeq}(f_{t})$. Therefore,
$V_{A}^{f_{t}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t+1}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t+2}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}\cdots$ (59)
and hence $v_{t+k}\stackrel{{\scriptstyle s}}{{=}}v_{t}$ for all $k\geq 0$, implying that $v_{t}=v_{\infty}$.
Proof of (c)
First, we show that $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
Suppose that there exists $v^{*}\in\mathcal{V}$ such that $\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))>\gamma_{A}\left(\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))\right)$ holds. Then, $v^{*}\succ v_{\infty}$ and there exists $f^{*}\in\mathcal{W}_{A}$ whose value function is $V_{A}^{f^{*}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v^{*}$. Let $\tau=\frac{\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))}{\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))}$. We have $1\geq\tau>\gamma_{A}$.
For an arbitrarily small $s>0$ there exists $t$ such that $\lVert v_{t}-v_{\infty}\rVert_{\infty}\leq s$ because $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$. Because of the continuity of $\mathcal{L}$, it also implies that for an arbitrarily small $s^{\prime}>0$ there exists $t^{\prime}$ such that $\mathcal{L}[v_{t^{\prime}}]>\mathcal{L}[v_{\infty}]-s^{\prime}$. Therefore, we can find $t$ satisfying
$\lVert v_{\infty}-v_{t}\rVert_{\infty}$ $\leq\delta\frac{\tau-\gamma_{A}}{\gamma_{A}}\lVert v^{*}-v_{\infty}\rVert_{\infty}\quad\text{for some $\delta\in(0,1)$},\quad\text{and}$ (60)
$\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}]$ $<(1-\epsilon)(\mathcal{L}[\bar{v}]-\mathcal{L}[v_{\infty}]),\quad\text{where}\quad\bar{v}(s):=v_{\infty}(s)+(1-\delta)(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}.$ (61)
Let $f_{t}$ be the corresponding policy whose value function is $V_{A}^{f_{t}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v_{t}$. Here, we assume that $\gamma_{A}>0$. For the case of $\gamma_{A}=0$, Statement (c) is an immediate consequence of Statement (d) because $v_{\infty}\in\mathcal{PV}$ and $\mathcal{PV}\subseteq\partial\mathcal{V}$.
Subsequently, we show $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}_{\infty}$ using Equation 60. In the proof of Statement (c) of Lemma 6.3, for all $s\in\mathcal{S}$,
$Q_{A}^{f^{\prime}{\dagger}}(s,f)-V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)-V_{A}^{f^{\prime}{\dagger}}(s)-\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$ (62)
Notably, $\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]\leq\lVert V_{A}^{f{\dagger}}-V_{A}^{f^{\prime}{\dagger}}\rVert_{\infty}$. Letting $f=f^{*}$ and $f^{\prime}=f_{t}$, we have, for all $s\in\mathcal{S}$,
$Q_{A}^{f_{t}{\dagger}}(s,f^{*})-V_{A}^{f_{t}{\dagger}}(s)$ $\geq V_{A}^{f^{*}{\dagger}}(s)-V_{A}^{f_{t}{\dagger}}(s)-\gamma_{A}\lVert V_{A}^{f^{*}{\dagger}}-V_{A}^{f_{t}{\dagger}}\rVert_{\infty}.$ (63)
By adding $V_{A}^{f_{t}{\dagger}}(s)-v_{\infty}(s)$ to both sides of the above inequality, we obtain, for all $s\in\mathcal{S}$,
$Q_{A}^{f_{t}{\dagger}}(s,f^{*})-v_{\infty}(s)$ $\geq v^{*}(s)-v_{\infty}(s)-\gamma_{A}\lVert v^{*}-v_{t}\rVert_{\infty}$ (64a)
$\geq\min_{s\in\mathcal{S}}\left(v^{*}(s)-v_{\infty}(s)\right)-\gamma_{A}\lVert v^{*}-v_{t}\rVert_{\infty}$ (64b)
$=\min_{s\in\mathcal{S}}\left(v^{*}(s)-v_{\infty}(s)\right)-\gamma_{A}\lVert v^{*}-v_{\infty}+v_{\infty}-v_{t}\rVert_{\infty}$ (64c)
$\geq\min_{s\in\mathcal{S}}\left(v^{*}(s)-v_{\infty}(s)\right)-\gamma_{A}\lVert v^{*}-v_{\infty}\rVert_{\infty}-\gamma_{A}\lVert v_{\infty}-v_{t}\rVert_{\infty}$ (64d)
$=(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}-\gamma_{A}\lVert v_{\infty}-v_{t}\rVert_{\infty}$ (64e)
$\geq(1-\delta)(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}.$ (64f)
Therefore, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}$. Because $\bar{v}\succ v_{\infty}\succcurlyeq V_{A}^{f_{t}{\dagger}}$, it implies that $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq V_{A}^{f_{t}{\dagger}}$. From Statement (a) of Lemma 6.3, we have $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})$.
Finally, we derive a contradiction. Let $\ell_{t}^{\sup}=\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}[Q_{A}^{f_{t}{\dagger}}(\cdot,f)]$. Because $f^{*}\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $\ell_{t}^{\sup}\geq\mathcal{L}[Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})]$. We also know that $\mathcal{L}[Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})]\geq\mathcal{L}[\bar{v}]$. Because $f_{t+1}\in\mathcal{W}_{1-\epsilon}(f_{t})$, we obtain
$\mathcal{L}[v_{t+1}]-\mathcal{L}[v_{t}]$ $\geq\mathcal{L}[Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})]-\mathcal{L}[v_{t}]$ (65a)
$\geq(1-\epsilon)\left(\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}[Q_{A}^{f_{t}{\dagger}}(\cdot,f)]-\mathcal{L}[v_{t}]\right)$ (65b)
$\geq(1-\epsilon)(\mathcal{L}[Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})]-\mathcal{L}[v_{\infty}])$ (65c)
$\geq(1-\epsilon)(\mathcal{L}[\bar{v}]-\mathcal{L}[v_{\infty}])$ (65d)
$>\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}],$ (65e)
where we used Equation 61 for the last inequality. It implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to the fact that $v_{\infty}\succcurlyeq v_{t+1}$. Therefore, $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
Subsequently, we show that $v_{\infty}\in\partial\mathcal{V}$.
Suppose that $v_{\infty}\in\mathcal{V}\setminus\partial\mathcal{V}$ (i.e., $v_{\infty}$ is an interior of $\mathcal{V}$). Then, there exists an $r>0$ such that =$\mathcal{E}=\{v\in\mathcal{F}_{\mathcal{S}}\mid\lVert v-v_{\infty}\rVert_{\infty}\leq r\}\subseteq\mathcal{V}$. Let $v_{\infty}^{r}$ be such that $v_{\infty}^{r}(s)=v_{\infty}(s)+r$ for all $s\in\mathcal{S}$. Then, $v_{\infty}^{r}\in\mathcal{E}$. Moreover, because $v_{\infty}^{r}\in\mathcal{V}$, there exists a policy $f_{\infty}^{r}$ such that $V_{A}^{f_{\infty}^{r}{\dagger}}=v_{\infty}^{r}$.
We lead to the contradiction to $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$. Using Equation 60, we obtain
$\frac{\min_{s\in\mathcal{S}}V_{A}^{f_{\infty}^{r}{\dagger}}(s)-v_{\infty}(s)}{\max_{s\in\mathcal{S}}V_{A}^{f_{\infty}^{r}{\dagger}}(s)-v_{\infty}(s)}=\frac{\min_{s\in\mathcal{S}}(v_{\infty}^{r}(s)-v_{\infty}(s))}{\max_{s\in\mathcal{S}}(v_{\infty}^{r}(s)-v_{\infty}(s))}=1,$ (66)
contradicting $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$ for all $v\in\mathcal{V}$.
Proof of (d)
In case of $\gamma_{A}=0$, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}$ for all $f_{t},f\in\mathcal{W}_{A}$.
Suppose that $v_{\infty}\notin\mathcal{PV}$. Then, there exists a value function $v^{*}\in\mathrm{cl}\mathcal{V}$ such that $v^{*}\succ v_{\infty}$. Similar to the proof of Statement (c), we choose $t$ such that
$\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}]<(1-\epsilon)(\mathcal{L}[v^{*}]-\mathcal{L}[v_{\infty}]).$ (67)
Let $\mathcal{V}_{\succcurlyeq}(v_{t}):=\{v\in\mathcal{V}\mid v\succcurlyeq v_{t}\}$ and $\mathrm{cl}\mathcal{V}_{\succcurlyeq}(v_{t}):=\{v\in\mathrm{cl}\mathcal{V}\mid v\succcurlyeq v_{t}\}$. Then, we have $v^{*}\in\mathrm{cl}\mathcal{V}_{\succcurlyeq}(v_{t})$ because $v^{*}\succ v_{\infty}\succcurlyeq v_{t}$. Moreover, because $\mathcal{V}_{\succcurlyeq}(v_{t})=\{V_{A}^{f{\dagger}}:f\in\mathcal{W}_{\succcurlyeq}(f_{t})\}$, we obtain
$\ell_{t}^{\sup}:=$ $\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}[V_{A}^{f{\dagger}}]$ (68a)
$=$ $~{}\sup_{v\in\mathcal{V}_{\succcurlyeq}(v_{t})}\mathcal{L}[v]$ (68b)
$=$ $\max_{v\in\mathrm{cl}\mathcal{V}_{\succcurlyeq}(v_{t})}\mathcal{L}[v]$ (68c)
$\geq$ $~{}\mathcal{L}[v^{*}].$ (68d)
Therefore,
$\mathcal{L}[v_{t+1}]-\mathcal{L}[v_{t}]$ $\geq(1-\epsilon)(\ell_{t}^{\sup}-\mathcal{L}[v_{t}])$ (69a)
$\geq(1-\epsilon)(\ell_{t}^{\sup}-\mathcal{L}[v_{\infty}])$ (69b)
$\geq(1-\epsilon)(\mathcal{L}[v^{*}]-\mathcal{L}[v_{\infty}])$ (69c)
$>\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}],$ (69d)
where the derivation of Equation 69a is due to the definition of $\mathcal{W}_{1-\epsilon}(f_{t})$ in Equation 9 and $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}$. This implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to $v_{\infty}\succcurlyeq v_{t+1}$. Therefore, $v_{\infty}\in\mathcal{PV}$. ∎
Appendix K Proposed Algorithm with Backtracking
Algorithm 1 may not output a PO policy because the terminate condition in line 5 of Algorithm 1 is the necessary condition for PO policies. In this case, Lemma 6.5 can be used to determine whether the output policy is a PO policy: provided (i) or (ii) in Lemma 6.5 holds for all $f\in\mathcal{W}_{A}$ with the output policy $f_{t}$ and its state value function $V_{A}^{f_{t}{\dagger}}$, $f_{t}$ is a PO policy. Notably, this judgment is not perfect because the policy may be determined not to be a PO policy when it really is a PO policy.
We proposed selecting another policy $f^{\prime}\in\mathcal{W}(f_{t-1})$ that has not been selected thus far and restarting the for loop with $f_{t}\leftarrow f^{\prime}$, which we call ”Backtracking”, to deal with the case where the output policy $f_{t}(t<M)$ is determined not to be a PO policy. If there is no policy left in $\mathcal{W}(f_{t-1})$, then it backtracks one by one, as $\mathcal{W}(f_{t-2}),\mathcal{W}(f_{t-3}),\dots$, until it reaches the set with policies that have never been selected. This method requires that $\mathcal{W}(f_{t})$ is saved at each iteration.
The entire proposed algorithm including the PO-policy judgement and the backtracking is shown in Algorithm 2. $W$ is a list to save each $\mathcal{W}(f_{t})$ and $M$ is the maximum number of iterations. This algorithm is designed such that the number of times to compute $V_{A}^{f_{t}{\dagger}}$ in line 5 and $\mathcal{W}_{\succcurlyeq}(f_{t})$ in line 7 is at most $M$. If the PO-policy judgment in line 9 does not return True during $M$ iterations, the algorithm outputs the policy that maximizes the objective $\sum_{s\in\mathcal{S}}\alpha_{s}V_{A}^{f_{t}{\dagger}}(s)$ among the ones that satisfy the necessary condition for PO policies thus far and $f_{M}$ (see lines 12, 25, and 26).
Algorithm 2 PIPOP with Backtracking
0:  Pareto-compliant scalarization $\mathcal{L}$, maximum number of iterations $M$, a $M$-length list of empty sets $W$.
1:  $W(0)\leftarrow\mathcal{W}_{A}$
2:  Randomly sample $f_{0}$ from $W(0)$ without replacement.
3:  $\mathcal{W}_{output}\leftarrow\{\}$
4:  for $t=0$ to $M-1$ do
5:     Compute $V_{A}^{f_{t}{\dagger}}\in\mathcal{F}(\mathcal{S})$ such that $(T_{A}^{f_{t}R_{B}^{*}(f_{t})}V_{A}^{f_{t}{\dagger}})(s)=V_{A}^{f_{t}{\dagger}}(s)$ for all $s\in\mathcal{S}$.
6:     $Q_{A}^{f_{t}{\dagger}}(s,a,b)\leftarrow r_{A}(s,a,b)+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b)}\left[V_{A}^{f_{t}{\dagger}}(s^{\prime})\right]$ for all $s\in\mathcal{S},a\in\mathcal{A},b\in\mathcal{B}$.
7:     $\mathcal{W}_{\succcurlyeq}(f_{t})\leftarrow\{f\in\mathcal{W}_{A}~{}|~{}\mathbb{E}_{a\sim f(s)}\left[Q_{A}^{f_{t}{\dagger}}(s,a,R_{B}^{*}(s,f))\right]\geq V_{A}^{f_{t}{\dagger}}(s)~{}\forall s\in\mathcal{S}\}$.
8:     if $\mathbb{E}_{a\sim f(s)}\left[Q_{A}^{f_{t}{\dagger}}(s,a,R_{B}^{*}(s,f))\right]=V_{A}^{f_{t}{\dagger}}(s)~{}\forall s\in\mathcal{S},\forall f\in\mathcal{W}_{\succcurlyeq}(f_{t})$ then
9:        if $f_{t}$ satisfies (i) or (ii) of Lemma 6.5 for all $f\in\mathcal{W}_{A}$ then
10:           return $f^{*}\leftarrow f_{t}$.
11:        end if
12:        Put $f_{t}$ into $\mathcal{W}_{output}$.
13:        # Backtracking
14:        $k\leftarrow t$
15:        while $|W(k)|=0$ do
16:           $k\leftarrow k-1$
17:        end while
18:        Randomly sample $f_{t+1}$ from $W(k)$ without replacement.
19:     else
20:        $\mathcal{W}(f_{t})\leftarrow\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}\left[\mathbb{E}_{a\sim f(\cdot)}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,a,R_{B}^{*}(\cdot,f))\right]\right]$.
21:        $W(t+1)\leftarrow\mathcal{W}(f_{t})$.
22:        Randomly sample $f_{t+1}$ from $W(t+1)$ without replacement.
23:     end if
24:  end for
25:  Put $f_{M}$ into $\mathcal{W}_{output}$.
26:  return $f^{*}\in\operatornamewithlimits{argmax}_{f_{t}\in\mathcal{W}_{output}}\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}(\cdot)\right]$
26:  A stationary policy $f^{*}$
Appendix L Relation between the method of Zhang et al. (2020a) and Bucarey et al. (2022)
Zhang et al.Zhang et al. (2020a) defined the optimal value functions as $V_{i}^{*}(i\in\{A,B\})$, which are the solution of the following equationsZhang et al. (2020a): let $a\in\mathcal{A},b\in\mathcal{B},\gamma\in[0,1)$, and for all $s\in\mathcal{S}$,
$Q_{i}^{*}(s,a,b)=r_{i}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b)}\left[V_{i}^{*}(s^{\prime})\right],$ (70)
$V_{i}^{*}(s)=Stackelberg_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s)),$ (71)
where $Q_{i}^{*}(s)$ is the Q-value $Q_{i}(a,b,s)$ of an action pair $(a,b)\in\mathcal{A}\times\mathcal{B}$ in state $s$. Given the fixed $s$, $Q_{i}^{*}(b,a,s)$ can be viewed as the payoff of $(a,b)$ for the agent $i$; thus, $Q_{i}^{*}(s)$ can be viewed as the payoff table of agent $i$ in the state $s$. $Stackelberg_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s))$ is the payoff for agent $i$ in the Stackelberg equilibrium of the normal-form game represented by the payoff tables $(Q_{A}^{*}(s),Q_{B}^{*}(s))$. Therefore, Equation 71 can be written as
$V_{i}^{*}(s)$ $=Q_{i}^{*}(s,R_{A}^{\prime}(s),R_{B}^{\prime}(s,R_{A}^{\prime}(s))),$ (72a)
$\mathrm{where~{}}$ $R_{A}^{\prime}(s):=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}Q_{A}^{*}(s,a,R_{B}(s,a)),$ (72b)
$R_{B}^{\prime}(s,a):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}Q_{B}^{*}(s,a,b).$ (72c)
By substituting Equation 70 to Equation 72, we obtain
$V_{i}^{*}(s)=r_{i}(s,R_{A}^{\prime}(s),R_{B}^{\prime}(s,R_{A}^{\prime}(s)))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}^{\prime}(s),R_{B}(s,R_{A}^{\prime}(s)))}\left[V_{i}^{*}(s^{\prime})\right],$ (73a)
$\mathrm{where~{}}$ $R_{A}^{\prime}(s):=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}r_{A}(s,a,R_{B}^{\prime}(s,a))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,R_{B}(s,a))}\left[V_{A}^{*}(s^{\prime})\right],$ (73b)
$R_{B}^{\prime}(s,a):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b))}\left[V_{B}^{*}(s^{\prime})\right].$ (73c)
Meanwhile, the operator used in Bucarey et al.Bucarey et al. (2022) is given by
$(Tv)_{i}(s)$ $=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{b}))}\left[v_{i}(s^{\prime})\right],$ (74a)
$\mathrm{where~{}}$ $R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$ (74b)
$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right].$ (74c)
Notably, $\gamma=\gamma_{A}=\gamma_{B}$. Let us consider only the deterministic leader policies in the fixed point of Equation 74 and replace $\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}$ to $\operatornamewithlimits{argmax}_{a\in\mathcal{A}}$ in the definition of $R_{A}$. Then, letting $(V_{A},V_{B})$ be the fixed point of $T$, it holds from Equation 74, for all $s\in\mathcal{S}$, which is expressed as
$V_{i}(s)$ $=r_{i}(s,R_{A}(s,V),R_{B}(s,R_{A}(s,V),V_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),R_{B}(s,R_{A}(s,V),V_{B}))}\left[V_{i}(s^{\prime})\right],$ (75a)
$\mathrm{where~{}}$ $R_{A}(s,V):=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}r_{A}(s,a,R_{B}(s,a,V_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,R_{B}(s,a,V_{B}))}\left[V_{A}(s^{\prime})\right],$ (75b)
$R_{B}(s,a,V_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b)}\left[V_{B}(s^{\prime})\right].$ (75c)
Comparing these with Equation 73, the definition of $(V_{A}^{*},V_{B}^{*})$ by Equation 73 is equivalent to the definition of $(V_{A},V_{B})$ by Equation 75. It follows that the fixed point in Zhang et al.Zhang et al. (2020a) is the same in Bucarey et al.Bucarey et al. (2022) when the leader policies in the fixed point are restricted to deterministic stationary policies.
By applying the discussion in the proof of Theorem 5.1 to $(V_{A},V_{B})$ defined by Equation 75, we obtain
$V_{A}(s)=\max_{f\in\mathcal{W}_{A}^{d}}V_{A}^{fR_{B}(f,V_{B})}(s),\quad V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s),$ (76)
for all $s\in\mathcal{S}$. Then, the problem of the method of Bucarey et al. discussed in Section 5.1 also occurs to $(V_{A},V_{B})$: $V_{A}$ does not coincide with the state value function of the deterministic SSE policies since $R_{B}(f,V_{B})$ is not guaranteed to be the best response against any $f\in\mathcal{W}_{A}^{d}$.
Appendix M Application: Policy Teaching by Intervention to the Transitions
In single-agent MDPs, the policy aimed by the agent can be changed by making changes to the agent’s observations, reward signals, and transition probabilities, among others. For example, in an MDP $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$, the changes in the observations and the transition probabilities are represented by replacement to another transition function $p^{\prime}$, which changes the agent’s aim from the optimal policies of $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$ to those of $(\mathcal{S},\mathcal{A},p^{\prime},\rho,r,\gamma)$. This is the theoretical foundation of poisoning attacks against RL Behzadan & Munir (2017); Huang & Zhu (2019); Zhang et al. (2020b); Rakhsha et al. (2020); Sun et al. (2020) and policy teaching Zhang & Parkes (2008); Zhang et al. (2009), which aims to guide the agent’s policy learning.
SSGs can be viewed as a model of intervention in transition probabilities by the leader where the follower is the victim of the attack or the teaching. Even if the follower takes the same action at each state, the transition can be shifted by the leader changing its actions because the transition probability depends on the actions of all the agents. Therefore, suppose the follower cannot observe both the attendance and the actions of the leader, then the follower’s learning can be represented as learning the optimal policies of the single-agent MDP $(\mathcal{S},\mathcal{A}_{B},p^{f},\rho,\hat{r}_{B},\gamma_{B})$, where $f\in\mathcal{W}_{A}$ is the leader’s policy, $p^{f}(s^{\prime}|s,b):=\mathbb{E}_{a\sim f(s)}[\hat{p}(s^{\prime}|s,a,b)]$, and $\hat{r}_{B}:\mathcal{S}\times\mathcal{B}\to\mathbb{R}$. Notably, $R_{B}^{*}(f)$ is its optimal policy. Then, the problem of guiding the follower’s aim to the leader’s objective is expressed as
$\max_{f\in\mathcal{W}_{A}}\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\right],$ (77)
where $\hat{r}_{A}$ is the leader’s reward function that represents the desired follower’s behavior intended by the leader. $f^{*}$ is the optimal solution of Equation 77, and the state-action sequences generated by the corresponding follower’s optimal policy $R_{B}^{*}(f)$ maximize not only the follower’s expected cumulative discounted reward but also that of the leader’s. This can be viewed as the follower attaining the leader’s desired behavior when the follower achieves its own learning goal .
The optimal solution of Equation 77 is given by the PO policy $f$ that maximizes $\mathcal{L}[V_{A}^{f{\dagger}}]=\sum_{s\in\mathcal{S}}\rho(s)V_{A}^{f{\dagger}}(s)$. Therefore, our proposed approach can be applied to this problem when the transition function is known and the follower’s best response can be computed. Although the optimality of the obtained solution is not guaranteed unless $\gamma_{A}=0$, our analysis guarantees the monotone improvement of the leader’s policy, which is of great importance in practice. In contrast, existing methods trying to obtain the SSE policy may be inadequate because it is usual that the follower is not myopic and hence the SSE policy does not necessarily exist.

### v2

@@
-
+Policy Iteration for Two-Player General-Sum Stochastic Stackelberg Games
 Abstract
-In general-sum stochastic games, a stationary Stackelberg equilibrium (SSE) does not always exist, in which the leader maximizes leader’s return for all the initial states when the follower takes the best response against the leader’s policy.
-Existing methods of determining the SSEs require strong assumptions to guarantee the convergence and the coincidence of the limit with the SSE.
-Moreover, our analysis suggests that the performance at the fixed points of these methods is not reasonable when they are not SSEs.
-Herein, we introduced the concept of Pareto-optimality as a reasonable alternative to SSEs.
-We derive the policy improvement theorem for stochastic games with the best-response follower and propose an iterative algorithm to determine the Pareto-optimal policies based on it.
-Monotone improvement and convergence of the proposed approach are proved, and its convergence to SSEs is proved in a special case.
+We address two-player general-sum stochastic Stackelberg games (SSGs), where the leader’s policy is optimized considering the best-response follower whose policy is optimal for its reward under the leader.
+Existing policy gradient and value iteration approaches for SSGs do not guarantee monotone improvement in the leader’s policy under the best-response follower.
+Consequently, their performance is not guaranteed when their limits are not stationary Stackelberg equilibria (SSEs), which do not necessarily exist.
+In this paper, we derive a policy improvement theorem for SSGs under the best-response follower and propose a novel policy iteration algorithm that guarantees monotone improvement in the leader’s performance.
+Additionally, we introduce Pareto-optimality as an extended optimality of the SSE and prove that our method converges to the Pareto front when the leader is myopic.
+keywords:
+Stochastic Stackelberg Games; Policy Iteration; Convergence Analysis.
 1 Introduction
-A Markov decision process (MDP) is a mathematical framework modeling the decision-making of the agent in dynamic environments, which serves as the theoretical foundation of reinforcement learning (RL).
-The evaluation value of the agent’s policy is determined by assessing the state-action sequences generated by the policy in the environment with the reward function.
-This evaluation represents the expectation of cumulative discounted reward.
-An optimal policy is defined as one that maximizes the evaluation value.
-One of the extensions of MDPs for the multi-agent is a stochastic game, also known as a Markov game, where multiple agents with individual reward functions simultaneously try to develop policies that maximize their expected cumulative discounted rewards in one environment.
-Stochastic games are initially proposed in game theory, and the solution is defined as a tuple of policies that form an equilibrium, such as a Nash equilibrium.The problem setting is characterized by the relationship in the reward functions, such as the zero-sum setting and the fully cooperative setting.
-In this study, we addressed the problem of determining a Stackelberg equilibrium in 2-player stochastic games with general relationships in the rewards.
-A Stackelberg equilibrium is defined as a tuple of strategies of two types of agents, “leader” and “follower,” such that the leader’s strategy maximizes the leader’s payoff when the follower always takes the strategy that is the best response with respect to the follower’s payoff against the leader’s strategy.
-We aim to find the stationary policy that forms the Stackelberg equilibrium in general-sum stochastic games, that is called SSE.
-We assume that the follower’s best responses are always computable and the environment is known.
-Under an identical setting as ours, Bucarey et al.
-(2022) proposed an algorithm for finding the SSE.
-However, this algorithm requires strong assumptions to guarantee convergence to the SSEs.
-Moreover, this algorithm can converge to low-quality policies when it does not converge to the SSEs Bucarey et al.
-(2019).
-As an SSE does not always exist in the general-sum setting Bucarey et al.
-(2022), the algorithm should converge to the SSE if it exists; otherwise, the algorithm must converge to policies that achieve reasonably good performance.
-We have developed such an algorithm.
+A Markov decision process (MDP) is a mathematical framework that models the decision-making of agents in dynamic environments, which serves as the foundation of reinforcement learning (RL).
+The performance of the agent’s policy is evaluated by assessing the state-action sequences it produces with the reward function.
+This performance is quantified by the expected cumulative discounted reward.
+An optimal policy is one that maximizes this expected return.
+One extension of MDPs for multi-agent systems is a stochastic game, also known as a Markov game, where multiple agents with individual reward functions simultaneously attempt to develop policies that maximize their expected cumulative discounted rewards within a single environment.
+Stochastic games were initially proposed in game theory, and the solution is defined as a tuple of policies that form a specific type of equilibrium, such as a Nash equilibrium.
+The problem setting is also characterized by the relationship between the reward functions, such as the zero-sum (i.e., competitive) and fully cooperative settings.
+In this study, we addressed the problem of determining stationary Stackelberg equilibria (SSEs) in 2-player stochastic games with general relationships in the rewards.
+An SSE is defined as a tuple of stationary policies of two types of agents, leader and follower, such that the leader’s policy maximizes the leader’s payoff when the follower takes the policy that is the best response to the leader’s policy with respect to the follower’s payoff.
+We assume that the follower always takes the best response to any leader’s policy.
+We refer to such a follower as a best-response follower.
+Additionally, we assume that the follower’s best responses are always computable to the leader and that the environment is known.
+No other assumptions are made on the follower’s algorithm.
+This situation arises, for example, in the design of e-commerce platforms aimed at the maximization of the site owner’s profit.
+The site user (follower) repeatedly makes purchasing decisions based on their preferences.
+In contrast, the site owner (leader) can configure various elements of the platform, such as page transitions, advertisements, and an incentive bonus.
+Leveraging knowledge of the user’s past behavior, the owner seeks to maximize long-term profit by anticipating the user’s responses.
+Assuming the user’s responses are always optimal for their preferences and the owner’s prediction is accurate enough, the resulting owner’s problem can then be formulated as an SSG with a known best-response follower.
+As an SSE does not always exist in the general-sum setting (Bucarey et al., 2022), the algorithm should converge a leader’s policy to the SSE if it exists; otherwise, it must converge to a policy that achieves reasonably satisfactory performance.
+However, existing methods based on dynamic programming operators (Bucarey et al., 2022; Zhang et al., 2020a) and policy gradient methods with total derivatives (Zheng et al., 2022; Vu et al., 2022) do not guarantee the convergence to the SSEs in our setting, even if they exist.
+Moreover, these algorithms may converge leaders’ policies to low-quality ones because they do not guarantee monotone improvement of the leaders’ performance under the best-response follower.
+In this work, we focus on policy improvement methods and develop an algorithm that satisfies the above requirements.
+1.1 Main Contributions
 Our contributions are summarized as follows.
-First, theoretical analyses of the method proposed by Bucarey et al.
-(2022) reveal that its fixed point is not necessarily a reasonable policy when it is not guaranteed to converge to the Stackelberg equilibrium.
-(Section 5.1)
-Second, a new sufficient condition for the existence of the SSE includes the one derived in Bucarey et al.
-(2022).
-A necessary and sufficient condition for the case of the myopic leader is derived.
-(Section 5.2)
-Third, a concept of the Pareto-optimality is introduced as a reasonable alternative to the SSE.
-There always exist policies that realize the Pareto-optimal value functions or their neighborhood with arbitrary precision.
-Moreover, it agrees with the SSE if the equilibrium exists.
-(Section 6.1)
-Lastly, a policy improvement theorem for general-sum SSGs is derived (Section 6.2), based on which a novel algorithm is proposed.
-Monotone improvement of the state values and convergence of our algorithm are proven in this study (Section 6.3).
+First, we explicitly derive the fixed point of the dynamic programming operator used in Bucarey et al.
+(2022); Zhang et al.
+(2020a).
+This result reveals that the fixed point does not necessarily bring a reasonable leader policy when it is not guaranteed to be an SSE (Section˜5).
+Second, a policy improvement theorem for general-sum stochastic Stackelberg games with the best-response follower is derived (Section˜6.2), based on which a novel policy iteration algorithm is proposed (Section˜6.3).
+Finally, we introduce the concept of Pareto-optimality as an extended optimality of the SSE.
+Policies that realize the Pareto-optimal value functions or their neighborhood with arbitrary precision always exist.
+Moreover, Pareto-optimality agrees with the SSE it exists (Section˜6.1).
+Then, we prove that the proposed method monotonically improves the state values toward the Pareto front and converges to the front when the leader is myopic (Section˜6.3).
+To the best of our knowledge, this is the first theoretical guarantee in general-sum myopic-leader SSGs.
 2 Preliminaries
 2.1 Markov Decision Process
-An MDP is a stochastic process where the state transition depends on the agent’s action and the rewards are defined.
-MDPs are defined as a tuple $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$: a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, an initial state distribution $\rho\in\Delta(\mathcal{S})$, a bounded reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a discount rate $\gamma\in[0,1)$, where $\Delta(\mathcal{S})$ is a set of probability distributions on $\mathcal{S}$.
+An MDP is a stochastic process with rewards in which the state transitions depend on the actions of an agent.
+An MDP is defined as a tuple $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$: a finite state space $\mathcal{S}$, a finite action space $\mathcal{A}$, a transition function $p:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, an initial state distribution $\rho\in\Delta(\mathcal{S})$, a bounded reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a discount rate $\gamma\in[0,1)$, where $\Delta(\mathcal{S})$ is a set of probability distributions on $\mathcal{S}$.
 The agent determines its action following its Markov policy.
-Let $\mathcal{W}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ be a set of (stochastic) decision rules, then at the time step $t\in\mathbb{N}$, the agent has the decision rule $f_{t}\in\mathcal{W}$ and selects the action $a_{t}\in\mathcal{A}$ in the current state $s_{t}\in\mathcal{S}$ as $a_{t}\sim f_{t}(s_{t})$.
+Let $\mathcal{W}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ be a set of (stochastic) decision rules.
+Then, at the time step $t\in\mathbb{N}$, the agent has the decision rule $f_{t}\in\mathcal{W}$ and selects the action $a_{t}\in\mathcal{A}$ in the current state $s_{t}\in\mathcal{S}$ as $a_{t}\sim f_{t}(s_{t})$.
 The conditional probability is denoted as $f_{t}(a_{t}|s_{t})$.
-If the decision rule is equal at any time step ($f_{t}=f$), the Markov policy is called a stationary policy and denoted by a single decision rule $f\in\mathcal{W}$.
+If the decision rule is invariant over time ($f_{t}=f$), the Markov policy is called a stationary policy and is simply denoted by $f\in\mathcal{W}$.
 The next state $s_{t+1}$ is stochastically determined as $s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})$ when the agent selects its action $a_{t}$ in the state $s_{t}$.
@@
 $V^{f}(s):=\mathbb{E}^{f}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\middle|s_{0}=s\right],$
-for all $s\in\mathcal{S}$, where $\mathbb{E}^{f}\left[\cdot\right]$ is the expectation over $\{(s_{t},a_{t})\}_{t=0}^{\infty}$ generated under $(f,p,\rho)$.
-$V^{f}(s)$ is called a state value function of $f$.
-There always exists a deterministic optimal policy Sutton & Barto (2018) such that, for all $s\in\mathcal{S}$,
-$f^{*}(s)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\bigg{\{}r(s,a)\\ +\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\max_{f\in\mathcal{W}}V^{f}(s^{\prime})\right]\bigg{\}}.$ (1)
-The main algorithms used to determine the optimal policy are value iteration and policy iteration when the transition function and the reward function are known and RL when they are unknown Sutton & Barto (2018).
+for all $s\in\mathcal{S}$, where $\mathbb{E}^{f}\left[\cdot|s_{0}=s\right]$ is the expectation over $\{(s_{t},a_{t})\}_{t=0}^{\infty}$ generated under $(f,p)$ given initial state $s$.
+$V^{f}(s)$ is called a state value function of stationary policy $f$.
+Hereafter, we consider only stationary policies and refer to them simply as policies.
+There always exists a deterministic optimal policy (Sutton and Barto, 2018) such that
+$f^{*}(s)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}\bigg\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\max_{f\in\mathcal{W}}V^{f}(s^{\prime})\right]\bigg\}$ (1)
+for all $s\in\mathcal{S}$.
+The main algorithms used to determine the optimal policy are value iteration and policy iteration when the transition and reward functions are known, and RL when they are unknown (Sutton and Barto, 2018).
 2.2 Stochastic Game and Stackelberg Game
-Stochastic games Shapley (1953); Solan (2022), also called Markov games, are the extension of MDPs for the multi-agent.
-When there are two agents, A and B, they have finite action spaces $\mathcal{A}$ and $\mathcal{B}$, bounded reward functions $r_{A}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$ and $r_{B}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$, discount rates $\gamma_{A}\in[0,1)$ and $\gamma_{B}\in[0,1)$ and sets of decision rules $\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ and $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\}$, respectively.
-The state $s\in\mathcal{S}$ is shared among the agents, and the transition function is defined as $p:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta(\mathcal{S})$, that is, the transition probability depends on the actions of all the agents.
+Stochastic games (Shapley, 1953; Solan, 2022), also known as Markov games, are an extension of MDPs for multi-agent settings.
+When there are two agents, A and B, they have finite action spaces $\mathcal{A}$ and $\mathcal{B}$, bounded reward functions $r_{A}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$ and $r_{B}:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\mathbb{R}$, discount rates $\gamma_{A}\in[0,1)$ and $\gamma_{B}\in[0,1)$ and sets of policies $\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\}$ and $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\}$, respectively.
+The state $s\in\mathcal{S}$ is shared among the agents, and the transition function is defined as $p:\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta(\mathcal{S})$; that is, the transition probability depends on the actions of all the agents.
 The expectation of the cumulative discounted reward under stationary policies $f\in\mathcal{W}_{A}$ and $g\in\mathcal{W}_{B}$ is defined for $r_{A}$ and $r_{B}$, respectively.
-Then, a pair of policies that are Nash equilibrium or other equilibria can be defined as the solutions to the problem by considering each cumulative reward as the payoff of the agent.
+Then, equilibrium policies, such as Nash equilibrium, can be identified as solutions to the problem by considering each cumulative reward as the agents’ payoff.
 The relationship of the reward functions between agents characterizes the game.
-In game theory, the game where $r_{A}=-r_{B}$ is called zero-sum game, and the others are called non-zero-sum game.
+In game theory, a game where $r_{A}=-r_{B}$ is called a zero-sum game, and the others are called non-zero-sum games.
 Both are called general-sum games.
 In the context of a multi-agent RL, situations where $r_{A}\propto-r_{B}$ and $r_{A}\propto r_{B}$ are called competitive and cooperative, respectively.
-The Stackelberg game is a game model where a leader first selects its strategy and then a follower selects its strategy such that it maximizes the follower’s payoff against the leader’s strategy.
+The Stackelberg game is a game model in which a leader selects its strategy first, and then a follower selects its strategy that maximizes the follower’s payoff against the leader’s strategy.
 This follower’s strategy is called the best response strategy against the leader’s strategy.
-Then, when the leader’s strategy maximizes the leader’s payoff under the follower’s best response, the pair of strategies is called a Stackelberg equilibrium.
-In this study, we consider the extension of Stackelberg games to stochastic games, such as stochastic Stackelberg game (SSG).
-An example of the application of SSG is provided in Appendix M.
-Unlike the single-agent MDP, the existence of the optimal policy, the SSE policy defined below, is not guaranteed, and the SSE policy is not deterministic even if it exists.
+When the leader’s strategy maximizes the leader’s payoff under the follower’s best response, the pair of strategies is called a Stackelberg equilibrium.
+In this study, we consider the extension of Stackelberg games to stochastic games, such as the stochastic Stackelberg game (SSG).
+Unlike the single-agent MDP, the optimal policy, SSE policy defined below, is not guaranteed to exist, and even if it does, it is not necessarily deterministic.
+An example scenario where the SSE policy does not exist is provided in the next section.
 3 Problem Setting
-A general-sum SSG is formulated. Its solutions are defined as the policies that are Stackelberg equilibria.
-In a 2-player stochastic game represented by $(\mathcal{S},(\mathcal{A},\mathcal{B}),\hat{p},\rho,(\hat{r}_{A},\hat{r}_{B}),(\gamma_{A},\gamma_{B}))$, agent A is called a “leader” and agent B is called a “follower.” The sets of decision rules are defined by
-$\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\},$ $\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\},$
-and the sets of deterministic decision rules are defined by
-$\mathcal{W}_{A}^{d}:=\{f:\mathcal{S}\to\mathcal{A}\},$ $\mathcal{W}_{B}^{d}:=\{g:\mathcal{S}\to\mathcal{B}\}.$
+A general-sum SSG is represented by $(\mathcal{S},(\mathcal{A},\mathcal{B}),\hat{p},\rho,(\hat{r}_{A},\hat{r}_{B}),(\gamma_{A},\gamma_{B}))$, where the subscripts $A$ and $B$ indicate the leader and follower, respectively.
+The sets of policies are defined by
+$\mathcal{W}_{A}:=\{f:\mathcal{S}\to\Delta(\mathcal{A})\},\quad\mathcal{W}_{B}:=\{g:\mathcal{S}\to\Delta(\mathcal{B})\},$
+and the sets of deterministic policies are defined by
+$\mathcal{W}_{A}^{d}:=\{f:\mathcal{S}\to\mathcal{A}\},\quad\mathcal{W}_{B}^{d}:=\{g:\mathcal{S}\to\mathcal{B}\}.$
 Let $\mathcal{F}_{\mathcal{S}}:=\{v:\mathcal{S}\to\mathbb{R}\}$ be the set of real-valued functions with a state as input.
@@
 $V_{i}^{fg}(s):=\mathbb{E}^{fg}\left[\sum_{t=0}^{\infty}\gamma_{i}^{t}\hat{r}_{i}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right],$
-where $\mathbb{E}^{fg}\left[\cdot\right]$ is the expectation over the state-action sequence $\{(s_{t},a_{t},b_{t})\}_{t=0}^{\infty}$ generated under $(f,g,\hat{p},\rho)$.
+[…]
 The follower’s best response is defined by employing the state value functions.
 Definition 3.2 (Follower’s best response).
-Let $s\in\mathcal{S}$ be a state and $f\in\mathcal{W}_{A}$ a leader’s decision rule. The set of the follower’s best responses $R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$ against $f$ is defined as
+Let $s\in\mathcal{S}$ be a state and $f\in\mathcal{W}_{A}$ a leader’s policy.
+The set of the follower’s best responses $R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$ against the leader’s policy $f$ is defined as
 $R_{B}^{*}(f):=\{g\in\mathcal{W}_{B}^{d}~|~g(s)\in R_{B}^{*}(s,f)~\forall s\in\mathcal{S}\},$
-where
-$R_{B}^{*}(s,f):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f(s),b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),b)}\left[\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s^{\prime})\right].$
-As the optimal policies in single-agent MDPs are represented as in Equation 1, we can observe that the set of best responses $R_{B}^{*}(f)$ is the set of the optimal policies with respect to single-agent MDPs when the leader’s policy is fixed to the stationary policy $f$.
+where the best response action at state $s$ is defined as
+$R_{B}^{*}(s,f):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f(s),b)+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),b)}\left[\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s^{\prime})\right].$
+As the optimal policies in single-agent MDPs are represented as in Equation˜1, we can observe that the set of best responses $R_{B}^{*}(f)$ with fixed leader’s stationary policy $f$ is the set of the optimal policies in terms of single-agent MDPs.
 It is also shown that $R_{B}^{*}(f)$ is non-empty for all $f\in\mathcal{W}_{A}$.
-The follower that always adopts the best response against the leader’s policy is called the “best-response follower.” If the leader adopts a stationary policy $f\in\mathcal{W}_{A}$, the best-response follower adopts the stationary policy $g\in R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$.
-We assume that $R_{B}^{*}(s,f)$ is a singleton for all $s\in\mathcal{S},f\in\mathcal{W}_{A}$.
-It follows that $R_{B}^{*}(f)$ is also singleton for all $f\in\mathcal{W}_{A}$; thus, the follower’s best response against $f$ is unique.
+We assume that the follower is a best-response follower, meaning that it always adopts the best response to the leader’s policy.
+If the leader adopts a stationary policy $f\in\mathcal{W}_{A}$, the best-response follower adopts the stationary policy $g\in R_{B}^{*}(f)\subseteq\mathcal{W}_{B}^{d}$.
+This setup is essentially equivalent to a situation where we can only control the leader’s policy, and the follower independently learns its policy with the ability to find the best responses in a finite amount of time for any leader’s policy.
+No other assumptions are made about the follower’s learning process.
+We assume that the set of the follower’s best responses, $R_{B}^{*}(s,f)$, is a singleton for each $s\in\mathcal{S}$ and for each $f\in\mathcal{W}_{A}$.
+Otherwise, we break ties deterministically.
+It follows that $R_{B}^{*}(f)$ is also a singleton for all $f\in\mathcal{W}_{A}$; thus, the follower’s best response against $f$ is unique.
 For simplicity, let $R_{B}^{*}(f)$ be the best response.
@@
 Let $s\in\mathcal{S}$ be a state. An SE value function is defined as
-$V_{A}^{*}(s):=\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$
+$V_{A}^{*}(s):=\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s).$
 Definition 3.4 (SSE policy).
 If a leader’s stationary policy $f^{*}\in\mathcal{W}_{A}$ satisfies $V_{A}^{f^{*}{\dagger}}(s)=V_{A}^{*}(s)$ for all $s\in\mathcal{S}$, then $f^{*}$ is an SSE policy.
-Notation
-[…]
-4 Related Works
+Game without SSEs
+We provide an example game where SSEs do not exist.
+This example is first introduced by Jean-Marie in his seminar talk (Jean-Marie et al., 2022) related to Bucarey et al.
+(2022).
+We consider a game with two states $\mathcal{S}=\{s_{1},s_{2}\}$, two leader’s actions $\mathcal{A}=\{a_{1},a_{2}\}$, and two follower’s actions $\mathcal{B}=\{b_{1},b_{2}\}$.
+The dynamics and the reward signals are deterministic and represented in Figure˜1.
+Let the leader’s policy be denoted as $f_{s_{1}}(a_{1})=p\in[0,1]$ and $f_{s_{2}}(a_{1})=q\in[0,1]$.
+Given the leader’s policy, it reduces to a standard MDP for the follower. The optimal value function for the follower is the solution to
+$V_{B}^{*}(s_{1})$ $=\max\{\gamma_{B}V_{B}^{*}(s_{1}),x+\gamma_{B}V_{B}^{*}(s_{2})\}\cdot p+(-y+\gamma_{B}V_{B}^{*}(s_{2}))\cdot(1-p);$
+$V_{B}^{*}(s_{2})$ $=\max\{\gamma_{B}V_{B}^{*}(s_{2}),x+\gamma_{B}V_{B}^{*}(s_{1})\}\cdot q+(-y+\gamma_{B}V_{B}^{*}(s_{1}))\cdot(1-q).$
+There are three cases depending on the leader’s policy $f$, summarized in Table˜1.
+If $\gamma_{B}y>x$, we know that $(p,q)=(1,0)$ leads to the case of $\gamma_{B}V_{B}^{*}(s_{1})>x+\gamma_{B}V_{B}^{*}(s_{2})$.
+Then, the maximum value of the leader’s value at state $s_{1}$ is obtained as $V_{A}^{f{\dagger}}(s_{1})=\frac{1}{1-\gamma_{A}}$.
+Similarly, we can check that $(p,q)=(0,1)$ leads to the case of $\gamma_{B}V_{B}^{*}(s_{2})>x+\gamma_{B}V_{B}^{*}(s_{1})$.
+Then, the maximum value of the leader’s value at state $s_{2}$ is obtained as $V_{A}^{f{\dagger}}(s_{2})=\frac{1}{1-\gamma_{A}}$.
+We can see that there is a tradeoff between the values at $s_{1}$ and $s_{2}$ and two Pareto optimal points exist: $(V_{A}^{f{\dagger}}(s_{1}),V_{A}^{f{\dagger}}(s_{2}))=(\frac{1}{1-\gamma_{A}},\frac{\gamma_{A}}{1-\gamma_{A}})$ and $(\frac{\gamma_{A}}{1-\gamma_{A}},\frac{1}{1-\gamma_{A}})$.
+That is, there does not exist a single optimal policy that maximizes the values at all states simultaneously, implying that an SSE does not exist.
+Table 1: The follower’s best response and the leader’s value under the best response follower.
+Case $R_{B}^{*}(s_{1},f),R_{B}^{*}(s_{2},f)$ $V_{A}^{f{\dagger}}(s_{1})$, $V_{A}^{f{\dagger}}(s_{2})$
+$\gamma_{B}V_{B}^{*}(s_{1})>x+\gamma_{B}V_{B}^{*}(s_{2})$ $b_{1},b_{2}$ $\frac{p}{1-\gamma_{A}},\frac{p\gamma_{A}}{1-\gamma_{A}}$
+$\gamma_{B}V_{B}^{*}(s_{2})>x+\gamma_{B}V_{B}^{*}(s_{1})$ $b_{2},b_{1}$ $\frac{q\gamma_{A}}{1-\gamma_{A}},\frac{q}{1-\gamma_{A}}$
+$\lvert V_{B}^{*}(s_{1})-V_{B}^{*}(s_{2})\rvert<x/\gamma_{B}$ $b_{2},b_{2}$ $0,0$
+$s_{1}$$s_{2}$($a_{1}$, $b_{1}$) / (1, 0)($a_{1}$, $b_{1}$) / (1, 0)($a_{1}$, $b_{2}$) / (0, $x$) ($a_{2}$, *) / (0, y-y)
+Figure 1: An example of a game without SSEs.
+Transitions are deterministic and represented by arrows.
+Each label shows (leader’s action, follower’s action) /  (leader’s reward, follower’s reward), where $x>0$ and $y>0$.
+4 Related Work
 Bucarey et al.
-(2022) proposed an algorithm for “strong” SSEs in the 2-player general-sum stochastic game.
-Assuming the uniqueness of the follower’s best response, the problem setting of Bucarey et al.
-(2022) is the same as ours.
-However, their algorithm requires strong assumptions to guarantee convergence to the SSEs, and there is little assurance of the performance of the obtained policy when it does not converge to the SSEs or when there is no SSE.
-This problem is because the fixed point of the operator used in Bucarey et al.
+(2022) proposed an algorithm for computing strong SSEs in the 2-player general-sum stochastic games, where strong SSEs are introduced to break ties for the follower’s best response, which corresponds to a specific tie-breaking mechanism for the follower’s best response in our setting.
+They provide sufficient conditions for their algorithm to converge to the SSEs.
+However, there is little assurance of the leader’s performance of the obtained policy when it does not converge to the SSEs or when there is no SSE.
+This problem is because the fixed point of the dynamic programming operator used in Bucarey et al.
 (2022) is different from the SE value function.
-We discuss this in detail in Section 5.1.
-A case where the algorithm of Bucarey et al.
-(2022) converges to the SSE is when the follower is myopic, which maximizes its immediate reward.
-There are some existing studies in the SSGs under the myopic follower.
+We discuss this in detail in Section˜5.
+A primary case where the algorithm of Bucarey et al.
+(2022) converges and the fixed point coincides with the SE values is when the follower is myopic, i.e., when the follower maximizes its immediate rewards.
+This setting is adopted in some existing SSG studies.
 For example, Zhao et al.
@@
 Zhong et al.
-(2023) proposed the guaranteed algorithm in the SSG where the group of myopic followers forms a Nash equilibrium among them.
-Different from these works, we focus on the situation where the follower is not myopic.
-Some studies concentrate on the cooperative setting.
+(2023) proposed an algorithm in the SSG where the group of myopic followers forms a Nash equilibrium among them with analyses on the upper bound of the regret and suboptimality of performance.
+In contrast to these studies, we focus on the situation where the follower is not myopic.
+Zhang et al.
+(2020a) also considered a similar problem setting, where the reward setting is general-sum and the follower is not myopic.
+The differences are that the transition function is unknown (i.e., in the RL setting), the follower’s policy takes the leader’s action as input, and only deterministic stationary policies are considered for the leader’s policies.
+The algorithm proposed by Zhang et al.
+(2020) shares its principal foundation with that of Bucarey et al.
+(2022) (see Appendix˜I).
+This means that they use the same operator, implying that the limitation of the method of Bucarey et al.
+(2022) applies to that of Zhang et al.
+(2020).
+Policy gradient methods have been proposed for SSGs with non-myopic followers (Zheng et al., 2022; Vu et al., 2022) based on the implicit function theorem (Fiez et al., 2020).
+They guarantee the convergence to differential SEs (DSEs), which are subsets of local SEs.
+However, these methods are not applicable in our setting, except when using the surrogate model of the follower, due to their centralized nature of learning.
+Moreover, the leader’s policy obtained by these algorithms does not necessarily form the DSEs under the best-response follower because the follower’s strategies in DSEs are not always the best response.
+Several studies concentrate on the cooperative setting.
 Kao et al.
 (2022) and Zhao et al.
-(2023) analyzed the cooperative tasks with information asymmetry, where the leader cannot observe the follower’s action Kao et al.
-(2022) and the reward functions are known only to the follower Zhao et al.
-(2023).
+(2023) analyzed the cooperative tasks with information asymmetry, where the leader cannot observe the follower’s action (Kao et al., 2022) and the reward functions are known only to the follower (Zhao et al., 2023).
 Kononen (2004) and Zhang et al.
 (2020a) point out Pareto efficiency and uniqueness as advantages of Stackelberg equilibria to Nash equilibria in cooperative tasks.
-Zhang et al.
-(2020a) is similar to ours and Bucarey et al.
-(2022) in that the problem setting is the general-sum and the follower is not myopic.
-The differences are that the transition function is unknown (i.e., in the RL setting), the follower’s policy takes the leader’s action as input, and only deterministic stationary policies are considered the leader’s policies.
-The algorithm proposed by Zhang et al.
-(2020a) shares its principal foundation with that of Bucarey et al.
-(2022) (see Appendix L).
-This means that these methods use the same operator, implying that the limitation of the method of Bucarey et al.
-(2022) applies to that of Zhang et al.
-(2020a).
-We proposed a novel algorithm for general-sum SSGs where the transition function and the reward functions are known and the follower is not myopic, which is the same as Bucarey et al.
-(2022).
-Different from the previous work, we proposed a novel approach with a convergence guarantee even when no SSE policy exists to deal with such a case.
-5 Fixed Point Analysis of Operators
-5.1 Operator defined by Bucarey et al. (2022)
-Bucarey et al.
-(2022) addressed the same problem as ours under the uniqueness of $R_{B}^{*}$ [^1] Bucarey et al.
-(2022), which defines Strong SSEs to avoid the ambiguity of the best responses and utilizes it as the optimal solution of the problem.
-In this study, instead, we assume that there is no ambiguity in the best responses; hence, the SSEs are indeed strong SSEs.
-.
-They defined an operator whose fixed point coincides with the SE value function under certain conditions.
-Then, they proposed a method to construct a stationary equilibrium policy corresponding to the fixed point of the defined operator, which is called fixed point equilibrium (FPE).
-However, they required strong assumptions that can potentially limit the applications of their approach (see for example Appendix M).
-First, they assumed that the follower is “myopic” or the reward functions are either cooperative or competitive to guarantee the convergence of their algorithm to the FPE Bucarey et al.
-(2022); Zhang et al.
-(2020a).
-Second, to guarantee the coincidence of the FPE and the SSE, the myopic follower was assumed Bucarey et al.
-(2022).
-Notably, the situation where the follower is “myopic” in Bucarey et al.
-(2022) is equivalent to the one where the follower’s discount rate $\gamma_{B}=0$ or the one where the transition function is not dependent on the follower’s actions Bucarey et al.
-(2022).
-That is, if the follower is not myopic, the algorithm may not converge and the fixed point of the algorithm may be different from the SSE.
-Even if the convergence of the algorithm is not guaranteed, it is of great significance from a practical viewpoint to understand the characteristics of the FPE if exists.
-Particularly, it is interesting to examine how effective the FPE is if it does not coincide with the SSE.
-Thus, we analyzed the fixed point of the operator defined in Bucarey et al.
-(2022).
-The operator $T:\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ defined by Bucarey et al.
-(2022) for the strong SSE policies can be simplified as one for the SSE policies as: for $v:=(v_{A},v_{B})\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$, $i\in\{A,B\}$ and $s\in\mathcal{S}$,
-$(Tv)_{i}(s):=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))\\ +\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))}\left[v_{i}(s^{\prime})\right],$ (2)
-where $R_{A}(s,v)$ is
-$R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$
-and is assumed to exist for all $s$ and $v$, [^2] This assumption is necessary to derive a policy corresponding to a value function and Bucarey et al.
-(2022) implicitly assume it.
-However, it is not guaranteed because the inside of $\operatornamewithlimits{argmax}$ is not necessarily continuous with respect to $f_{s}$, even though $\Delta(\mathcal{A})$ is compact.
-$R_{B}$ always exists because $\mathcal{B}$ is finite, and
-$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right]$
-is assumed to be a singleton for all $s$, $f_{s}$, and $v_{B}$.
+We proposed a novel algorithm for general-sum SSGs with the non-myopic best-response follower under the assumptions that the transition function is known and that the follower’s best response is computable in a reasonable time.
+Unlike the previous methods, iterative methods by the operator (Bucarey et al., 2022; Zhang et al., 2020a) and policy gradient methods (Zheng et al., 2022; Vu et al., 2022), our algorithm guarantees monotone improvement of the leader’s state values under the best-response follower.
+Furthermore, we give a convergence guarantee even when no SSE policy exists.
+5 Analysis on DP Operators
+Dynamic programming operators are the core of the solutions of single-agent MDPs and the foundation of RL algorithms (Sutton and Barto, 2018).
+For stochastic games, the operator is extended as an operation of solving a one-step game by multiple players, and methods for Nash equilibria are proposed with convergence guarantees (Hu and Wellman, 2003).
+For SSGs, however, existing methods based on such a one-step game operator (Bucarey et al., 2022; Zhang et al., 2020a) have problems in terms of the leader’s performance in our setting.
+This section describes it in detail.
+In summary, we analyze the fixed points of the operator and show that, while the equilibrium formed by the fixed point holds under the best-response follower, the obtained leader’s policy may perform poorly when the fixed point is not the SSE, since monotone improvement in the leader’s performance is not guaranteed.
+The one-step game operator $T:\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ is defined for SSGs as
+$(Tv)_{i}(s):=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))\\ +\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))}\left[v_{i}(s^{\prime})\right]\quad(i\in\{A,B\}),$ (2)
+where $v:=(v_{A},v_{B})\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$,
+$R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$
+which is assumed to exist for all $s$ and vv, [^1] This assumption is necessary to derive a policy corresponding to a value function, and Bucarey et al.
+(2022) implicitly assumes it.
+However, it is not guaranteed in general because the inside of $\operatornamewithlimits{argmax}$ is not necessarily continuous with respect to $f_{s}$, even though $\Delta(\mathcal{A})$ is compact.
+Conversely, $R_{B}$ always exists because $\mathcal{B}$ is finite.
+and
+$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)\\ +\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right],$
+which is assumed to be a singleton for all $s$, $f_{s}$, and $v_{B}$.
 [^2] Bucarey et al.
-(2022) defines the “strong” version of $R_{B}$ similarly to the strong SSE.
-Thus, we assume that $R_{B}$ is a singleton to simplify the discussion in accordance with our goal to find the SSE policies.
+(2022) defines the strong version of $R_{B}$ similarly to the strong SSE.
+In this paper, we assume that $R_{B}$ is a singleton to simplify the discussion, in accordance with our goal of finding the SSE policies.
 As the core of our analysis does not depend on the uniqueness of $R_{B}$, the result can be easily extended to the method of Bucarey et al.
 (2022).
-If there exists a fixed point $V:=(V_{A},V_{B})$ of $T$, the pair of the value functions of the FPE $(\bar{f},\bar{g})$ defined as follows coincides with $V$ Bucarey et al.
-(2022): for any $s\in\mathcal{S}$,
-$\bar{f}(s)\in R_{A}(s,V),$ $\bar{g}(s)\in R_{B}(s,\bar{f}(s),V_{B}).$
-$T$ is a contraction mapping for $v_{A}$ under certain conditions [^4] Bucarey et al.
+Equation˜2 is regarded as a simplified version of the operator defined in Bucarey et al.
 (2022).
-An example is the case when $r_{A}$ is an affine transformation of $r_{B}$.
-In this situation, the game is either cooperative or competitive..
-Then, we can obtain the fixed point asymptotically by repeatedly applying such $T$ to an arbitrary initial function $v_{A}^{(0)}\in\mathcal{F}_{\mathcal{S}}$ Bucarey et al.
-(2022).
-In this case, a single update by $T$ can be interpreted as solving a normal-form Stackelberg game on each state $s\in\mathcal{S}$ by considering the current Q functions
-$q_{i}^{(t)}(s,a,b):=r_{i}(s,a,b)+\gamma_{i}\mathbb{E}_{s^{\prime}}\left[v_{i}^{(t)}(s^{\prime})\right]\quad(i\in\{A,B\})$
-as the gain tables on $s$ for the leader and the follower.
-This operation is also used in Nash-Q Learning Hu & Wellman (2003), which is guaranteed to converge to the Nash equilibrium in stochastic games.
-However, the convergence to the Stackelberg equilibrium by the same operation is not guaranteed even if the equilibrium exists.
-Let us consider the characteristics of the fixed point $(V_{A},V_{B})$ of the operator $T$ in general situations where the fixed point exists but it is different from the SE value functions.
-Theorem 5.1 reveals a general property of $(V_{A},V_{B})$.
-The proof is in Appendix A.
+The single update of current values $v$ with $T$ as $v_{i}^{\prime}(s)=(Tv)_{i}(s)$ can be seen as solving a normal-form (i.e., one-step) Stackelberg game for player $i$ at given state $s\in\mathcal{S}$ where the payoff functions are given by current Q functions
+$q_{i}(s,f_{s},b):=r_{i}(s,f_{s},b)+\gamma_{i}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{i}(s^{\prime})\right]\quad(i\in\{A,B\})$
+for the leader’s mixed strategy $f_{s}$ and the follower’s pure strategy $b$.
+In this context, $R_{B}(s,f_{s},v_{B})$ is the follower’s best response to $f_{s}$, $R_{A}(s,v)$ is the leader’s Stackelberg equilibrium strategy, the pair $(R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))$ is the Stackelberg equilibrium, and $((Tv)_{A}(s),(Tv)_{B}(s))$ is the equilibrium payoffs of the normal-form game for given $s$ and $v$.
+Bucarey et al.
+(2022) proposed a method to construct a stationary equilibrium policy corresponding to the fixed point of the operator, which is called fixed point equilibrium (FPE).
+If there exists a fixed point $V:=(V_{A},V_{B})$ of $T$, the pair of the value functions of the FPE $(\bar{f},\bar{g})$ defined as, for any $s\in\mathcal{S}$,
+$\bar{f}(s)\in R_{A}(s,V),\quad\bar{g}(s)\in R_{B}(s,\bar{f}(s),V_{B})$
+coincides with $V$ (Bucarey et al., 2022).
+Since $T$ is a contraction mapping for $v_{A}$ under certain conditions (Bucarey et al., 2022) [^3] One example is a condition that $r_{A}$ is an affine transformation of $r_{B}$ (i.e., cooperative or competitive with $r_{B}$).
+In general, $T$ does not always converge, which is empirically demonstrated for the aforementioned example game (Figure 1) in Jean-Marie et al.
+(2022)., we can obtain the fixed point asymptotically by repeatedly applying $T$ with arbitrary initial values.
+Let us consider the characteristics of the fixed point $(V_{A},V_{B})$ of the operator $T$ in general situations where the fixed point exists.
+Theorem˜5.1 reveals a general property of $(V_{A},V_{B})$.
+The proof is in Appendix˜A.
+(Appendices are provided as supplementary material.)
 Theorem 5.1.
@@
 Then, if $(TV)_{i}(s)=V_{i}(s)$ holds for all $i\in\{A,B\}$ and for all $s\in\mathcal{S}$, it holds that
-$V_{A}(s)$ $=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)\quad\forall s\in\mathcal{S};$
-$V_{B}(s)$ $=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s)\quad\forall s\in\mathcal{S}.$
+$V_{A}(s)=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)\kern 5.0pt\forall s\in\mathcal{S};\quad V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s)\kern 5.0pt\forall s\in\mathcal{S}.$
 If $V$ is the fixed point, because $\bar{f}=R_{A}(V)$ and $\bar{g}=R_{B}(\bar{f},V_{B})$ , it holds that
@@
 $V_{A}^{\bar{f}\bar{g}}(s)=V_{A}(s)=\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}(s)$ (4)
-for all $s\in\mathcal{S}$ in light of Theorem 5.1.
-Equation 3 shows that the follower’s FPE $\bar{g}=R_{B}(\bar{f},V_{B})$ is the best response to the leader’s FPE $\bar{f}$.
-However, $R_{B}(f,V_{B})$ is not a best response to an arbitrary $f\in\mathcal{W}_{A}$, except when $V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s)~{}\forall s\in\mathcal{S}$.
-In addition, determining any meaning from $R_{B}(f,V_{B})$ is difficult.
-Therefore, $V_{A}^{\bar{f}\bar{g}}$ does not coincides with the SE value function.
-Moreover, there the performance of the FPE is not guaranteed, which is also empirically demonstrated in Bucarey et al.
+for all $s\in\mathcal{S}$ in light of Theorem˜5.1.
+Equation˜3 shows that the follower’s FPE $\bar{g}=R_{B}(\bar{f},V_{B})$ is the best response to the leader’s FPE $\bar{f}$.
+However, $R_{B}(f,V_{B})$ is not the best response to arbitrary $f\in\mathcal{W}_{A}$, except when $V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{fg}(s)~\forall s\in\mathcal{S}$.
+The results of Theorem˜5.1 illuminate that the leader’s performance $V_{A}$ of FPEs is the maximum of the leader’s state value on $f\in\mathcal{W}_{A}$ under $R_{B}(f,V_{B})$, but not the maximum under the follower’s best responses $R_{B}^{*}(f)$ like SE value functions.
+Therefore, the leader’s performance of FPEs at each state can be less than the SE value, which is empirically demonstrated in Bucarey et al.
 (2019).
-5.2 State-dependent Markov policy and SSE Policy
-An attempt to design another operator that has an SE value function as a fixed point is made.
-From this attempt, we derive a novel sufficient condition for the existence of the SSE policy that is also a necessary condition for a special case.
-The game class that is guaranteed to have SSE policies by our sufficient condition includes the game class guaranteed by Bucarey et al.
-(2022).
-First, we define a new policy class, state-dependent Markov policy, in which the whole decision rule for each state is determined depending on the state.
-Definition 5.2 ((Leader’s) State-dependent Markov policy).
-Let $s\in\mathcal{S}$ be a state visited by the agent.
-A decision rule that the agent follows at $s$ is determined depending on $s$ when the agent follows a state-dependent Markov policy.
-In other words, let $\hat{\mathbf{f}}:\mathcal{S}\times\mathcal{S}\to\Delta(\mathcal{A})$ be the agent’s state-dependent Markov policy and $\hat{\mathbf{f}}(\cdot;s)\in\mathcal{W}_{A}$ be the agent’s decision rule at $s$.
-Then, the agent’s action on $s$ is sampled as $a\sim\hat{\mathbf{f}}(s;s)$ and the follower’s best response action is $R_{B}^{*}(s,\hat{\mathbf{f}}(\cdot;s))\in\mathcal{B}$.
-Let $\mathcal{W}_{A}^{\mathcal{S}},\mathcal{W}_{B}^{\mathcal{S}}$ be the sets of state-dependent Markov policies for the leader and the follower, respectively.
-For simplicity, let $\hat{f}_{s}:=\hat{\mathbf{f}}(\cdot;s)$.
-State-dependent Markov policies are an extension of the stationary policy concept, such that $\mathcal{W}_{A}^{\mathcal{S}}\supset\mathcal{W}_{A}$.
-Similar to the stationary policies, we can define the leader’s state value function under the best-response follower as
-$V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$ $:=V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s)$
-$:=\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right],$
-where $R_{B}^{*}(\hat{\mathbf{f}})\in\mathcal{W}_{B}^{\mathcal{S}}$ is the follower’s state-dependent Markov policy in which the decision rule on the visited state $s\in\mathcal{S}$ is $R_{B}^{*}(\hat{f}_{s})\in\mathcal{W}_{B}$, and $\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}[\cdot]$ is an expectation for the trajectory $\{(s_{t},a_{t},b_{t})\}_{t=0}^{\infty}$, which is generated under $(\hat{\mathbf{f}},R_{B}^{*}(\hat{\mathbf{f}}),\hat{p},\rho)$.
-Let the maximization of $V_{A}^{\hat{\mathbf{f}}{\dagger}}$ on $\mathcal{W}_{A}^{\mathcal{S}}$ be defined as
-$\hat{V}_{A}^{*}(s):=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$
-for each $s\in\mathcal{S}$. Subsequently, Theorem 5.3 shows that $\hat{V}^{*}$ is the unique fixed point of the operator $\hat{T}_{A}$ defined in Equation 5. The proof is in Appendix B.
-Theorem 5.3.
-Let $\hat{T}_{A}:\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}$ be defined as
-$(\hat{T}_{A}v)(s):=\sup_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\quad\forall s\in\mathcal{S}.$ (5)
-Then, $\hat{T}_{A}$ is a contraction mapping and its unique fixed point is $\hat{V}_{A}^{*}$.
-We define the set of decision rules $R_{A}^{*}(s)$ for each state $s\in\mathcal{S}$ as
-$R_{A}^{*}(s):=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))\\ +\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (6)
-and we suppose $R_{A}^{*}(s)$ is non-empty for all $s\in\mathcal{S}$.
-Lemma 5.4 provides a way to construct a state-dependent Markov policy by using $R_{A}^{*}$ that realizes $\hat{V}_{A}^{*}$.
-This proof is in Appendix C.
-Lemma 5.4.
-Let $\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}$ be a leader’s state-dependent Markov policy that satisfies $\hat{\mathbf{f}}(\cdot;s)\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$.
-Then, $V_{A}^{\hat{\mathbf{f}}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}\hat{V}_{A}^{*}$ holds.
-Finally, Theorem 5.5 provides a sufficient condition for the existence of SSE policies. The proof is in Appendix D.
-Theorem 5.5.
-Let $R_{A}^{*}(s)$ be a subset of $\mathcal{W}_{A}$ defined in Equation 6 for each $s\in\mathcal{S}$.
-Then, if there exists $\hat{f}\in\mathcal{W}_{A}$ such that $\hat{f}\in\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$, $\hat{f}$ is an SSE policy.
-If we assume the sufficient conditions for the existence of SSE policies in Bucarey et al.
-(2022), $\gamma_{B}=0$ holds, or the transition does not depend on the follower’s action.
-Then, for $f\in\mathcal{W}_{A}$, because $R_{B}^{*}(s,f)$ depends only on $s\in\mathcal{S}$ and the leader’s action distribution $f(s)\in\Delta(\mathcal{A})$ on state $s$, the part to the right of the $\max$ operator in the right-hand side of Equation 6 depends only on $f(s)\in\Delta(\mathcal{A})$ and not on all $f(s^{\prime})\in\Delta(\mathcal{A})$ such that $s^{\prime}\neq s$.
-Thus, $R_{A}^{*}(s)$ is a restricted set with respect only to the action distribution $f(s)\in\Delta(\mathcal{A})$, which shows that there exists $f^{\prime}\in\mathcal{W}_{A}$ such that $f^{\prime}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$.
-Then, the existence of SSE policies is guaranteed by Theorem 5.5.
-This means that the game class whose existence of SSE policies is guaranteed by Theorem 5.5 includes the game class whose existence is guaranteed by Bucarey et al.
-(2022).
-Rather than considering a myopic follower, if we consider the situation that the leader’s discount rate $\gamma_{A}=0$, the sufficient condition in Theorem 5.5 becomes a necessary condition.
-Its proof is provided in Appendix E.
-Corollary 5.6.
-If $\gamma_{A}=0$, then an SSE policy exists if and only if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is non-empty.
-Even if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is empty, we can construct stationary policies utilizing $R_{A}^{*}$, which is asymptotically obtained using $\hat{T}_{A}$ in light of Theorem 5.3.
-A possible approach is to choose $f$ contained in $R_{A}^{*}(s)$ for most of the states or to extract one non-dominated by stationary policies in $R_{A}^{*}$.
-However, the performance is not guaranteed for the stationary policies obtained as aforementioned.
-This is because any stationary policy not included in $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ does not achieve $\hat{V}_{A}^{*}$, and its performance may be significantly worse than $\hat{V}_{A}^{*}$ as $R_{B}^{*}(f)$; hence, $\hat{V}_{A}^{*}$ is not necessarily continuous with respect to $f$.
-6 Policy Iteration for Pareto-Optimal Policies
+In addition, it is difficult to determine any meaning from $R_{B}(f,V_{B})$ for arbitrary $f\in\mathcal{W}_{A}$, which makes the performance of FPEs unpredictable.
+Although Bucarey et al.
+(2022) derives the sufficient condition for SSEs to exist and for FPEs to coincide with the SSEs, this condition is not compatible with our problem setting or can be a strong assumption.
+The sufficient condition is equivalent to the union of two conditions (Bucarey et al., 2022).
+The first is that the follower’s discount rate $\gamma_{B}=0$, which is a condition on the follower’s learning algorithm, but we admit no other assumptions on the follower than the best responsivity.
+The second is that the transition function does not depend on the follower’s actions, which can limit the applications of their approach.
+In the end, finding FPEs does not guarantee the leader’s performance in our setting.
+To overcome the problem of SSGs that SSEs do not always exist, in the next section, we propose an algorithm that guarantees monotone improvement in the leader’s performance toward alternative equilibria, Pareto-optimal policies.
+6 Pareto-Optimal Policy Iteration
 SSE policies do not always exist in general-sum stochastic games.
-Therefore, it is desired that an algorithm converges to an SSE policy if it exists, and to a policy that is “reasonable” provided an SSE policy does not exist.
+Therefore, it is desired that an algorithm converges to an SSE policy if it exists and to a policy that is “reasonable” provided an SSE policy does not exist.
 First, we introduce the notion of Pareto-optimality as the reasonable target.
-A policy that approximates a Pareto-optimal value function with an arbitrary precision always exists and it admits an SSE policy if an SSE policy exists.
-Then, we derive a policy improvement theorem for SSGs.
-Based on the policy improvement theorem for SSGs, we design an algorithm that monotonically improves the policy toward Pareto-optimal policies.
+A policy that approximates a Pareto-optimal value function with an arbitrary precision always exists, and it admits an SSE policy if an SSE policy exists.
+Then, we derive a policy improvement theorem for SSGs under best-response followers.
+Based on the policy improvement theorem, we design an algorithm that monotonically improves the policy toward Pareto-optimal policies.
+Notation
+We introduce the notation for the dominance relation.
+Let $v,v^{\prime}\in\mathcal{F}_{\mathcal{S}}$.
+If $v(s)=v^{\prime}(s)$ for all $s\in\mathcal{S}$, we express $v\stackrel{{\scriptstyle s}}{{=}}v^{\prime}$.
+If $v(s)\geq v^{\prime}(s)$ for all $s\in\mathcal{S}$, then $v$ weakly dominates $v^{\prime}$ and $v\succcurlyeq v^{\prime}$.
+If $v\succcurlyeq v^{\prime}$ and $\not\stackrel{{\scriptstyle s}}{{=}}v^{\prime}, vv$ strictly dominates $v^{\prime}$ and $v\succ v^{\prime}$.
 6.1 Pareto-Optimal Policy
@@
 Meanwhile, there is no stationary policy that realizes a value function better than a PO value function.
-Moreover, an arbitrary PO policy is an SSE policy if there exist SSE policies, as shown in Proposition 6.2.
+Moreover, an arbitrary PO policy is an SSE policy if there exist SSE policies, as shown in Proposition˜6.2.
 Therefore, PO value functions are a reasonable alternative to the SE value function.
-The proof is presented in Appendix F.
+The proof is presented in Appendix˜B.
 Proposition 6.2.
@@
 Consequently, the SSE policies exist if and only if $\mathcal{PV}$ is a singleton and $\mathcal{PV}\subseteq\mathcal{V}$.
+Figure 2: Pareto optimal value functions in the two-state value function space.
+Thick blue line: Pareto optimal value functions.
+Red shaded area (the upper right of the dot in the box): the area of value functions dominating the value function represented by the black dot, which is empty for PO value functions.
 6.2 Policy Improvement Theorem
 We derive three important lemmas based on which we design the proposed approach.
-Lemma 6.3 is an extension of the policy improvement theorem Sutton & Barto (2018) in single-agent MDPs.
+Theorem˜6.3 is an extension of the policy improvement theorem (Sutton and Barto, 2018) in single-agent MDPs.
 With this theorem, we can confirm if a policy $f$ is better than a baseline $f^{\prime}$ without computing the value function for $f$.
-The proof is shown in Appendix G.
-Lemma 6.3.
+The proof is shown in Appendix˜C.
+The key idea behind its proof is to express the follower’s best response in terms of the leader’s policy and then apply reasoning analogous to the policy improvement theorem for single-agent MDPs.
+Theorem 6.3.
 Let $f\in\mathcal{W}_{A},f^{\prime}\in\mathcal{W}_{A}$ be stationary policies, and a Q-function $Q_{A}^{f^{\prime}{\dagger}}(s,f):\mathcal{S}\times\mathcal{W}_{A}\to\mathbb{R}$ is expressed as
@@
 where $\mathbb{E}_{s^{\prime}}$ is considered under $p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))$. Then, the following conditions hold:
-  • (a)
-    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\succcurlyeq Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}$;
-  • (b)
-    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}$;
-  • (c)
-    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}\iff\\ V_{A}^{f{\dagger}}-V_{A}^{f^{\prime}{\dagger}}(s)\succ\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$.
-A necessary condition for PO policies is provided below. It shows that any improvement by Lemma 6.3 (a) is not possible from PO policies. This proof is in Appendix H.
-Lemma 6.4 (Necessary condition for PO policies).
+$(a)\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\succcurlyeq Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}};$
+$(b)\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}};$
+$(c)\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}-V_{A}^{f^{\prime}{\dagger}}\succ\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$
+A necessary condition for PO policies is given below, which shows that no further performance improvement can be made by Theorem˜6.3 (a) from PO policies. This proof is in Appendix˜D.
+Theorem 6.4 (Necessary condition for PO policies).
 If $f^{\prime}\in\mathcal{W}_{A}$ is a PO policy, it holds that for any $f\in\mathcal{W}_{A}$
 $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}.$ (5)
-The following result shows a sufficient condition for PO policies. This proof is presented in Appendix I.
-Lemma 6.5 (Sufficient condition for PO policies).
+The following result shows a sufficient condition for PO policies. This proof is presented in Appendix˜E.
+Theorem 6.5 (Sufficient condition for PO policies).
 $f^{\prime}\in\mathcal{W}_{A}$ is a PO policy if, for any $f\in\mathcal{W}_{A}$, either of the following two conditions hold:
-  • i)
-    $Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$, $\exists s\in\mathcal{S}$, where $\delta(f,f^{\prime})=\max\limits_{s\in\mathcal{S}}\mathbb{E}_{s^{\prime}}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ and the expectation is considered for $s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))$;
-  • ii)
-    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}$.
+$(\mathrm{i})\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}};$
+$(\mathrm{ii})\quad\exists s\in\mathcal{S},Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime}),$
+where $\delta(f,f^{\prime}):=\max\limits_{s\in\mathcal{S}}\mathbb{E}_{s^{\prime}\sim p(\cdot|s,f(s),R_{B}^{*}(s,f))}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$
+When $\gamma_{A}=0$, Theorem˜6.5 implies that no further policy update by Theorem˜6.3 (a) with strict improvement is a sufficient condition for PO policies.
+Combined with Theorem˜6.4, $f^{\prime}$ is a PO policy if and only if there is no room for policy improvement when $\gamma_{A}=0$.
 6.3 Policy Iteration for PO Policies
-We propose an approach that monotonically improves a policy toward PO policies and is guaranteed to converge.
+We propose an approach that monotonically improves policy toward PO policies and is guaranteed to converge.
 PO value functions are not always unique.
 To guide the algorithm, we introduce scalarization.
-Given a stationary policy $f\in\mathcal{W}_{A}$, we maximize a Pareto-compliant scalarization $\mathcal{L}$ of the state value on $f$, namely, $\mathcal{L}[V_{A}^{f{\dagger}}]$, where the Pareto-compliant scalarization555 For example, a weighted sum scalarization $\mathcal{L}[v]=\sum_{s\in\mathcal{S}}\alpha_{s}v(s)$ is a Pareto-compliant scalarization for $\boldsymbol{\alpha}:=\{\alpha_{s}\in\mathbb{R}_{>0}\}_{s\in\mathcal{S}}$ such that $\sum_{s\in\mathcal{S}}\alpha_{s}=1$.
+Given a stationary policy $f\in\mathcal{W}_{A}$, we maximize a Pareto-compliant scalarization $\mathcal{L}$ of the state value on $f$, namely, $\mathcal{L}[V_{A}^{f{\dagger}}]$, where the Pareto-compliant scalarization [^4] For example, a weighted sum scalarization $\mathcal{L}[v]=\sum_{s\in\mathcal{S}}\alpha_{s}v(s)$ is a Pareto-compliant scalarization for $\boldsymbol{\alpha}:=\{\alpha_{s}\in\mathbb{R}_{>0}\}_{s\in\mathcal{S}}$ such that $\sum_{s\in\mathcal{S}}\alpha_{s}=1$.
 If we choose $\alpha$ such that $\alpha_{s}=\rho(s)$ for all $s\in\mathcal{S}$, then $\mathcal{L}[v]$ coincides with the expected return under the initial state distribution $\rho$.
@@
 $\mathcal{W}_{\succcurlyeq}(f_{t}):=\left\{f\in\mathcal{W}_{A}~|~Q_{A}^{f_{t}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f_{t}{\dagger}}\right\}.$ (6)
-Selecting the next policy $f_{t+1}$ from $\mathcal{W}_{\succcurlyeq}(f_{t})$ guarantees $\mathcal{L}[V_{A}^{f_{t+1}{\dagger}}]\geq\mathcal{L}[V_{A}^{f_{t}{\dagger}}]$ in light of Lemma 6.3 (a).
-Ideally,666 In practice, $\operatornamewithlimits{argmax}$ is not necessary, and sometimes it does not exist or is not computable.
-An alternative approach is to select a policy $f_{t+1}$ from a set $\mathcal{W}_{1-\epsilon}(f_{t})$ of policies that are $\epsilon$-optimal in terms of $\mathcal{L}$, such as $\mathcal{W}_{1-\epsilon}(f_{t}):=\bigg{\{}f\in\mathcal{W}_{\succcurlyeq}(f_{t})\mid\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\\ \geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\bigg{(}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\bigg{)}\bigg{\}}$ (9) for some $\epsilon\in[0,1)$.
+Selecting the next policy $f_{t+1}$ from $\mathcal{W}_{\succcurlyeq}(f_{t})$ guarantees $\mathcal{L}[V_{A}^{f_{t+1}{\dagger}}]\geq\mathcal{L}[V_{A}^{f_{t}{\dagger}}]$ in light of Theorem˜6.3 (a).
+Ideally, [^5] In practice, $\operatornamewithlimits{argmax}$ is not necessary, and sometimes it does not exist or is not computable.
+An alternative approach is to select a policy $f_{t+1}$ from a set $\mathcal{W}_{1-\epsilon}(f_{t})$ of $\epsilon$-optimal policies in terms of $\mathcal{L}$, such as $\mathcal{W}_{1-\epsilon}(f_{t}):=\bigg\{f\in\mathcal{W}_{\succcurlyeq}(f_{t})\mid\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\\ \geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\bigg(\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\bigg)\bigg\}$ (7) for some $\epsilon\in[0,1)$.
 Notably, $\mathcal{W}(f_{t})=\mathcal{W}_{1}(f_{t})$.
@@
 $f_{t+1}\in\mathcal{W}(f_{t}):=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right].$ (8)
-Theorem 6.6 guarantees that the proposed algorithm always converges and that the limit satisfies the necessary condition for PO policies (i.e., $v_{\infty}\in\partial\mathcal{V}$).
-The proof is shown in Appendix J.
+Theorem˜6.6 guarantees that the proposed algorithm always converges and that the limit satisfies the necessary condition for PO policies (i.e., $v_{\infty}\in\partial\mathcal{V}$).
+The proof is shown in Appendix˜F.
 Theorem 6.6.
-Let $f_{0}\in\mathcal{W}_{A}$ be an arbitrary initial policy and a policy sequence $\{f_{t}\in\mathcal{W}(f_{t-1})\}_{t=1}^{\infty}$ be obtained by Equation 9.
+Let $f_{0}\in\mathcal{W}_{A}$ be an arbitrary initial policy and a policy sequence $\{f_{t}\in\mathcal{W}(f_{t-1})\}_{t=1}^{\infty}$ be obtained by Equation˜7.
 Let $\{v_{t}=V_{A}^{f_{t}{\dagger}}\}_{t=0}^{\infty}$ be the corresponding value functions.
 Then, the following statements hold:
-  • a)
+  • (a)
     $\{v_{t}\}_{t=0}^{\infty}$ monotonically increases in the sense that $v_{t+1}\succcurlyeq v_{t}$ and converges to $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$;
-  • b)
+  • (b)
     $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$ if and only if $v_{t}\stackrel{{\scriptstyle s}}{{=}}v_{\infty}$;
-  • c)
-    $v_{\infty}\in\partial\mathcal{V}$ and for all $v\in\mathcal{V}$,
-    $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$;
-  • d)
+  • (c)
+    $v_{\infty}\in\partial\mathcal{V}$ and $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$ for all $v\in\mathcal{V}$;
+  • (d)
     $v_{\infty}\in\mathcal{PV}$ if $\gamma_{A}=0$.
-As shown in the proof (Appendix J), the algorithm ceases to improve if and only if $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$.
+As shown in the proof (Appendix˜F), the algorithm ceases to improve if and only if $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$.
 The algorithm terminates when it holds and returns $f_{t}$.
-The whole proposed algorithm summarizing the above is shown in Algorithm 1.
-Algorithm 1 Policy Iteration for PO Policies (PIPOP)
+The entire proposed algorithm is shown in Algorithm˜1.
+Algorithm 1 Pareto-Optimal Policy Iteration
 0: Maximum number of iterations $M$.
 1: Randomly initialize $f_{0}\in\mathcal{W}_{A}$.
-2:  for $t=0$ to $M-1$ do
+2:for $t=0$ to M1M-1 do
 3:  Compute $V_{A}^{f_{t}{\dagger}}$ by repeatedly applying $T_{A}^{f_{t}R_{B}^{*}(f_{t})}$
-4:     Compute $\mathcal{W}_{\succcurlyeq}(f_{t})$ in Equation 8
+4:  Compute $\mathcal{W}_{\succcurlyeq}(f_{t})$ in Equation˜6
 5:  if $Q_{A}^{f_{t}{\dagger}}(s,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$ then
@@
 7:  else
-8:        Compute $\mathcal{W}(f_{t})$ in Equation 10 or Equation 9
+8:   Compute $\mathcal{W}(f_{t})$ in Equation˜8 or Equation˜7
 9:   Randomly sample $f_{t+1}$ from $\mathcal{W}(f_{t})$
@@
 12: A stationary policy $f^{*}$
-The convergence guarantee and the monotone improvement of Algorithm 1 (Statements (a) and (b) in Theorem 6.6) show the advantages of the proposed approach.
-However, compared with the policy iteration approach in single-agent MDPs, where the convergence toward the optimal policy is guaranteed, the fixed point of Algorithm 1 is not assured to be PO value functions.
-Rather, Theorem 6.6 (c) guarantees that the limit is at the boundary $\partial\mathcal{V}$, including all the PO value functions.
-In a special case of $\gamma_{A}=0$, the limit is guaranteed to be a PO value function (Theorem 6.6 (d)).
-Furthermore, there exists an algorithmic difference to the policy iteration for the single-agent MDP.
+The convergence guarantee and the monotone improvement of Algorithm˜1 (Statements (a) and (b) in Theorem˜6.6) show the advantages of the proposed approach.
+However, compared with the policy iteration approach in single-agent MDPs, where the convergence to the optimal policy is guaranteed, the fixed point of Algorithm˜1 is not assured to be PO value functions.
+Rather, Theorem˜6.6 (c) guarantees that the limit is at the boundary $\partial\mathcal{V}$, including all the PO value functions.
+In a special case of $\gamma_{A}=0$, the limit is guaranteed to be a PO value function (Theorem˜6.6 (d)).
+The visualization of (c) and (d) is in Figure˜3.
+This condition, the leader’s discount rate $\gamma_{A}=0$, is a condition on the leader’s learning process, meaning Algorithm˜1 does not need any assumptions on the follower’s learning process to guarantee the convergence and the leader’s “reasonable” performance.
+This property allows us to apply this algorithm to various applications modeled by decentralized SSGs in which we are the leader.
+We give an example of such applications in Appendix˜J.
+There is another algorithmic difference from the policy iteration for the single-agent MDP.
 In single-agent MDPs, a state-wise maximization of the action value function of $f_{t}$, $a_{s}\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}Q^{f_{t}}(s,a)$ is performed to construct the next (deterministic) policy.
-This is invalid in SSGs because such a policy does not necessarily satisfy the condition of Lemma 6.3 (a) and may not improve the values.
+This is invalid in SSGs because such a policy does not necessarily satisfy the condition of Theorem˜6.3 (a) and may not improve the values.
 Moreover, the optimal policy is not necessarily a deterministic one.
-Because of these differences, Equation 10 cannot be simplified as the state-wise maximization.
-This can be a limitation from a practical viewpoint when Equation 10 (or Equation 9) is intractable.
-In practice, it is helpful to restart the policy iteration process from other initial policies to obtain a better policy unless the obtained policy satisfies the sufficient condition Lemma 6.5.
-In this situation, one can randomly pick an initial policy or select a policy $f\in\mathcal{W}(f_{t-1})$ that is not yet explored.
-Because $f\in\mathcal{W}(f_{t-1})$ is already improved from the previous initial policy, we can expect such an initialization to be more efficient.
-The detail of the proposed approach with the restart strategy is provided in Appendix K.
+Because of these differences, Equation˜8 cannot be simplified to the state-wise maximization.
+This can be a limitation from a practical viewpoint when Equation˜8 (or Equation˜7) is intractable due to the large search space of $\mathcal{W}_{A}$.
+We propose a practical strategy of splitting the policy space to find the next policy efficiently, avoiding exhaustive search over the entire space $\mathcal{W}_{A}$.
+Its details are provided in Appendix˜G
+Figure 3: Visualization of Theorem˜6.6 (c) and (d).
+For $v_{\infty}$, there must not exist a value function in the red-shaded cone representing the area of $v$ violating the inequality in (c).
+The upper and lower slopes of the cone are $1/\gamma_{A}$ and $\gamma_{A}$, respectively.
+$v_{\infty}$ can be on the non-Pareto boundary (grey solid line) when $\gamma_{A}>0$ (left), but $v_{\infty}$ is on $\mathcal{PV}$ when $\gamma_{A}=0$ (right) because the cone equals the area of dominating value functions.
 7 Limitation and Future Work
-This study proposed a novel algorithm with a convergence guarantee in general-sum SSGs.
-We introduced the notion of the Pareto-optimal value function and developed an algorithm that monotonically improves a policy toward the Pareto-optimal value functions.
-While existing methods have little guarantee, especially in games where the SSEs do not exist, our proposed approach can monotonically improve the leader’s performance in general, which is of great importance in practice.
-A potential application of the proposed approach is the intervention to an adaptive or learning follower agent as described in Appendix M, where the follower is not usually myopic; hence the existence of the SSEs is not guaranteed.
-On such an application, the proposed approach is advantageous to the existing approaches discussed in this paper.
-However, there is room for improvement both from theoretical and practical viewpoints. Further research must focus on addressing the limitations of the current work listed below.
+This study proposed a novel algorithm with a convergence guarantee in two-player general-sum SSGs under best-response followers.
+We introduced the notion of the Pareto-optimal value function to target it even if there are no SSEs, and developed an algorithm that monotonically improves the policy toward the Pareto front.
+While existing methods have little guarantee under best-response followers, especially in games where the SSEs do not exist, our proposed approach can monotonically improve the leader’s performance and guarantees its limit to be Pareto optimal in state values when the leader is myopic.
+To the best of our knowledge, this is the first theoretical guarantee in general-sum myopic-leader SSGs.
+However, there is room for improvement both from theoretical and practical viewpoints.
+Further research must focus on addressing the limitations of the current work listed below.
 First, the Pareto-optimality of the proposed algorithm is not guaranteed.
@@
 The development of a restart strategy to satisfy such a sufficient condition is desired.
-Second, a computationally efficient and algorithmic update rule to select the next policy (cf., Equation 10 or Equation 9) is required.
-Practically, such an update similar to the policy iteration in single-agent MDPs, is desired.
-Finally, our approach requires the knowledge of the follower’s best response and transition probability as in the previous studies.
+Second, a computationally efficient update rule to select the next policy (cf.
+Equation˜8 or Equation˜7) is required.
+Practically, a state-wise update similar to the policy iteration in single-agent MDPs is desired.
+Finally, our approach requires knowledge of the follower’s best response and transition probability, as in the previous studies.
 A sample approximation of the algorithm by reinforcement learning is desired to widen the applicability of the proposed approach.
-Appendix A Proof of Theorem 5.1
-Proof.
-First, letting $V_{B}$ be fixed, we prove that $(TV)_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}$ implies $V_{A}\stackrel{{\scriptstyle s}}{{=}}\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
+Appendix A Proof of Theorem˜5.1
+Proof A.1.
+First, we prove that $(TV)_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}$ implies $V_{A}\stackrel{{\scriptstyle s}}{{=}}\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$ for a fixed $V_{B}$.
 Suppose that $V_{A}\stackrel{{\scriptstyle s}}{{=}}(TV)_{A}$. Given a policy $f\in\mathcal{W}_{A}$, let $f_{s}=f(s)$ for each $s\in\mathcal{S}$. Then, we obtain for all $s\in\mathcal{S}$
@@
 $V_{A}(s)$ $\geq r_{A}(s,f(s),R_{B}(s,f(s),V_{B}))+\gamma_{A}\mathbb{E}_{s_{1}}\left[V_{A}(s_{1})\right]$ (10a)
-=$\end{multlined}r_{A}(s,f(s),R_{B}(s,f(s),V_{B}))+\gamma_{A}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},f(s_{1}),R_{B}(s_{1},f(s_{1}),V_{B}))+\gamma_{A}\mathbb{E}_{s_{2}}\left[V_{A}(s_{2})\right]\Bigl{]}$ (12c)
 […]
-$=\mathbb{E}^{fR_{B}(f,V_{B})}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}(f,V_{B})}\Bigl{[}V_{A}(s_{n})\Bigl{|}s_{0}=s\Bigl{]}.$ (12d)
-In the limit for $n\to\infty$, the first term of the right-hand side of Equation 12d converges to $V_{A}^{fR_{B}(f,V_{B})}(s)$, while the second term converges to 0777Since $\hat{r}_{A}$ is bounded and $\gamma_{A}\in[0,1)$, the inside of the expectation of Equation 12d is always bounded, which enables the exchange of the extreme and the expectation.
-Therefore, the same exchanges are performed in Equation 16d and Equation 37e..
+$\geq\cdots$
+[…]
+In the limit for $n\to\infty$, the first term of the right-hand side of Equation˜10f converges to $V_{A}^{fR_{B}(f,V_{B})}(s)$, while the second term converges to 0666Since $\hat{r}_{A}$ is bounded and $\gamma_{A}\in[0,1)$, the inside of the expectation of Equation 10f is always bounded, which enables the exchange of the extreme and the expectation.
+Therefore, the same exchanges are performed in Equation 14e and Equation 19e..
 As the choice of $f$ is arbitrary, we obtain $V_{A}\succcurlyeq\sup_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
@@
 $=r_{A}(s,f_{s}^{*},R_{B}(s,f_{s}^{*},V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s}^{*},R_{B}(s,f_{s}^{*},V_{B}))}\left[V_{A}(s^{\prime})\right]$ (11c)
-$=r_{A}(s,f^{*}(s),R_{B}(s,f^{*}(s),V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f^{*}(s),R_{B}(s,f^{*}(s),V_{B}))}\left[V_{A}(s^{\prime})\right].$ (13d)
+[…]
 Therefore, with the repeated application of this equality, analogously to the above argument, we obtain $V_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{*}R_{B}(f^{*},V_{B})}$.
@@
 $V_{B}(s)\geq r_{B}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),g(s))}\left[V_{B}(s^{\prime})\right]$ (13)
-for all $s\in\mathcal{S}$ and for all $g\in\mathcal{W}_{B}^{d}$. By recursively applying Equation 15 for $V_{B}$ on the right-hand side, we have for all $s\in\mathcal{S}$
+for all $s\in\mathcal{S}$ and for all $g\in\mathcal{W}_{B}^{d}$. By recursively applying Equation˜13 for $V_{B}$ on the right-hand side, we have for all $s\in\mathcal{S}$
 $V_{B}(s)$ $\geq r_{A}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s_{1}}\left[V_{B}(s_{1})\right]$ (14a)
-$\end{multlined}r_{A}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},R_{A}(s_{1},V),g(s_{1}))+\gamma_{B}\mathbb{E}_{s_{2}}\left[V_{B}(s_{2})\right]\Bigl{]}$ (16c)
 […]
-$\geq\mathbb{E}^{R_{A}(V)g}\left[\sum_{t=0}^{n-1}\gamma_{B}^{t}\hat{r}_{B}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]+\gamma_{B}^{n}\mathbb{E}^{R_{A}(V)g}\Bigl{[}V_{B}(s_{n})\Bigl{|}s_{0}=s\Bigl{]}.$ (16d)
-In the limit for $n\to\infty$, the first term of the right-hand side of Equation 16d converges to $V_{B}^{R_{A}(V)g}(s)$, and the second term converges to 0.
+$\geq\cdots$
+[…]
+In the limit for $n\to\infty$, the first term of the right-hand side of Equation˜14e converges to $V_{B}^{R_{A}(V)g}(s)$, and the second term converges to 0.
 Therefore, $V_{B}(s)\geq V_{B}^{R_{A}(V)g}(s)$ holds for all $s\in\mathcal{S}$.
@@
 $V_{B}(s)=r_{B}(s,R_{A}(s,V),g^{*}(s))+\gamma_{B}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),g^{*}(s))}\left[V_{B}(s^{\prime})\right].$ (17)
-With the repeated application of this equality, analogous to the above argument, we obtain $V_{B}\stackrel{{\scriptstyle s}}{{=}}V_{B}^{R_{A}(V)g^{*}}$. ∎
-Appendix B Proof of Theorem 5.3
-Proof.
-First, we prove that $\hat{T}_{A}$ is a contraction mapping; hence it has a unique fixed point.
-For an arbitrary $(v,u)\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$ and $s\in\mathcal{S}$, we obtain
-$(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)=\sup_{f\in\mathcal{W}_{A}}\left\{r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\right\}\\ -\sup_{f\in\mathcal{W}_{A}}\left\{r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[u(s^{\prime})\right]\right\}$ (20)
-from the definition. Let the first term on the right-hand side of the above equation be $v^{*}(s)$. Then, for an arbitrary $\delta>0$, there exists $f\in\mathcal{W}_{A}$ such that
-$r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[v(s^{\prime})\right]\geq v^{*}(s)-\delta.$ (21)
-Let $\delta=\frac{1-\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}$ and $f_{v}$ satisfy the above condition. Then,
-$(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)$ $\leq v^{*}(s)-\left\{r_{A}(s,f_{v}(s),R_{B}^{*}(s,f_{v}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{v}(s),R_{B}^{*}(s,f_{v}))}\left[u(s^{\prime})\right]\right\}$ (22a)
-[…]
-$=\delta+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{v}(s),R_{B}^{*}(s,f_{v}))}\left[v(s^{\prime})-u(s^{\prime})\right]$ (22e)
-$\leq\delta+\gamma_{A}\lVert v(s)-u(s)\rVert_{\infty}$ (22f)
-$=\frac{1+\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}.$ (22g)
-Analogously, we have
-$(\hat{T}_{A}u)(s)-(\hat{T}_{A}v)(s)\leq\frac{1+\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}.$ (23)
-Therefore, since $s$ is arbitrary, we have
-$\lVert(\hat{T}_{A}v)(s)-(\hat{T}_{A}u)(s)\rVert_{\infty}\leq\frac{1-\gamma_{A}}{2}\lVert v(s)-u(s)\rVert_{\infty}$ (24)
-for any $(v,u)\in\mathcal{F}_{\mathcal{S}}\times\mathcal{F}_{\mathcal{S}}$.
-This implies that $\hat{T}_{A}$ is a contraction mapping.
-According to the Banach fixed-point theorem, $\hat{T}_{A}$ has a unique fixed point.
-Subsequently, we prove that $\hat{V}_{A}^{*}$ is the fixed point of $\hat{T}_{A}$. From the definition, we obtain
-$\hat{V}_{A}^{*}(s)$ $=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)$ (25a)
-$=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[V_{A}^{\hat{\mathbf{f}}{\dagger}}(s^{\prime})\right]$ (25b)
-$\leq\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (25c)
-$=\sup_{\hat{f}_{s}\in\mathcal{W}_{A}}r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (25d)
-$=(\hat{T}_{A}\hat{V}_{A}^{*})(s)$ (25e)
-for all $s\in\mathcal{S}$.
-We show that the equality in Equation 25c holds for all $s\in\mathcal{S}$.
-Suppose that there exists $s_{0}\in\mathcal{S}$ such that the above inequality strictly holds.
-Let the discrepancy between the left-hand side and the right-hand side of the above inequality at state $s_{0}$ be $2\epsilon$.
-Then, there must exists $\hat{\mathbf{f}}\in\mathcal{W}_{\mathcal{A}}^{\mathcal{S}}$ such that
-$\hat{V}_{A}^{*}(s)$ $\leq r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]-\epsilon\quad\text{for }s=s_{0}$ (26)
-$\hat{V}_{A}^{*}(s)$ $\leq r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]+\epsilon(1-\gamma_{A})\quad\text{for }s\neq s_{0}$ (27)
-holds, where $\hat{f}_{s}$ is the policy derived from $\hat{\mathbf{f}}$ at $s$. Then, starting $s_{0}$ with Equation 26 and repeatedly applying Equation 27, we obtain
-$\hat{V}_{A}^{*}(s_{0})$
-$\leq r_{A}(s_{0},\hat{f}_{s_{0}}(s_{0}),R_{B}(s_{0},\hat{f}_{s_{0}})+\gamma_{A}\mathbb{E}_{s_{1}}\left[\hat{V}_{A}^{*}(s_{1})\right]-\epsilon$ (28a)
-$\end{multlined}r_{A}(s_{0},\hat{f}_{s_{0}}(s_{0}),R_{B}(s_{0},\hat{f}_{s_{0}}))+\gamma_{A}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},\hat{f}_{s_{1}}(s_{1}),R_{B}(s_{1},\hat{f}_{s_{1}}))+\gamma_{A}\mathbb{E}_{s_{2}}\left[\hat{V}_{A}^{*}(s_{2})\right]\Bigl{]}-\epsilon+\epsilon\gamma_{A}(1-\gamma_{A})$ (28c)
-[…]
-$\leq\mathbb{E}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\left[\sum_{t=0}^{\infty}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]-\epsilon+\epsilon(1-\gamma_{A})\sum_{t=0}^{\infty}\gamma_{A}^{t+1}$ (28d)
-$=V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)-\epsilon(1-\gamma_{A}).$ (28e)
-That is, $V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)>\hat{V}_{A}^{*}(s_{0})$, contradicting $\hat{V}_{A}^{*}\stackrel{{\scriptstyle s}}{{=}}\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}$.
-Therefore, the equality must hold in Equation 25c for all $s\in\mathcal{S}$.
-This implies
-$\hat{V}_{A}^{*}(s)=\max_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]=(\hat{T}_{A}\hat{V}_{A}^{*})(s).$ (29)
-This completes the proof. ∎
-Appendix C Proof of Lemma 5.4
-Proof.
-Let $\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}$ satisfy $\hat{f}_{s}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$.
-When the follower follows the best response against $\hat{\mathbf{f}}$, from the definition of $R_{A}^{*}$, we have
-$r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]\\ =\max_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (30)
-for all $s\in\mathcal{S}$. Therefore, from Equation 29, it holds that
-$\hat{V}_{A}^{*}(s)=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[\hat{V}_{A}^{*}(s^{\prime})\right]$ (31)
-for all $s\in\mathcal{S}$. Let an operator $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}:\mathcal{F}_{\mathcal{S}}\to\mathcal{F}_{\mathcal{S}}$ be defined as
-$(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}v)(s)=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[v(s^{\prime})\right],$ (32)
-then $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$ is a contraction mapping, and as it holds that
-$V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s)$ $=r_{A}(s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,\hat{f}_{s}(s),R_{B}^{*}(s,\hat{f}_{s}))}\left[V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}(s^{\prime})\right]$ (33a)
-$=(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})})(s)$ (33b)
-for all $s\in\mathcal{S}$ from the definition of $V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$, $V_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$ is the unique fixed point of $T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}$.
-From Equation 31, since $\hat{V}_{A}^{*}$ satisfies $\hat{V}_{A}^{*}(s)=(T_{A}^{\hat{\mathbf{f}}R_{B}^{*}(\hat{\mathbf{f}})}\hat{V}_{A}^{*})(s)$ for all $s\in\mathcal{S}$, it holds that $V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)=\hat{V}_{A}^{*}(s)$ for all $s\in\mathcal{S}$.
-∎
-Appendix D Proof of Theorem 5.5
-Proof.
-From Lemma 5.4, it follows that if there exists $f^{\prime}\in\mathcal{W}_{A}$ such that $\hat{f}\in\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$, it holds that $V_{A}^{\hat{f}{\dagger}}(s)=\hat{V}_{A}^{*}(s)$ for all $s\in\mathcal{S}$.
-Then, because it holds that
-$\hat{V}_{A}^{*}(s)=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)\geq\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ (34)
-from $\mathcal{W}_{A}^{\mathcal{S}}\supset\mathcal{W}_{A}$, it holds that
-$\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)\geq V_{A}^{\hat{f}{\dagger}}(s)=\sup_{\hat{\mathbf{f}}\in\mathcal{W}_{A}^{\mathcal{S}}}V_{A}^{\hat{\mathbf{f}}{\dagger}}(s)\geq\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ (35)
-for all $s\in\mathcal{S}$. Therefore, it holds that $V_{A}^{\hat{f}{\dagger}}(s)=\sup_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$. ∎
-Appendix E Proof of Corollary 5.6
-Proof.
-From Theorem 5.5, if $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is non-empty, then a SSE policy exists.
-If $\hat{f}\in\mathcal{W}_{A}$ is a SSE policy, it is holds that $\hat{f}\in\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)$ for all $s\in\mathcal{S}$ from the definition.
-If $\gamma_{A}=0$, because
-$R_{A}^{*}(s)=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}V_{A}^{f{\dagger}}(s)=\operatornamewithlimits{argmax}_{f\in\mathcal{W}_{A}}r_{A}(s,f(s),R_{B}^{*}(s,f))$
-holds, $\hat{f}\in R_{A}^{*}(s)$ for all $s\in\mathcal{S}$. It implies that $\bigcap_{s\in\mathcal{S}}R_{A}^{*}(s)$ is nonempty. ∎
-Appendix F Proof of Proposition 6.2
-Proof.
+With the repeated application of this equality, analogous to the above argument, we obtain $V_{B}\stackrel{{\scriptstyle s}}{{=}}V_{B}^{R_{A}(V)g^{*}}$.
+Appendix B Proof of Proposition˜6.2
+Proof B.1.
 First, suppose that $V_{A}^{*}\in\mathcal{PV}$.
@@
 Hence, $\mathcal{PV}=\{V_{A}^{*}\}$.
-[…]
+Subsequently, suppose that $\mathcal{PV}=\{v^{*}\}$ is a singleton.
+Then, $v^{*}\succ v$ for all $v\in\mathcal{V}\setminus\mathcal{PV}$ because, for all $v\in\mathcal{V}\setminus\mathcal{PV}$, there exists $\hat{v}\in\mathcal{PV}$ such that $\hat{v}\succ v$.
+Suppose that $v^{*}\not\stackrel{{\scriptstyle s}}{{=}}V_{A}^{*}$.
+Then, for some $s\in\mathcal{S}$, $v^{*}(s)<\sup_{f\in\mathcal{W}}V_{A}^{f{\dagger}}(s)$ holds.
+It implies that there exists $v\in\mathcal{V}\setminus\mathcal{PV}$ such that $v^{*}(s)<v(s)$ for some $s$, which contradicts $v^{*}\succ v$.
+Hence, $v^{*}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{*}$.
 Altogether, we have $V_{A}^{*}\in\mathcal{PV}$ if and only if $\mathcal{PV}$ is a singleton.
 Because PO policies exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$, the SE policy exists if and only if $\mathcal{PV}=\{v^{*}\}$ is a singleton and $v^{*}\in\mathcal{V}$.
-∎
-Appendix G Proof of Lemma 6.3
-Proof.
+Appendix C Proof of Theorem˜6.3
+Proof C.1.
 We prove each statement one by one.
@@
 […]
-[…]
+$\leq\dots$
 $\leq\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}^{*}(f)}\left[V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}\right].$ (19e)
@@
 Proof of (b)
-If it holds that $Q_{A}^{f^{\prime}{\dagger}}(s,f)=V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$, by the similar derivation of Equation 37, we obtain
+If it holds that $Q_{A}^{f^{\prime}{\dagger}}(s,f)=V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}$, by the similar derivation of Equation˜19, we obtain
 $V_{A}^{f^{\prime}{\dagger}}(s_{0})=\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}^{*}(f)}\left[V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}\right]$ (20)
@@
 The policy improvement is possible if and only if $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}$.
-Because of the above equality, equivalently we can say that the policy improvement from $f^{\prime}$ is possible if and only if there exists a policy $f$ whose value function satisfies
+Because of the above equality, equivalently, we can say that the policy improvement from $f^{\prime}$ is possible if and only if there exists a policy $f$ whose value function satisfies
 $V_{A}^{f{\dagger}}(s)-V_{A}^{f^{\prime}{\dagger}}(s)\geq\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ (25)
-for all $s\in\mathcal{S}$, and there exists a state $s$ where the inequality strictly holds. ∎
-Appendix H Proof of Lemma 6.4
-Proof.
-We prove the counterpart of Lemma 6.4. Taking the negation of Equation 7 for all $f\in\mathcal{W}_{A}$, we have
+for all $s\in\mathcal{S}$, and there exists a state $s$ where the inequality strictly holds.
+Appendix D Proof of Theorem˜6.4
+Proof D.1.
+We prove the counterpart of Theorem˜6.4. Taking the negation of Equation˜5 for all $f\in\mathcal{W}_{A}$, we have
 $\neg\forall f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}\implies V_{A}^{f{\dagger}}(s)=V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}\right\}$
@@
 $\iff\exists f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}\land V_{A}^{f{\dagger}}(s)\neq V_{A}^{f^{\prime}{\dagger}}(s)~\exists s\in\mathcal{S}\right\}.$ (26b)
-From Lemma 6.3 (a), since $V_{A}^{f{\dagger}}(s)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$ holds and $Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~{}\forall s\in\mathcal{S}$ holds, we have
+From Theorem˜6.3 (a), since $V_{A}^{f{\dagger}}(s)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}$ holds and $Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}$ holds, we have
 $\exists f\in\mathcal{W}_{A}\left\{Q_{A}^{f^{\prime}{\dagger}}(s,f)\geq V_{A}^{f^{\prime}{\dagger}}(s)~\forall s\in\mathcal{S}\land V_{A}^{f{\dagger}}(s)\neq V_{A}^{f^{\prime}{\dagger}}(s)~\exists s\in\mathcal{S}\right\}$
@@
 $\implies V_{A}^{f^{\prime}{\dagger}}\notin\mathcal{PV}.$ (30)
-Therefore, $f^{\prime}$ is not a PO policy. This completes the proof. ∎
-Appendix I Proof of Lemma 6.5
-Proof.
+Therefore, $f^{\prime}$ is not a PO policy. This completes the proof.
+Appendix E Proof of Theorem˜6.5
+Proof E.1.
 First, we prove that (i) $\implies V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~\exists s\in\mathcal{S}$.
@@
 $Q_{A}^{f^{\prime}{\dagger}}(s,f)$ $=r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ (32a)
-$\geq r_{A}(s,f(s),R_{B}^{*}(s,f))+\gamma_{A}\Bigl{(}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)\right]-\delta(f,f^{\prime})\Bigl{)}$ (50b)
 […]
-$\geq\mathbb{E}^{fR_{B}^{*}(f)}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}r_{A}(s_{t},f(s_{t}),R_{B}^{*}(s_{t},f))+\gamma_{A}^{n}V_{A}^{f^{\prime}{\dagger}}(s_{n})\middle|s_{0}=s\right]-\left(\sum_{t=0}^{n-1}\gamma_{A}^{t}-\gamma_{A}^{0}\right)\delta(f,f^{\prime}).$ (50c)
+$\geq\dots$
+[…]
 for all $s\in\mathcal{S}$.
-In the limit for $n\to\infty$, the first term converges to $V_{A}^{f{\dagger}}(s)$, the second term converges to 0, and the third term converges to $-\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$.
+In the limit for $n\to\infty$, the first term converges to $V_{A}^{f{\dagger}}(s)$, the second term converges to 0, and the third term converges to $\frac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$.
 Therefore, we have
@@
 $\therefore V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)\quad\exists s\in\mathcal{S}.$ (35)
-If (ii) holds, because of Lemma 6.3 (b), it holds that
+If (ii) holds, because of Theorem˜6.3 (b), it holds that
 $V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)\quad\forall s\in\mathcal{S}.$ (36)
-Therefore,
-Considering this, it holds that
-$\forall f\in\mathcal{W}_{A}\{\mathrm{(i)~{}or~{}(ii)}\}\implies\forall f\in\mathcal{W}_{A}\left\{V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~{}~{}\exists s\in\mathcal{S}\mathrm{~{}~{}or~{}~{}}V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)~{}~{}\forall s\in\mathcal{S}\right\},$ (55)
-where the right-hand side is equivalent to the definition of PO policies (Definition 6.1). ∎
-Appendix J Proof of Theorem 6.6
-Proof.
+Therefore, it holds that
+$\forall f\in\mathcal{W}_{A}\{\mathrm{(i)~or~(ii)}\}\\ \implies\forall f\in\mathcal{W}_{A}\left\{V_{A}^{f^{\prime}{\dagger}}(s)>V_{A}^{f{\dagger}}(s)~~\exists s\in\mathcal{S}\mathrm{~~or~~}V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)~~\forall s\in\mathcal{S}\right\},$ (37)
+where the right-hand side is equivalent to the definition of PO policies (Definition˜6.1).
+Appendix F Proof of Theorem˜6.6
+Proof F.1.
 We prove each statement one by one.
+The proof of (a) is based on the monotonicity of $\{v_{t}\}$ and the compactness of $\mathrm{cl}\mathcal{V}$.
+The proof of (b) is based on the statement (b) of Theorem˜6.3.
+The statements (c) and (d) are proved by contradiction using the statement (c) of Theorem˜6.3.
 Proof of (a)
@@
 Second, we prove $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}\implies v_{t}\stackrel{{\scriptstyle s}}{{=}}v_{\infty}$. As $f_{t+1}\in\mathcal{W}_{1-\epsilon}(f_{t})$, it holds that
-$\mathcal{L}\left[V_{A}^{f_{t+1}{\dagger}}\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\geq\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\Bigl{(}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\Bigl{)}.$ (57)
+$\mathcal{L}\left[V_{A}^{f_{t+1}{\dagger}}\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$ $\geq\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]$ (39)
+$\geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\Bigl(\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\Bigl).$ (40)
 If $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$, we have $\mathcal{L}[v_{t+1}]-\mathcal{L}[v_{t}]=0$; hence, the above inequality implies
@@
 However, because $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f_{t}{\dagger}}$, implying $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$.
-From Statement (b) of Lemma 6.3, we have $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$.
+From Statement (b) of Theorem˜6.3, we have $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$.
 (Conversely, if $V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$, we have $v_{t+1}\stackrel{{\scriptstyle s}}{{=}}v_{t}$.) Thus, $\mathcal{W}_{\succcurlyeq}(f)=\mathcal{W}_{\succcurlyeq}(f_{t})$ holds for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$.
@@
 Proof of (c)
-First, we show that $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
-Suppose that there exists $v^{*}\in\mathcal{V}$ such that $\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))>\gamma_{A}\left(\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))\right)$ holds.
+First, we show that
+$\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s)))$
+for all $v\in\mathcal{V}$.
+Suppose that there exists $v^{*}\in\mathcal{V}$ such that
+$\min_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s))>\gamma_{A}(\max_{s\in\mathcal{S}}(v^{*}(s)-v_{\infty}(s)))$
+holds.
 Then, $v^{*}\succ v_{\infty}$ and there exists $f^{*}\in\mathcal{W}_{A}$ whose value function is $V_{A}^{f^{*}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v^{*}$.
@@
 We have $1\geq\tau>\gamma_{A}$.
-For an arbitrarily small $s>0$ there exists $t$ such that $\lVert v_{t}-v_{\infty}\rVert_{\infty}\leq s$ because $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$.
-Because of the continuity of $\mathcal{L}$, it also implies that for an arbitrarily small $s^{\prime}>0$ there exists $t^{\prime}$ such that $\mathcal{L}[v_{t^{\prime}}]>\mathcal{L}[v_{\infty}]-s^{\prime}$.
+For an arbitrarily small $\xi>0$ there exists $t$ such that $\lVert v_{t}-v_{\infty}\rVert_{\infty}\leq\xi$ because $v_{\infty}\stackrel{{\scriptstyle s}}{{=}}\lim_{t\to\infty}v_{t}$.
+Because of the continuity of $\mathcal{L}$, it also implies that for an arbitrarily small $\xi^{\prime}>0$ there exists $t^{\prime}$ such that $\mathcal{L}[v_{t^{\prime}}]>\mathcal{L}[v_{\infty}]-\xi^{\prime}$.
 Therefore, we can find $t$ satisfying
 $\lVert v_{\infty}-v_{t}\rVert_{\infty}$ $\leq\delta\frac{\tau-\gamma_{A}}{\gamma_{A}}\lVert v^{*}-v_{\infty}\rVert_{\infty}\quad\text{for some $\delta\in(0,1)$},\quad\text{and}$ (43)
-$\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}]$ $<(1-\epsilon)(\mathcal{L}[\bar{v}]-\mathcal{L}[v_{\infty}]),\quad\text{where}\quad\bar{v}(s):=v_{\infty}(s)+(1-\delta)(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}.$ (61)
-Let $f_{t}$ be the corresponding policy whose value function is $V_{A}^{f_{t}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v_{t}$.
+$\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}]$ $<(1-\epsilon)(\mathcal{L}[\bar{v}]-\mathcal{L}[v_{\infty}]),$ (44)
+where $\bar{v}(s):=v_{\infty}(s)+(1-\delta)(\tau-\gamma_{A})\lVert v^{*}-v_{\infty}\rVert_{\infty}.$ Let $f_{t}$ be the corresponding policy whose value function is $V_{A}^{f_{t}{\dagger}}\stackrel{{\scriptstyle s}}{{=}}v_{t}$.
 Here, we assume that $\gamma_{A}>0$.
 For the case of $\gamma_{A}=0$, Statement (c) is an immediate consequence of Statement (d) because $v_{\infty}\in\mathcal{PV}$ and $\mathcal{PV}\subseteq\partial\mathcal{V}$.
-Subsequently, we show $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}_{\infty}$ using Equation 60.
-In the proof of Statement (c) of Lemma 6.3, for all $s\in\mathcal{S}$,
+Subsequently, we show $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq\bar{v}_{\infty}$ using Equation˜43.
+In the proof of Statement (c) of Theorem˜6.3, for all $s\in\mathcal{S}$,
 $Q_{A}^{f^{\prime}{\dagger}}(s,f)-V_{A}^{f^{\prime}{\dagger}}(s)=V_{A}^{f{\dagger}}(s)-V_{A}^{f^{\prime}{\dagger}}(s)-\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$ (45)
@@
 Because $\bar{v}\succ v_{\infty}\succcurlyeq V_{A}^{f_{t}{\dagger}}$, it implies that $Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})\succcurlyeq V_{A}^{f_{t}{\dagger}}$.
-From Statement (a) of Lemma 6.3, we have $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})$.
+From Statement (a) of Theorem˜6.3, we have $V_{A}^{f^{*}{\dagger}}\succcurlyeq Q_{A}^{f_{t}{\dagger}}(\cdot,f^{*})$.
 Finally, we derive a contradiction.
@@
 $>\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}],$ (48e)
-where we used Equation 61 for the last inequality.
+where we used Equation˜44 for the last inequality.
 It implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to the fact that $v_{\infty}\succcurlyeq v_{t+1}$.
-Therefore, $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
+Therefore, we have $\min_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}(v(s)-v_{\infty}(s))\right)$ for all $v\in\mathcal{V}$.
 Subsequently, we show that $v_{\infty}\in\partial\mathcal{V}$.
 Suppose that $v_{\infty}\in\mathcal{V}\setminus\partial\mathcal{V}$ (i.e., $v_{\infty}$ is an interior of $\mathcal{V}$).
-Then, there exists an $r>0$ such that =$\mathcal{E}=\{v\in\mathcal{F}_{\mathcal{S}}\mid\lVert v-v_{\infty}\rVert_{\infty}\leq r\}\subseteq\mathcal{V}$.
+Then, there exists an $r>0$ such that $\mathcal{E}=\{v\in\mathcal{F}_{\mathcal{S}}\mid\lVert v-v_{\infty}\rVert_{\infty}\leq r\}\subseteq\mathcal{V}$.
 Let $v_{\infty}^{r}$ be such that $v_{\infty}^{r}(s)=v_{\infty}(s)+r$ for all $s\in\mathcal{S}$.
@@
 Moreover, because $v_{\infty}^{r}\in\mathcal{V}$, there exists a policy $f_{\infty}^{r}$ such that $V_{A}^{f_{\infty}^{r}{\dagger}}=v_{\infty}^{r}$.
-We lead to the contradiction to $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$. Using Equation 60, we obtain
+We lead to the contradiction to $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$. Using Equation˜43, we obtain
 $\frac{\min_{s\in\mathcal{S}}V_{A}^{f_{\infty}^{r}{\dagger}}(s)-v_{\infty}(s)}{\max_{s\in\mathcal{S}}V_{A}^{f_{\infty}^{r}{\dagger}}(s)-v_{\infty}(s)}=\frac{\min_{s\in\mathcal{S}}(v_{\infty}^{r}(s)-v_{\infty}(s))}{\max_{s\in\mathcal{S}}(v_{\infty}^{r}(s)-v_{\infty}(s))}=1,$ (49)
@@
 $>\mathcal{L}[v_{\infty}]-\mathcal{L}[v_{t}],$ (52d)
-where the derivation of Equation 69a is due to the definition of $\mathcal{W}_{1-\epsilon}(f_{t})$ in Equation 9 and $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}$.
+where the derivation of Equation˜52a is due to the definition of $\mathcal{W}_{1-\epsilon}(f_{t})$ in Equation˜7 and $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f{\dagger}}$.
 This implies $\mathcal{L}[v_{t+1}]>\mathcal{L}[v_{\infty}]$, which contradicts to $v_{\infty}\succcurlyeq v_{t+1}$.
 Therefore, $v_{\infty}\in\mathcal{PV}$.
-∎
-Appendix K Proposed Algorithm with Backtracking
-Algorithm 1 may not output a PO policy because the terminate condition in line 5 of Algorithm 1 is the necessary condition for PO policies.
-In this case, Lemma 6.5 can be used to determine whether the output policy is a PO policy: provided (i) or (ii) in Lemma 6.5 holds for all $f\in\mathcal{W}_{A}$ with the output policy $f_{t}$ and its state value function $V_{A}^{f_{t}{\dagger}}$, $f_{t}$ is a PO policy.
+Appendix G Splitting the Search Space
+The proposed algorithm in Algorithm˜1 can be impractical when the leader’s policy space $\mathcal{W}_{A}$ is prohibitively large to search for the next policy in Equation˜6 and Equation˜8.
+To avoid the exhaustive search, we propose a splitting strategy for the search space to obtain improved policies from each subspace efficiently.
+We first split $\mathcal{W}_{A}$ into disjoint sets $\mathcal{W}_{1}\cup\dots\cup\mathcal{W}_{K}$, where the best response is $R_{B}^{*}(f)=g_{i}^{*}\in\mathcal{W}_{B}^{d}$ for $f\in\mathcal{W}_{i}$.
+Because the set of deterministic policies, $\mathcal{W}_{B}^{d}$, is finite, we can always find such a separation.
+Additionally, we argue that the number of deterministic policies that form the best response is rather limited, enabling us to enumerate them easily.
+We consider replacing (8) with locating a Pareto optimal $f$ with respect to $Q_{A}^{f_{t}\dagger}(\cdot,f)$.
+Because (8) also locates Pareto optimal $f$, this replacement corresponds to selecting a different $\mathcal{L}$ at each $t$.
+This is realized in two steps.
+First, we locate a policy satisfying $Q_{A}^{f_{t}\dagger}(\cdot,f)\succcurlyeq V_{A}^{f_{t}\dagger}$.
+It is nontrivial as the best response $R_{B}^{*}$ changes as $f$ changes.
+However, if we limit our attention to a subset $\mathcal{W}_{k}$ for some $k$, the best response does not change with $f\in\mathcal{W}_{k}$.
+Then, locating $f\in\mathcal{W}_{k}$ satisfying $Q_{A}^{f_{t}\dagger}(\cdot,f)\succcurlyeq V_{A}^{f_{t}\dagger}$ is casted as the following constrained optimization problem:
+$\min_{z,f\in\mathcal{W}_{k}}z\quad\text{s.t.}\quad V_{A}^{f_{t}{\dagger}}(s)-Q_{A}^{f_{t}\dagger}(s,f)\leqslant z\quad\text{for all }s\in\mathcal{S},$ (53)
+where $Q_{A}^{f_{t}\dagger}(s,f)$ is differentiable with respect to $f$ and is easily computable as the best response is constant.
+Therefore, it will be solved by, e.g., a gradient-based solver with projection onto $\mathcal{W}_{k}$.
+A solution with $z\leq 0$ satisfies $Q_{A}^{f_{t}\dagger}(\cdot,f)\succcurlyeq V_{A}^{f_{t}\dagger}$.
+Such a solution must exist in $\mathcal{W}_{k}$ for some $k$ because $f_{t}$ satisfies this condition.
+Once we find such a solution, denoted as $(z_{t},f_{t}^{\prime})$, we find a Pareto optimal policy for $Q_{A}^{f_{t}\dagger}(\cdot,f)$ in $\mathcal{W}_{k}\cap\mathcal{W}_{\succcurlyeq}(f_{t})$ by improving $f_{t}^{\prime}$.
+It is done by performing a Pareto ascent method with projection,
+$f\leftarrow\Pi_{\mathcal{W}_{k}}(f+\eta\Delta),\quad\text{where}\quad\Delta^{\top}\nabla Q_{A}^{f_{t}\dagger}(s,f)\geqslant 0\quad\text{for all }s\in\mathcal{S},$ (54)
+or its variants, e.g., Harada et al.
+(2006), with $f_{t}$ as a feasible initial solution of this step.
+If we find $f_{t}^{*}$ such that $Q_{A}^{f_{t}\dagger}(\cdot,f)\succ V_{A}^{f_{t}\dagger}$, we let $f_{t+1}=f_{t}^{*}$.
+These steps will be continued until a predefined termination criterion is satisfied, or no policy strictly dominating the current policy is found.
+The whole algorithm is shown in Algorithm˜2.
+Algorithm 2 Practical Pareto-Optimal Policy Iteration
+0: Maximum number of iterations $M$.
+1: Split the leader’s policy space $\mathcal{W}_{A}$ into disjoint sets $\mathcal{W}_{1}\cup\dots\cup\mathcal{W}_{K}$, where the best response is $R_{B}^{*}(f)=g_{i}^{*}\in\mathcal{W}_{B}^{d}$ for $f\in\mathcal{W}_{i}$
+2: Randomly initialize $f_{0}\in\mathcal{W}_{A}$ and compute $g_{0}=R_{B}^{*}(f_{0})$
+3:for $t=0$ to M1M-1 do
+4:  Compute $V_{A}^{f_{t}{\dagger}}$ by repeatedly applying $T_{A}^{f_{t}R_{B}^{*}(f_{t})}$
+5:  Let $\mathcal{I}_{t}=\{1,\dots,K\}$
+6:  while $\lvert\mathcal{I}_{t}\rvert>0$ do
+7:   Randomly select $k\in\mathcal{I}_{t}$
+8:   Solve (53) to obtain $z_{t},f_{t}^{\prime}$
+9:   if $z_{t}\leq 0$ then
+10:    Perform the Pareto ascent (54) to obtain $f_{t}^{*}$
+11:    Let $f_{t+1}=f_{t}^{*}$ and break if $Q_{A}^{f_{t}\dagger}(\cdot,f_{t}^{*})\succ V_{A}^{f_{t}{\dagger}}$
+12:   end if
+13:   Update $\mathcal{I}_{t}\leftarrow\mathcal{I}_{t}\setminus\{k\}$
+14:  end while
+15:   Let $f_{M}=\dots=f_{t+1}=f_{t}$ and break if $\mathcal{I}_{t}=\emptyset$
+16:end for
+16: A stationary policy $f_{M}$
+Appendix H Backtracking
+Algorithm˜1 may not output a PO policy because the terminate condition in line 5 of Algorithm˜1 is the necessary condition for PO policies.
+In this case, Theorem˜6.5 can be used to determine whether the output policy is a PO policy: provided (i) or (ii) in Theorem˜6.5 holds for all $f\in\mathcal{W}_{A}$ with the output policy $f_{t}$ and its state value function $V_{A}^{f_{t}{\dagger}}$, $f_{t}$ is a PO policy.
 Notably, this judgment is not perfect because the policy may be determined not to be a PO policy when it really is a PO policy.
-We proposed selecting another policy $f^{\prime}\in\mathcal{W}(f_{t-1})$ that has not been selected thus far and restarting the for loop with $f_{t}\leftarrow f^{\prime}$, which we call ”Backtracking”, to deal with the case where the output policy $f_{t}(t<M)$ is determined not to be a PO policy.
+We proposed selecting another policy $f^{\prime}\in\mathcal{W}(f_{t-1})$ that has not been selected thus far and restarting the for loop with $f_{t}\leftarrow f^{\prime}$, which we call "Backtracking", to deal with the case where the output policy $f_{t}(t<M)$ is determined not to be a PO policy.
 If there is no policy left in $\mathcal{W}(f_{t-1})$, then it backtracks one by one, as $\mathcal{W}(f_{t-2}),\mathcal{W}(f_{t-3}),\dots$, until it reaches the set with policies that have never been selected.
 This method requires that $\mathcal{W}(f_{t})$ is saved at each iteration.
-The entire proposed algorithm including the PO-policy judgement and the backtracking is shown in Algorithm 2.
+The entire proposed algorithm with the PO-policy judgement and the backtracking is shown in Algorithm˜3.
 $W$ is a list to save each $\mathcal{W}(f_{t})$ and $M$ is the maximum number of iterations.
@@
 If the PO-policy judgment in line 9 does not return True during $M$ iterations, the algorithm outputs the policy that maximizes the objective $\sum_{s\in\mathcal{S}}\alpha_{s}V_{A}^{f_{t}{\dagger}}(s)$ among the ones that satisfy the necessary condition for PO policies thus far and $f_{M}$ (see lines 12, 25, and 26).
-Algorithm 2 PIPOP with Backtracking
+Algorithm 3 Pareto-Optimal Policy Iteration with Backtracking
 0: Pareto-compliant scalarization $\mathcal{L}$, maximum number of iterations $M$, a $M$-length list of empty sets $W$.
-1:  $W(0)\leftarrow\mathcal{W}_{A}$
+[…]
 2: Randomly sample $f_{0}$ from $W(0)$ without replacement.
-3:  $\mathcal{W}_{output}\leftarrow\{\}$
-4:  for $t=0$ to $M-1$ do
+[…]
+4:for $t=0$ to M1M-1 do
 5:  Compute $V_{A}^{f_{t}{\dagger}}\in\mathcal{F}(\mathcal{S})$ such that $(T_{A}^{f_{t}R_{B}^{*}(f_{t})}V_{A}^{f_{t}{\dagger}})(s)=V_{A}^{f_{t}{\dagger}}(s)$ for all $s\in\mathcal{S}$.
@@
 8:  if $\mathbb{E}_{a\sim f(s)}\left[Q_{A}^{f_{t}{\dagger}}(s,a,R_{B}^{*}(s,f))\right]=V_{A}^{f_{t}{\dagger}}(s)~\forall s\in\mathcal{S},\forall f\in\mathcal{W}_{\succcurlyeq}(f_{t})$ then
-9:        if $f_{t}$ satisfies (i) or (ii) of Lemma 6.5 for all $f\in\mathcal{W}_{A}$ then
+9:   if $f_{t}$ satisfies (i) or (ii) of Theorem˜6.5 for all $f\in\mathcal{W}_{A}$ then
 10:    return $f^{*}\leftarrow f_{t}$.
@@
 26: A stationary policy $f^{*}$
-Appendix L Relation between the method of Zhang et al. (2020a) and Bucarey et al. (2022)
-Zhang et al.Zhang et al.
-(2020a) defined the optimal value functions as $V_{i}^{*}(i\in\{A,B\})$, which are the solution of the following equationsZhang et al.
-(2020a): let $a\in\mathcal{A},b\in\mathcal{B},\gamma\in[0,1)$, and for all $s\in\mathcal{S}$,
+Appendix I Relation between the methods of Zhang et al. (2020) and Bucarey et al. (2022)
+Zhang et al.
+(2020a) defined the optimal value functions as $V_{i}^{*}(i\in\{A,B\})$, which are the solution of the following equations (Zhang et al., 2020a): let $a\in\mathcal{A},b\in\mathcal{B},\gamma\in[0,1)$, and for all $s\in\mathcal{S}$,
 $Q_{i}^{*}(s,a,b)=r_{i}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b)}\left[V_{i}^{*}(s^{\prime})\right],$ (55)
-$V_{i}^{*}(s)=Stackelberg_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s)),$ (71)
+$V_{i}^{*}(s)=\textsc{Stackelberg}_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s)),$ (56)
 where $Q_{i}^{*}(s)$ is the Q-value $Q_{i}(a,b,s)$ of an action pair $(a,b)\in\mathcal{A}\times\mathcal{B}$ in state $s$.
 Given the fixed $s$, $Q_{i}^{*}(b,a,s)$ can be viewed as the payoff of $(a,b)$ for the agent $i$; thus, $Q_{i}^{*}(s)$ can be viewed as the payoff table of agent $i$ in the state $s$.
-$Stackelberg_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s))$ is the payoff for agent $i$ in the Stackelberg equilibrium of the normal-form game represented by the payoff tables $(Q_{A}^{*}(s),Q_{B}^{*}(s))$.
-Therefore, Equation 71 can be written as
+$\textsc{Stackelberg}_{i}(Q_{A}^{*}(s),Q_{B}^{*}(s))$ is the payoff for agent $i$ in the Stackelberg equilibrium of the normal-form game represented by the payoff tables $(Q_{A}^{*}(s),Q_{B}^{*}(s))$.
+Therefore, Equation˜56 can be written as
 $V_{i}^{*}(s)$ $=Q_{i}^{*}(s,R_{A}^{\prime}(s),R_{B}^{\prime}(s,R_{A}^{\prime}(s))),$ (57a)
@@
 $R_{B}^{\prime}(s,a):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}Q_{B}^{*}(s,a,b).$ (57c)
-By substituting Equation 70 to Equation 72, we obtain
+By substituting Equation˜55 to Equation˜57, we obtain
 $V_{i}^{*}(s)$ $=r_{i}(s,R_{A}^{\prime}(s),R_{B}^{\prime}(s,R_{A}^{\prime}(s)))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}^{\prime}(s),R_{B}(s,R_{A}^{\prime}(s)))}\left[V_{i}^{*}(s^{\prime})\right],$ (58a)
@@
 $R_{B}^{\prime}(s,a):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b))}\left[V_{B}^{*}(s^{\prime})\right].$ (58c)
-Meanwhile, the operator used in Bucarey et al.Bucarey et al. (2022) is given by
-$(Tv)_{i}(s)$ $=r_{i}(s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,v),R_{B}(s,R_{A}(s,v),v_{b}))}\left[v_{i}(s^{\prime})\right],$ (74a)
-$\mathrm{where~{}}$ $R_{A}(s,v):=\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}r_{A}(s,f_{s},R_{B}(s,f_{s},v_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},R_{B}(s,f_{s},v_{B}))}\left[v_{A}(s^{\prime})\right],$ (74b)
-$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right].$ (74c)
+Meanwhile, the operator used in Bucarey et al. (2022) is given by
+[…]
+[…]
+$R_{B}(s,f_{s},v_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,f_{s},b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f_{s},b)}\left[v_{B}(s^{\prime})\right].$ (59g)
 Notably, $\gamma=\gamma_{A}=\gamma_{B}$.
-Let us consider only the deterministic leader policies in the fixed point of Equation 74 and replace $\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}$ to $\operatornamewithlimits{argmax}_{a\in\mathcal{A}}$ in the definition of $R_{A}$.
-Then, letting $(V_{A},V_{B})$ be the fixed point of $T$, it holds from Equation 74, for all $s\in\mathcal{S}$, which is expressed as
-$V_{i}(s)$ $=r_{i}(s,R_{A}(s,V),R_{B}(s,R_{A}(s,V),V_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,R_{A}(s,V),R_{B}(s,R_{A}(s,V),V_{B}))}\left[V_{i}(s^{\prime})\right],$ (75a)
-$\mathrm{where~{}}$ $R_{A}(s,V):=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}r_{A}(s,a,R_{B}(s,a,V_{B}))+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,R_{B}(s,a,V_{B}))}\left[V_{A}(s^{\prime})\right],$ (75b)
-$R_{B}(s,a,V_{B}):=\operatornamewithlimits{argmax}_{b\in\mathcal{B}}r_{B}(s,a,b)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a,b)}\left[V_{B}(s^{\prime})\right].$ (75c)
-Comparing these with Equation 73, the definition of $(V_{A}^{*},V_{B}^{*})$ by Equation 73 is equivalent to the definition of $(V_{A},V_{B})$ by Equation 75.
-It follows that the fixed point in Zhang et al.Zhang et al.
-(2020a) is the same in Bucarey et al.Bucarey et al.
+Let us consider only the deterministic leader policies in the fixed point of Equation˜59 and replace $\operatornamewithlimits{argmax}_{f_{s}\in\Delta(\mathcal{A})}$ with $\operatornamewithlimits{argmax}_{a\in\mathcal{A}}$ in the definition of $R_{A}$.
+Then, letting $(V_{A},V_{B})$ be the fixed point of $T$, it holds from Equation˜59, for all $s\in\mathcal{S}$, which is expressed as
+[…]
+[…]
+[…]
+Comparing these with Equation˜58, the definition of $(V_{A}^{*},V_{B}^{*})$ by Equation˜58 is equivalent to the definition of $(V_{A},V_{B})$ by Equation˜60.
+It follows that the fixed point in Zhang et al.
+(2020) is the same in Bucarey et al.
 (2022) when the leader policies in the fixed point are restricted to deterministic stationary policies.
-By applying the discussion in the proof of Theorem 5.1 to $(V_{A},V_{B})$ defined by Equation 75, we obtain
+By applying the discussion in the proof of Theorem˜5.1 to $(V_{A},V_{B})$ defined by Equation˜60, we obtain
 $V_{A}(s)=\max_{f\in\mathcal{W}_{A}^{d}}V_{A}^{fR_{B}(f,V_{B})}(s),\quad V_{B}(s)=\max_{g\in\mathcal{W}_{B}}V_{B}^{R_{A}(V)g}(s),$ (61)
@@
 Then, the problem of the method of Bucarey et al.
-discussed in Section 5.1 also occurs to $(V_{A},V_{B})$: $V_{A}$ does not coincide with the state value function of the deterministic SSE policies since $R_{B}(f,V_{B})$ is not guaranteed to be the best response against any $f\in\mathcal{W}_{A}^{d}$.
-Appendix M Application: Policy Teaching by Intervention to the Transitions
+discussed in Section˜5 also occurs to $(V_{A},V_{B})$: $V_{A}$ does not coincide with the state value function of the deterministic SSE policies since $R_{B}(f,V_{B})$ is not guaranteed to be the best response against any $f\in\mathcal{W}_{A}^{d}$.
+Appendix J Application: Policy Teaching by Intervention to the Transitions
 In single-agent MDPs, the policy aimed by the agent can be changed by making changes to the agent’s observations, reward signals, and transition probabilities, among others.
 For example, in an MDP $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$, the changes in the observations and the transition probabilities are represented by replacement to another transition function $p^{\prime}$, which changes the agent’s aim from the optimal policies of $(\mathcal{S},\mathcal{A},p,\rho,r,\gamma)$ to those of $(\mathcal{S},\mathcal{A},p^{\prime},\rho,r,\gamma)$.
-This is the theoretical foundation of poisoning attacks against RL Behzadan & Munir (2017); Huang & Zhu (2019); Zhang et al.
-(2020b); Rakhsha et al.
-(2020); Sun et al.
-(2020) and policy teaching Zhang & Parkes (2008); Zhang et al.
-(2009), which aims to guide the agent’s policy learning.
-SSGs can be viewed as a model of intervention in transition probabilities by the leader where the follower is the victim of the attack or the teaching.
+This is the theoretical foundation of poisoning attacks against RL (Behzadan and Munir, 2017; Huang and Zhu, 2019; Zhang et al., 2020b; Rakhsha et al., 2020; Sun et al., 2020) and policy teaching (Zhang and Parkes, 2008; Zhang et al., 2009), which aims to guide the agent’s policy learning.
+SSGs can be viewed as a model of intervention in transition probabilities by the leader, where the follower is the victim of the attack or the teaching.
 Even if the follower takes the same action at each state, the transition can be shifted by the leader changing its actions because the transition probability depends on the actions of all the agents.
@@
 where $\hat{r}_{A}$ is the leader’s reward function that represents the desired follower’s behavior intended by the leader.
-$f^{*}$ is the optimal solution of Equation 77, and the state-action sequences generated by the corresponding follower’s optimal policy $R_{B}^{*}(f)$ maximize not only the follower’s expected cumulative discounted reward but also that of the leader’s.
+$f^{*}$ is the optimal solution of Equation˜62, and the state-action sequences generated by the corresponding follower’s optimal policy $R_{B}^{*}(f)$ maximize not only the follower’s expected cumulative discounted reward but also that of the leader’s.
 This can be viewed as the follower attaining the leader’s desired behavior when the follower achieves its own learning goal.
-The optimal solution of Equation 77 is given by the PO policy $f$ that maximizes $\mathcal{L}[V_{A}^{f{\dagger}}]=\sum_{s\in\mathcal{S}}\rho(s)V_{A}^{f{\dagger}}(s)$.
+The optimal solution of Equation˜62 is given by the PO policy $f$ that maximizes $\mathcal{L}[V_{A}^{f{\dagger}}]=\sum_{s\in\mathcal{S}}\rho(s)V_{A}^{f{\dagger}}(s)$.
 Therefore, our proposed approach can be applied to this problem when the transition function is known and the follower’s best response can be computed.
 Although the optimality of the obtained solution is not guaranteed unless $\gamma_{A}=0$, our analysis guarantees the monotone improvement of the leader’s policy, which is of great importance in practice.
-In contrast, existing methods trying to obtain the SSE policy may be inadequate because it is usual that the follower is not myopic and hence the SSE policy does not necessarily exist.
+In contrast, existing methods trying to obtain the SSE policy may be inadequate because it is usual that the follower is not myopic, and hence, the SSE policy does not necessarily exist.